A Complete Bibliography of Statistics in Medicine (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

26 June 2021 Version 1.01

Title word cross-reference

$106 [785]. $12 [788]. $26.75 [708]. $29 [693]. 2 × 2 [745, 334, 462, 200, 618, 813]. $32.95 [363]. $34.50 [456]. $35 [151]. $45.00 [234]. $46 [749]. $48.95 [518]. $49.90 [150]. $69.75 [692, 621]. $85.00 [253]. $88.00 [125]. $89.25 [253]. F [265]. K [813, 426]. λ [704]. P [814, 803, 618]. t [250, 348, 424].

-ratio [426]. -test [348, 424]. -value [618, 803]. -values [814].

13.95 [691]. 14.95/$29.60 [516]. 15.00 [269]. 1842 [459]. 19.65 [607]. 1978 [253, 253]. 1982 [67]. 1983 [150]. 1984 [217]. 1985 [308]. 1986 [441, 485]. 1987 [518]. 1987p [595]. 1988 [793, 668].

20 [581, 645]. 20.00 [806]. 209 [137]. 22.50 [560, 709]. 22.95 [235]. 24-hour [548]. 29.95 [364]. 2nd [789, 788, 708].

30 [690, 723, 619]. 33.90 [561]. 35 [786]. 38.35/$55.00 [517].

4 [789]. 40 [706]. 40.00 [305]. 45 [787]. 4XY [789].

50.00 [529]. 58.00 [365]. 580 [34].

7 [855]. 7.50 [214]. 7TQ [790].

8 [854]. 80 [316].

9.25 [790]. 9.95 [707, 620].

A. [39]. abdominal [465]. aberrations [660]. aborter [326]. abortion [286, 326, 244]. absolute [712]. accident [588]. according [231]. account [447]. accrual [316]. Accumulating [371]. accuracy [70]. Acheson [253]. Acknowledgement [648]. acquired [508, 289]. action [456]. active [687]. acute [105, 465, 7]. addition [744]. additive [229]. Adelin [621]. adequacy [240]. adjust [652]. adjusted [361]. Adjusting [771, 111]. Adjustment [329, 110]. Adjustments [413]. administration [709]. admissions [286]. adolescence [45]. adverse [20, 585, 602, 454]. aetiology [122]. affect [472, 332]. affects [683, 684]. after [189]. against [618]. Age [410, 411, 29, 322, 437, 110, 564, 328, 454, 587]. age-adjustment [110]. age-cohort [410]. Age-period [410]. Age-period-cohort [411]. age-specific [328]. Agency [723]. agents [312]. aggregate [551]. aggregation [486]. aging [472]. agree [442]. agreement [773, 191]. aid [186, 105, 278, 449, 7]. aided [320]. aids [743, 79, 632, 634, 626, 300, 633, 818, 628, 625, 627]. air [760, 759]. Aitkin [807]. Albert [621]. alcoholic [223]. Aldershot [269]. Algorithm [276, 90]. alleged [814]. Allmut [365]. allocated [603]. allocating [850]. Allocation [296, 141, 63, 284, 202, 685, 354]. allowance [533]. alpha [673]. alpha-fetoprotein [673]. alternative [520, 265, 618]. Altman [51]. ambulatory [548]. American [693]. among [699, 289, 191, 579, 823, 15]. anaemia [81, 837]. anaesthesia [575]. analyse [773]. analyses [248, 231, 146, 323, 398, 528, 104, 11, 5, 123]. Analysing [109, 481, 500, 478, 473, 108, 639, 448, 474, 21, 524]. Analysis [502, 765, 87, 344, 655, 142, 249, 414, 777, 657, 498, 122, 597, 739, 247, 63, 517, 545, 394, 97, 578, 232, 239, 81, 756, 387, 442, 316, 303, 514, 735, 425, 265, 82, 242, 563, 778, 196, 797, 850, 787, 810, 723, 326, 133, 736, 88, 148, 522, 284, 114, 555, 417, 482, 226, 643, 137, 464, 462, 136, 134, 287, 447, 480, 579, 118, 211, 731, 826, 746, 85, 548, 611, 817, 318, 269, 223, 94, 209, 748, 610, 427, 602, 144, 14, 493, 654, 823, 259, 243, 476, 645, 614, 193, 78, 663, 798]. analysis [537, 700, 33, 124, 310, 342, 358, 57, 820, 587, 15, 523, 437, 72, 780, 331, 194, 263, 479, 76, 812, 98, 664, 529, 692, 693]. analytical [234, 71]. Anastasios [854]. Ancestry [37]. ancillary [842]. Anderson [807]. Andrew [269]. animal [11, 645]. Announcement [12, 35, 106, 183, 198, 236, 254, 268, 292, 418, 515, 540, 592, 608, 769, 808, 646, 571]. Announcements [64, 92, 281, 350, 367, 530]. answers [381]. anthropometric [232]. antiarrhythmic [97, 801]. antibody [109]. antigen [673]. antigens [61, 513]. Anwendung [365]. Apology [23]. apparent [218, 104]. appearance [94]. applicability [852, 851]. Application [765, 447, 85, 71, 310, 120, 587, 15, 640, 247, 77, 232, 242, 805, 148, 662, 543, 11, 244, 652, 564, 227, 123, 56, 79, 200, 356, 812]. Applications [364, 718, 331, 150]. applied [97, 348, 600, 29, 692]. approach [63, 653, 449, 712, 431, 743, 70, 826, 508, 85, 318, 506, 56, 257, 330, 356, 315]. Approaches [139, 213, 248, 451, 821, 567]. approximation [362]. arbitrarily [96]. arising [344]. ARMA [359]. Armitage [560]. Arnold [691]. art [591]. arteriographic [333]. artery [168, 192]. arthritis [513]. Arthur [825]. ascertained [205]. ascertainment [205]. ascertainment-bias [205]. Aspects [493, 444, 668]. aspirin [377, 200]. assay [249]. assays [559, 69]. Assessing [218, 509, 703, 240, 309, 80, 801, 353, 572, 301, 602, 438, 627]. Assessment [228, 264, 338, 445, 407, 658, 221, 375, 805, 760, 759, 69, 58, 543, 162, 615]. assistance [224]. associated [509, 300, 277, 508, 818]. Association [51, 790, 773, 61, 300, 492, 704, 122]. associations [513]. assumptions [433]. assurance [366, 844]. asymptotic [555]. atlas [706, 253, 253]. attention [284]. attitudes [834]. attributable [549, 131, 436, 589, 525, 28]. August [793]. Austin [37]. Australia [788]. Australian [788]. Austria [793]. Author [716]. Automatic [145]. Autoregressive [476, 477, 243, 820]. average [538]. avoidable [706]. Avoidance [375].

B [291, 456, 709, 39, 807, 673]. bacillus [45]. background [33]. Bailar [581]. Bailey [125]. balanced [613]. balancing [141]. bank [271]. Barker [253]. base [329, 777]. based [63, 658, 109, 653, 512, 449, 459, 17, 508, 746, 319, 147, 802, 659, 257, 836, 225]. Basel [708, 137, 749]. baseline [489, 488, 771, 462, 285]. baselines [796]. bases [672]. basic [25]. Basingstoke [305]. basis [775, 294]. batch [735]. batch-to-batch [735]. Bayes [465, 320, 208]. Bayesian [247, 248, 611, 748, 288]. BCG [45]. be [174]. behaviour [346]. behavioural [826]. belts [667]. benefits [4]. Benjamin [785]. Berlin [485]. Berry [560]. best [70]. Betty [707]. Between [68, 61, 88, 188, 300, 8, 583, 428, 666, 701, 124, 120]. Beyond [617]. Bias [131, 361, 60, 212, 387, 698, 613, 192, 783, 205, 574, 342, 120, 510]. biased [284]. Biases [407, 375]. Bibliography [50]. Bimodality [84]. binary [578, 422, 564, 777, 534, 763]. binding [734]. binomial [573, 604, 300]. bioassay [749]. biological [359, 274, 275, 21, 206, 207]. biology [693]. biomathematics [125]. biomedical [787, 583, 145]. biopsies [57]. Biostatistics [67, 217, 308, 441, 595, 793, 607, 235]. Birth [193, 143, 267]. birthweight [305, 15]. black [511]. Blackwell [560, 707, 619]. Bland [620]. bleeding [320]. Bloch [364, 234]. blocked [191]. blood [230, 475, 548, 84, 511]. BMA [516]. Book [291, 581, 517, 789, 788, 806, 51, 560, 607, 456, 708, 518, 706, 529, 690, 125, 305, 709, 691, 364, 787, 561, 786, 790, 692, 723, 34, 363, 151, 234, 621, 785, 825, 235, 516, 137, 693, 826, 548, 707, 269, 150, 253, 620, 749, 619, 645, 214, 807, 365]. Books [214, 581]. Bootstrap [725, 190, 301]. Bootstrapped [740]. born [437]. Boston [561]. both [436]. bovine [719]. Box [788, 261]. Bradford [37, 48, 38, 44, 43]. Bradley [785]. brain [421]. breast [4, 159, 14]. Breslow [723]. brief [444, 832]. British [51]. broken [17]. broken-line [17]. bronchodilators [796]. Brown [561]. business [709]. Buyse [151]. bypass [168].

C [581, 607, 826, 253]. C.T. [224]. Calculating [297, 149, 102, 197]. calculation [565, 424]. Calendar [772]. Calibrated [313]. Cambridge [291, 456, 825]. Campbell [619]. Can [100, 177, 371]. Cancer [679, 385, 765, 229, 220, 410, 411, 4, 108, 25, 723, 88, 366, 10, 799, 137, 447, 227, 159, 312, 29, 150, 157, 14, 253, 402, 523, 208, 324, 460, 76, 490, 788, 151, 366]. cancers [459]. carbamazepine [795]. carcinogenic [565]. carcinoma [228, 673]. cardiac [57]. Cardiff [441, 455]. cardiologists [834]. Cardiovascular [680, 672, 383]. care [166, 162]. caries [451, 509]. Carlo [143, 267, 346, 433, 133]. Carr [305]. Carr-Hill [305]. carriers [673]. carry [104, 591]. carry-over [104, 591]. Carter [137]. Case [105, 273, 119, 288, 765, 316, 82, 345, 7, 821, 228, 329, 555, 525, 505, 652, 611, 781, 240, 296, 777, 748, 657, 260, 262, 834, 331, 463, 631]. case-base [329, 777]. Case-control [273, 288, 765, 345, 821, 555, 525, 505, 611, 240, 296, 748, 119, 260]. case-heterogeneity [262]. cases [289, 460]. categorical [774, 484, 757, 577, 692, 726, 104, 480, 479]. causal [556, 15]. Cause [263, 416, 204]. Cause-deleted [263, 416, 204]. caused [613, 783, 574, 500]. causes [111]. cautionary [820]. Cautions [715, 99, 713]. cell [837]. censored [336, 302, 265, 779, 89, 32, 96, 90, 598]. Censoring [287, 741, 614, 500]. central [845, 523, 747]. centre [141]. centres [271]. cerebral [224, 98]. certain [61, 602]. certificates [671]. cervical [447]. Chairman [503, 471, 853]. Chalmers [595]. Change [196, 771, 699, 572, 104, 427, 507, 804]. change-over [104]. Change-point [196, 507, 804]. changeover [736, 522]. changes [500]. changing [834, 324]. Chapman [826]. characterize [699, 508]. Charles [125]. CHD [596]. chemistry [231, 408]. chemosensitivity [249]. chemotherapy [137]. chest [123]. chi [146, 413]. chi-square [146, 413]. Chichester [517, 364, 787, 234, 516, 253]. childhood [475]. choice [286, 566, 443, 124]. choices [310]. Cholesterol [658]. Cholesterol-based [658]. Choosing [260, 5]. chronic [221, 316, 226, 223, 641]. cigarette [490]. Claire [786]. class [242, 417, 226, 144, 259]. Classical [248]. classification [221, 321, 144, 56, 145, 356]. classifying [123]. clinic [846]. Clinical [67, 217, 308, 441, 595, 793, 121, 46, 790, 49, 312, 517, 338, 788, 47, 186, 636, 248, 246, 278, 337, 231, 638, 154, 835, 423, 218, 174, 377, 387, 442, 698, 220, 316, 303, 514, 601, 797, 842, 73, 177, 392, 364, 850, 30, 139, 213, 567, 847, 679, 854, 108, 48, 16, 74, 266, 832, 568, 17, 433, 398, 396, 284, 151, 621, 313, 678, 443, 825, 681, 166, 142, 535, 366, 612, 448, 235, 10, 202, 58, 249, 162, 85, 301, 729, 615, 5, 319, 833, 506, 223, 179, 343, 116, 533, 309, 844, 408]. clinical [852, 682, 701, 688, 71, 619, 75, 811, 685, 796, 421, 700, 310, 79, 257, 311, 298, 848, 534, 463, 843, 76, 315, 812, 851, 839, 680, 381, 789, 529, 790]. clinically [88]. clinician [42]. CLL [316]. closed [147]. Closing [675]. Cluster [855, 344, 702, 605, 822]. clustered [413]. clustering [135, 600, 524]. Co [125, 561, 235, 269, 387, 141]. co-operative [387]. co-ordinating [141]. Coates [788]. cohort [100, 229, 410, 411, 723, 222, 296, 29, 767, 493, 744, 322, 193, 56, 342, 437]. cohorts [461]. coin [284]. Colebatch [788]. Colin [305]. collect [843]. collected [174]. collection [846, 847]. College [441]. colon [765]. combination [821, 815, 432, 194]. combining [636, 423, 392, 295]. Comment [715, 714]. Commentary [489, 170, 760, 494, 400, 661, 501, 95, 495, 182, 389, 390, 656, 665, 181, 483, 169]. comments [655, 761]. committee [601]. common [702]. commonly [844]. community [706, 187, 836]. community-based [836]. compact [601]. company [607]. Comparability [671]. Comparative [209, 473]. compare [164, 271]. Comparing [573, 584, 604]. Comparison [143, 267, 230, 816, 488, 102, 222, 733, 357, 205, 91, 670, 149, 745, 302, 636, 239, 387, 442, 734, 196, 146, 197, 10, 687, 304, 813, 117, 78, 446, 696, 251, 76, 812, 489]. comparisons [683, 426, 361, 695, 666, 684, 347, 721]. compartment [747]. competing [447]. Complementary [579, 823, 445]. complex [82]. compliance [638, 841]. complications [447]. components [815]. computation [521, 96]. Computer [278, 224, 601, 558, 320, 331]. computer-aided [320]. computerized [309]. concentrations [795]. concept [278]. Concluding [853, 503]. conditions [187, 19]. conduct [529]. Confidence [566, 556, 336, 740, 728, 434, 779, 655, 147, 754]. confidential [629]. Confirmatory [464]. confounding [590, 717]. Confronting [342]. congenital [327, 54]. Congress [459]. Congressi [217]. Congruity [428]. consent [140]. consideration [569]. Considerations [290, 427, 260, 836, 314, 801]. consistent [436]. Constructing [800, 728]. consultant [456]. consulting [116]. contact [289, 631]. contiguous [276]. Contingency [536, 334]. continuation [512]. continuous [576, 505]. contraceptive [317]. contribution [459, 43]. contributions [48]. Control [387, 765, 231, 273, 838, 345, 821, 228, 156, 555, 69, 525, 505, 611, 240, 296, 408, 748, 157, 717, 119, 260, 288]. controlled [637, 840]. controls [229, 142, 744, 260]. conventional [776]. Converting [712]. cooperative [845]. coronary [105, 174, 377, 658, 7, 168, 192]. Correcting [353, 743, 783, 574, 438, 510, 730]. Correction [127, 185, 256, 420, 855, 754, 205, 60, 212, 752]. Corrections [647]. correlated [816, 188]. correlation [114, 8]. correlations [429, 816]. Correspondence [797, 579, 223, 823]. Cost [835, 836, 337, 850, 431, 847, 849]. cost-effectiveness [850]. cost-function [431]. Costs [832]. counting [318]. countries [836]. course [318]. Covariance [756, 498, 33, 124]. Covariate [685, 489, 765, 218, 613, 488, 343, 767, 614, 264]. cover [807]. Cox [725, 819, 740, 239, 190, 264]. criteria [682, 421, 840]. criterion [302]. Critical [221, 75]. Cross [591, 422, 462, 547]. Cross-over [591, 422, 462]. cross-sectional [547]. crossover [810, 687]. crosssectional [193]. cubic [697]. curability [14]. Cure [412, 133]. Curtis [529]. curve [569, 277, 11]. curve-fitting [277]. curves [63, 802, 780, 477]. cusum [415, 580, 244, 71]. cut [800]. cut-off [800].

D [607, 456, 708, 826, 253, 645, 807, 234]. D27.25 [34]. D7.00 [51]. Daniel [692]. Darby [690]. data [247, 336, 774, 502, 765, 642, 773, 741, 87, 578, 672, 239, 113, 273, 451, 174, 109, 546, 757, 601, 573, 258, 265, 242, 563, 653, 146, 444, 473, 538, 413, 344, 846, 718, 779, 847, 692, 25, 699, 17, 164, 359, 436, 148, 655, 528, 348, 89, 422, 558, 69, 137, 277, 464, 462, 104, 299, 525, 497, 480, 564, 95, 85, 21, 96, 129, 8, 90, 666, 781, 94, 271, 777, 91, 657, 767, 654, 671, 243, 476, 193, 498, 78, 663, 349, 33, 507, 820, 660, 523, 298, 72, 640, 641, 843, 28, 804, 664, 627]. data [787, 34, 621]. database [166, 159]. databases [99, 713]. dates [818]. David [291, 364, 619]. Davis [607]. Day [723]. death [569, 111, 272, 671, 421, 358, 500, 706]. deaths [423]. decision [352, 825, 643, 321, 319, 310, 79]. decision-aids [79]. decision-making [319]. decline [547]. decline-further [547]. Dedication [52]. defects [143, 267]. Deficiencies [73]. Deficiency [289]. Definition [682, 119]. definitive [70]. Dekker [692, 621, 137]. deleted [416, 204, 263]. delusion [591]. demand [588]. dementia [794, 822]. demographic [632]. demonstrate [567]. Density [524]. Dental [509, 451, 59]. departments [588]. dependent [435]. depends [572]. deriving [131]. Descriptive [528, 63]. Design [529, 517, 683, 139, 213, 344, 431, 723, 284, 10, 104, 687, 140, 269, 684, 493, 619, 686, 645, 310, 342, 200]. designed [121, 18]. Designing [450, 567]. designs [246, 232, 638, 838, 518, 702, 559, 736, 329, 522, 424, 136, 296, 534]. detect [600]. detectable [201]. detected [88, 460]. detecting [227]. Detection [22, 660, 105, 7, 20]. determination [201, 314]. determinations [637]. Determining [77, 103]. develop [356]. developing [836]. development [305]. developments [48]. deviations [603]. devices [230]. diabetes [545, 546]. diagnose [465]. diagnosis [449, 797, 417, 62, 320, 259, 818, 224, 225]. diagnostic [407, 250, 509, 532, 551, 357, 703, 552, 800]. dialogue [319]. diastolic [511]. dichotomized [576]. dichotomous [246]. diet [765]. difference [188]. differences [394, 88]. different [100, 451, 4, 584, 271]. differential [699, 589, 328, 752]. digestive [447]. dilution [109, 528, 140]. direct [6]. disability [499]. disagree [442]. discharge [196]. discipline [387, 442]. discordance [466]. discrete [577]. discrete-time [577]. Discriminant [63, 239, 82, 78]. Discrimination [299, 527, 78]. discriminatory [526]. discussants [370, 470, 830]. Discussion [155, 158, 161, 163, 165, 167, 171, 173, 175, 178, 180, 372, 374, 376, 378, 380, 382, 384, 386, 388, 391, 393, 395, 397, 399, 401, 403, 405, 502, 478, 484, 496, 487, 755, 758, 762, 764, 766, 768, 847, 221, 104, 776]. Disease [629, 105, 672, 273, 377, 658, 135, 778, 7, 509, 130, 192, 482, 226, 58, 536, 552, 112, 849, 654, 238, 224, 335, 641]. diseases [125, 61, 261, 55, 253, 328, 660, 122, 437, 262, 680]. disorders [681]. displaying [17]. distributed [642]. distribution [300, 11, 495, 491, 362]. distribution-free [11]. distributional [433]. distributions [720, 718, 287, 481]. districts [588]. DM [365]. DNA [693]. do [373, 341, 172]. Doctor [459]. doctors [341, 790]. Doll [690]. Donald [137]. donors [300]. Dordrecht [785]. Dose [277, 194, 100, 729]. Dose-interpolation [277]. Dose-response [194, 100]. dose-toxicity [729]. doses [314]. Doubling [618]. Douglas [51]. Dowie [825]. Dr [495]. drop [574]. drop-outs [574]. dropouts [783]. Drug [596, 97, 174, 559, 815, 585, 454, 801]. drugs [73, 20, 301, 602, 314]. dual [753]. Duchenne [285]. due [361]. duration [637, 598]. durations [584]. during [475]. Düsseldorf [308]. Dynamic [189, 63, 452]. dyspepsia [79]. dystrophy [285].

each [815]. Early [463, 177, 433, 396, 37, 833, 437]. East [453]. ECOG [854, 688]. ecological [21]. Ed [706, 693]. edited [456, 364]. edition [788, 708]. Editor [623, 705, 644, 722, 233, 689, 854]. Editorial [571, 2, 468, 828, 128]. editors [583, 373]. Eds [788, 151, 150, 581, 234, 365]. Edward [691]. Effect [741, 4, 343, 490, 542, 461, 613, 453, 192, 201, 553, 589, 32, 107, 782, 312, 603, 606, 263, 499]. Effective [434]. effectiveness [850, 667]. Effects [396, 492, 638, 375, 452, 584, 130, 639, 799, 58, 104, 548, 86, 686, 591, 454, 640]. efficacy [97, 616, 432, 801]. Efficiency [819, 744, 835, 423, 415, 177, 850, 847, 832, 580, 179, 844, 849, 260, 836, 839]. Efficient [842, 848, 833]. Eighth [595]. elderly [822]. eligibility [840]. Elstein [825]. Elston [607]. Emergency [669, 105, 7, 588]. Empirical [695, 208, 305]. employed [844]. end [335]. end-stage [335]. endpoints [679, 678, 681, 682, 848, 680]. England [253]. enhance [841]. enhanced [558]. enhancing [849]. Ensuring [129]. environmental [544]. epidemic [634, 599, 289, 628, 625]. epidemics [22]. epidemiologic [295, 83, 329, 553, 118, 211, 753, 713, 103, 777, 243, 537, 752]. epidemiological [718, 72, 539, 490]. Epidemiology [54, 757, 226, 244, 44, 294, 324, 561, 363]. epilepsy [795]. equality [250]. equalized [524]. equivalence [615, 701]. Era [437]. Erasmus [67]. error [765, 756, 357, 62, 343, 754, 145, 763, 752]. Errors [751, 375, 760, 359, 759, 738, 767, 56]. Errors-in-variables [751]. Essential [707]. Essentials [607]. establish [701]. establishing [300]. Estimability [732]. estimate [616]. estimated [568, 696]. estimates [242, 466, 131, 436, 553, 589, 603, 147, 754]. Estimating [667, 506, 328, 499, 613, 509, 554, 28]. Estimation [720, 642, 429, 702, 799, 775, 298, 818, 566, 273, 196, 805, 139, 213, 27, 553, 188, 612, 643, 662, 536, 543, 525, 732, 349, 587, 208, 641, 547, 241, 510]. estimator [413, 776]. estimators [745, 436, 715]. etiologic [732]. Eugene [621]. European [706]. evaluate [121, 159]. Evaluating [526, 432, 460, 97, 187, 450, 535]. Evaluation [347, 79, 160, 156, 189, 285, 729, 552, 845, 272]. event [636, 415, 563, 580, 120]. events [447, 435]. Everitt [456]. evidence [371]. evolution [495, 491, 223]. Exact [555, 334]. Examination [590, 844]. examinations [476]. Examining [590]. example [590, 159, 598, 507, 804]. examples [773]. excess [112, 732, 544]. exclusion [30]. expected [149, 102, 197, 297, 245]. Experience [157, 174]. experiment [170, 191, 11, 18]. experimental [719]. experiments [517, 332, 733, 645]. expert [338, 225]. Explanatory [603, 492]. exploitation [305]. Exponential [742]. exposed [100]. Exposure [576, 323, 639, 760, 759, 589, 262, 752]. Exposure-response [576]. exposures [4, 505]. Extended [824, 285]. Extension [426, 54]. external [229, 69]. extraneous [111]. extrapolation [549]. extreme [511].

F [607, 708, 787, 34, 855, 269]. face [324]. facial [356]. factor [821, 495, 491]. Factorial [638, 838, 232, 200]. factors [496, 239, 472, 332, 219, 220, 190, 228, 274, 275, 564, 90, 206, 207, 262, 335, 141]. facts [459]. fallacy [164]. false [720]. Familial [486, 429]. family [265]. Fantasy [176]. Farewell [708]. fast [602]. Fatal [670]. female [631]. fetoprotein [673]. fibre [18]. fibres [600]. Fifth [217]. filtering [477]. Fischer [365]. Fisher [788, 334]. fit [147]. Fitting [359, 89, 718, 277]. fixed [524]. Fleiss [517]. Flexible [697]. Fluoride [451]. follow [247, 718, 287, 523]. follow-up [247, 718, 287, 523]. following [698, 447, 719, 271, 57]. Forecasting [588, 662]. form [740]. Formulae [855, 416, 204, 605]. formulation [491, 495]. four [678, 272]. fourfold [241]. fraction [131, 436, 589]. fractions [732]. Framingham [489, 488]. France [599]. Francis [691, 807]. Frederick [581, 235]. free [11]. Freeman [692]. French [316]. friend [41]. function [105, 472, 546, 7, 473, 431, 733, 564, 640, 547]. funds [850]. Further [767, 547].

G [51, 560]. gains [814, 744]. gait [98]. Galen [137]. game [26]. Gardner [253]. Gart [645]. Gary [854]. gastrointestinal [227, 320]. General [789, 474, 480, 495, 491]. generalized [756, 265, 718, 412]. generation [130]. Genetic [294, 536, 122]. genetics [429]. genstat [81]. geographic [654]. geographical [276]. gestation [15]. gissi [834]. given [90]. Gleser [761]. glim [301, 807, 806]. Goldstone [214]. gonorrhea [663]. Goodman [704]. Goodness [147]. Goodness-of-fit [147]. Gore [51, 785]. Gothenburg [595]. Govindarajulu [749]. Gower [269]. GPO [788]. Grading [811]. Graphical [718, 712, 568, 120, 330]. Griffin [125]. Group [789, 772, 698, 82, 854, 192, 361, 688, 245]. grouping [276]. groups [302, 538, 90]. growth [472, 360, 11, 477]. guide [786, 516]. Guidelines [788]. Gustav [365].

H [788, 692, 235, 137]. H. [39]. habitual [326]. haemophilia [818]. haemophilia-associated [818]. Haenszel [413, 821, 85, 6]. Hall [826]. Halperin [495]. Hand [456, 826]. Handbook [709]. handwashing [304]. haplotypes [290]. hard [826, 807]. hardback [291, 825]. Harris [621]. Harvard [291]. hazard [100]. hazards [819, 190, 265, 538, 695, 696, 264]. head [309, 271]. Health [485, 590, 455, 668, 835, 709, 691, 786, 579, 21, 269, 588, 59, 823, 43]. healthy [461, 263]. Healy [806]. Heart [239, 489, 377, 658, 488, 39, 112, 238, 358]. Heinemann [214]. Held [485, 67, 217, 308, 441, 595, 793]. hemodialysis [335]. hepatitis [673]. hepatocellular [228, 673]. Heterogeneity [597, 631, 27, 446, 262]. hidden [542, 499]. high [300]. high-risk [300]. Hill [305, 48, 37, 38, 44, 43]. Hinde [807]. hip [805, 543]. Historical [156, 142, 157]. historically [637]. histories [97]. history [273, 563, 778]. HIV [629, 632, 630, 731, 628, 631, 627]. HLA [61, 513]. Hodgkin [135, 318]. Holland [706]. Home [653]. homes [794]. homicide [671]. homogeneity [244, 813]. Hörmann [365]. hormonal [8]. hormone [559]. Hospital [789]. hour [548]. Howard [709]. human [429, 629, 360, 101]. hybrid [329]. hypertension [596, 783, 781, 574]. hypospadias [507, 804]. hypotheses [728]. Hypothesis [227, 435, 346, 520].

Ian [786]. IARC [645]. ICDA [19]. identification [190]. Identifying [564, 586, 784, 472]. II [364, 723, 442, 698, 411, 313, 590, 684, 207, 315]. III [581, 274, 700, 315]. illness [269]. images [811]. imbalance [685]. Immune [289]. immunoassay [277]. immunodeficiency [629, 508]. immunotherapy [11, 157]. impact [632, 177, 854, 309, 688, 446, 331]. Implications [404, 310, 840, 634]. importance [69]. important [586, 784]. impossibility [421]. improve [242, 850]. Improved [466, 132]. Improving [601, 551, 179]. imputation [653]. Inborn [610]. Inborn/outborn [610]. incidence [626, 509, 743, 482, 112, 328, 818, 627]. incidences [653]. incomplete [642, 573, 55, 498]. increasing [839]. incubation [508]. independence [617, 618, 320]. independent [465, 348, 371, 272]. index [81, 557, 347, 575]. indicants [797]. indicators [232, 81]. indirect [346]. indirectly [361]. individual [322, 818, 463, 510]. individuals [289, 21, 511]. Indoor [760, 759]. induced [799]. industrial [461]. industries [670]. inevitability [421]. Infant [668, 274, 275, 206, 207, 272, 358]. infants [144]. infarction [189, 78]. infection [632, 630, 134, 719, 731, 628, 627]. infections [54]. infectious [261]. inference [288]. influence [394, 44, 111]. influenza [599, 587]. information [569, 177, 772, 733, 746, 145]. informative [500]. Ingelfinger [235]. Ingram [364, 234]. inherent [421]. initial [572]. injuries [817, 670]. injury [653, 309, 271]. Innsbruck [793]. inosiplex [142]. insights [547]. Instrumental [552]. intake [81, 15]. integrated [129, 315]. Intelligent [319]. intended [86]. Inter [394, 191]. Inter-rater [191]. Inter-study [394]. interaction [83, 495, 491, 614]. Interactions [583, 218, 521, 264]. intercourse [631]. Interim [248, 316, 303, 514, 5]. intermediate [717]. Internal [408, 69, 839]. International [67, 217, 308, 441, 595, 793, 723, 271, 67, 217, 308, 441, 595, 793]. interobserver [805, 543, 123]. interpolation [17, 277]. Interpretation [130, 398, 27, 394, 186, 429, 621, 299, 557, 224, 379]. interpreting [283]. interval [566, 188, 598]. interval-censored [598]. intervals [336, 740, 728, 113, 434, 655, 299, 147, 556, 754]. Intervention [855, 453, 605, 414]. Introduction [751, 471, 678, 620, 806]. Introductory [831]. investigate [274, 275, 206, 207]. Investigating [760, 759]. Investigation [513, 118, 211, 725, 114, 112]. Investigations [544]. involving [535, 539]. ionizing [323]. irradiation [447]. irregular [193]. irreversibility [421]. irreversible [328]. ischemic [238]. Island [693]. Issues [423, 485, 639, 269, 479, 150, 752]. Italian [459, 834]. IV [459, 275]. ix [806].

J [291, 788, 806, 456, 125, 561, 790, 151, 826, 253, 619, 645, 807, 517]. Jack [825]. James [235]. Japan [437, 818]. Jay [761]. Jenkins [261]. Jerald [34]. Jewell [745]. John [581]. Johnson [607]. joints [805, 543]. Joseph [235]. journal [583]. journals [9]. Jr [692, 137]. judgement [825]. June [485].

Kalman [477]. kappa [123]. kappa-type [123]. Karger [708, 749]. Kaunas [414]. Kenneth [561, 150]. Kensington [790]. kinetics [734]. Kirkwood [707]. Kiryluk [690]. Kish [518]. Kluwer [785]. know [341]. Knowledge [225]. known [141]. Konferenz [365]. Krewski [645]. KRIS [414]. Kruskal [704].

L [517, 529, 235, 137]. laboratories [845, 71]. laboratory [621, 298, 72]. lag [652]. large [387, 375, 436, 602, 849, 172]. Latent [578, 226, 242, 417, 144, 259]. law [68]. Lawless [34]. Lectures [693]. Lee [645]. left [464]. Lehmacher [365]. Leicestershire [82]. Leon [761]. Leonard [214]. Lesley [691]. Leslie [518]. Lessons [383]. Letter [705, 644, 854, 722]. Letters [233, 689]. leukaemia [719, 45]. leukemia [221, 316]. levels [639, 84]. LF67 [789]. Life [317, 693, 739, 569, 108, 37, 464, 416, 101, 204]. Life-table [317]. lifetime [34]. ligand [734]. ligand-binding [734]. Light [291]. Likelihood [482, 745, 821, 31, 60, 212, 349]. limitations [392]. limits [77, 779]. line [17]. Linear [232, 739, 740, 756, 699, 59, 498]. linkage [672, 709, 122]. List [370, 470, 830]. litter [95, 94]. litter-matched [95, 94]. Little [561]. liver [546, 564]. lives [667]. Living [516]. locations [669, 654]. Log [59, 699, 241]. Log-linear [59, 699]. logical [449]. Logistic [105, 819, 794, 239, 7, 526, 210, 727, 695, 579, 527, 118, 211, 227, 240, 823, 754, 60, 212, 78, 241]. logit [146, 274, 275, 206, 207]. logrank [741, 16, 266, 412, 424, 94]. London [51, 125, 691, 790, 826, 214]. Long [269, 14, 316, 323, 17, 645, 511]. long-term [316, 323, 17, 645, 511]. Longitudinal [485, 547, 819, 473, 497, 480, 475, 285, 427, 243, 476, 820, 640]. look [704]. lose [371]. Loss [520, 177, 31]. low [639]. lower [814]. Ltd [269, 214]. lung [472, 220, 10, 733, 547, 76, 490]. lymphocytic [221, 316]. lymphoma [318].

M [806, 51, 690, 691, 151, 785, 137, 693, 253, 807, 151]. Machin [619]. Macmillan [305, 235]. made [854, 688]. Magnitude [31]. major [437]. making [352, 321, 319, 825]. malaria [453, 125]. Malaysia [453]. male [631]. male-to-female [631]. malformations [327]. mammographic [4]. management [691, 129, 311]. Mantel [413, 821, 85, 6]. map [524]. Marcel [692, 621, 137]. marginal [481]. Marino [217]. Mark [691]. markers [80]. Markov [731, 557]. Martin [620]. Mass [291, 581]. Masthead [1, 13, 24, 36, 53, 65, 93, 115, 126, 138, 152, 184, 199, 215, 237, 255, 270, 282, 293, 306, 325, 339, 351, 368, 406, 419, 430, 439, 457, 467, 504, 519, 531, 541, 550, 562, 570, 582, 593, 609, 622, 635, 649, 676, 694, 710, 724, 737, 750, 770, 791, 809, 827]. matched [345, 730, 555, 95, 94, 240, 744, 260]. matched-pair [730]. Maternal [668, 15]. maternity [428]. mathematical [693, 652, 364, 234]. Mathematics [693]. matrices [498]. Matthews [708]. Maximum [349, 745, 821, 60, 212]. McDowell [786]. mean [511]. meaningful [381]. means [842, 245]. measure [489, 186, 488, 287, 527, 704, 802]. measured [576]. measurement [765, 756, 19, 343, 767, 754, 763, 752]. measurement-error [763]. measurements [478, 337, 462, 481]. measures [578, 114, 136, 474, 826]. Measuring [108, 738, 8, 230, 786]. mechanisms [573]. medelling [823]. Medical [789, 51, 683, 142, 247, 9, 560, 578, 708, 449, 352, 417, 643, 136, 99, 129, 707, 669, 179, 150, 620, 259, 214, 225, 356, 379, 581]. medically [381]. medicine [77, 364, 68, 234, 321, 854, 855, 235, 38]. Medico [789]. Medico-Pharmaceutical [789]. Meeting [67, 217, 308, 441, 595, 793]. Meinert [529]. melanoma [219]. memory [659]. Meta [387, 442, 610, 342]. Meta-analysis [387, 442, 610, 342]. metastases [564]. Method [103, 797, 426, 805, 295, 17, 424, 543, 447, 729, 602, 524, 72, 641]. methodological [445, 314]. Methodology [206, 366, 455, 585]. Methods [478, 392, 151, 572, 110, 150, 149, 560, 302, 636, 626, 109, 757, 709, 364, 473, 787, 317, 723, 327, 34, 131, 102, 197, 234, 202, 662, 480, 551, 70, 474, 26, 321, 844, 272, 78, 537, 262, 28, 625]. Michael [619]. Michalek [95]. Miettinen [363]. might [332]. Mihalko [95]. Mike [150]. mild [596]. minimal [22]. minor [334]. Misclassification [821, 553, 730, 589]. misclassified [757]. mismeasuring [492]. mismodelling [492]. Missing [497, 502, 642, 653, 349, 481]. misuse [803]. Miura [693]. Mixture [781]. mixtures [747]. modalities [720]. model [739, 725, 794, 642, 740, 87, 632, 239, 429, 190, 258, 265, 653, 699, 526, 210, 333, 535, 448, 54, 731, 508, 652, 227, 240, 59, 243, 264, 763, 640, 747]. model-based [653, 508]. Modelling [546, 453, 567, 599, 422, 634, 132, 261, 360, 579, 476, 481, 807]. Models [461, 410, 411, 21, 491, 203, 280, 628, 774, 545, 819, 773, 578, 232, 229, 756, 613, 512, 577, 538, 697, 629, 34, 359, 655, 412, 89, 695, 482, 719, 274, 275, 781, 319, 206, 207, 91, 29, 493, 322, 498, 696, 820, 633, 335, 495]. moderate [375]. Modern [561]. Modified [355, 373]. Monitoring [433, 428, 666, 415, 444, 580, 299, 71, 625]. Monte [143, 267, 346, 433, 133]. morbidity [131, 557, 641, 499]. Mortality [82, 545, 149, 100, 229, 102, 197, 222, 482, 662, 297, 26, 274, 275, 746, 495, 491, 111, 414, 206, 207, 557, 29, 253, 328, 347, 117, 101, 587, 45, 437, 208, 641, 263, 499]. Mosteller [581, 235]. mother [428]. mothers [332]. motivated [822]. much [174, 107, 782, 606, 843]. multi [82, 482]. multi-group [82]. multi-state [482]. Multicentre [48]. multifactorial [575]. Multinomial [512]. Multiple [426, 532, 526, 19, 776, 205, 753, 586, 347, 784, 587, 721]. Multiplicative [229]. Multistate [545]. Multivariate [801, 63, 201, 85, 96, 498, 621, 826]. Munro [790]. Muscatine [475]. muscle [600, 18]. muscular [285]. myocardial [189, 78]. myocarditis [852].

N [690, 723, 645]. narrow [840]. National [590, 668, 69, 671, 670]. natural [273, 778]. need [375, 832, 86, 373, 172]. negative [720, 283]. neighbourhood [590]. Nejm [581]. Neonatal [26, 274, 275, 610]. nephropathy [545]. neuronal [196]. neurotransmitter [559]. Newcombe [709]. Newell [786]. Ninth [793]. NO [125, 291, 789, 788, 806, 51, 560, 456, 708, 518, 706, 529, 690, 305, 709, 691, 364, 787, 561, 786, 790, 692, 723, 34, 363, 785, 825, 235, 137, 693, 826, 707, 269, 253, 749, 645, 214, 807, 365, 581, 607, 234, 621, 150, 620, 619]. Non [429, 168, 614, 336, 642, 638, 573, 589, 104, 318, 610, 31, 120, 752]. non-compliance [638]. non-differential [589, 752]. non-Hodgkin [318]. non-normally [642]. Non-parametric [614, 336, 104, 120]. Non-random [429, 573]. Non-randomized [168, 610]. non-randomly [642]. non-responder [120]. non-stratified [31]. nonlinear [359]. normal [77, 298]. normally [642]. Norman [125]. Note [188, 623, 779, 568, 521, 193, 820, 251]. notes [790, 789]. notifiable [660]. NSW [788]. Number [517, 516, 630, 16, 266, 513, 366, 5, 260]. numerical [356]. nurses [579, 823]. nutrient [15]. Nutrition [590].

objective [506]. Observation [170, 463]. observational [160, 164, 166]. observations [535, 598, 481]. observed [312]. obstetric [565, 428]. obtained [109]. Obtaining [381]. occupational [672, 639, 669, 43, 670]. occur [603]. occurrence [415, 580, 120, 660]. odds [745, 87, 413, 702, 568, 554, 813, 241]. off [800]. old [794, 704]. Olli [363]. omitting [613]. omnimetrics [164]. Oncological [788]. oncology [313, 701]. one [821, 618, 701]. one-sided [618, 701]. open [142]. Opening [650]. operational [682]. operative [387]. Ophthalmologic [681]. opinion [257]. Opportunities [849, 385]. Optimal [246, 604, 534, 251, 559]. optimistic [520]. oral [373]. ordered [774, 794, 726, 245]. ordinal [258, 146, 699, 348, 763]. ordinating [141]. organ [124]. organs [539]. orthodontists [191]. osteoarthrosis [805, 543]. other [730, 299, 78]. outborn [610]. outcome [564, 271, 575]. outcomes [683, 684, 479]. outs [574]. outstanding [563]. overview [377, 832, 752, 381]. overviews [375, 383, 398, 396, 404, 373, 402, 379]. Oxford [806, 560, 706, 529, 690, 709, 786, 723, 151, 707, 620, 619, 645, 807].

P [560, 253, 645]. package [601]. paediatric [508, 701]. pages [291, 581, 517, 789, 788, 806, 51, 560, 607, 456, 708, 518, 706, 529, 690, 125, 305, 709, 691, 364, 787, 561, 786, 790, 692, 723, 34, 363, 234, 621, 785, 825, 235, 516, 137, 693, 826, 707, 269, 150, 253, 620, 749, 619, 645, 214, 807, 365]. pain [258, 146, 465, 356]. pair [345, 730]. pair-matched [345]. paired [573, 702, 11, 539]. paired-cluster [702]. Palazzo [217]. palsy [98]. pancreatitis [223]. panencephalitis [142]. paper [826]. paperback [291, 825]. papers [847, 678]. Parameters [101, 569, 642, 536, 556, 747]. parametric [336, 104, 614, 120]. Part [364, 234]. partially [781]. participation [455]. particular [284, 62]. Partitioning [321, 146]. partners [630]. passive [523]. pathology [506]. Patient [316, 176, 25, 443, 166, 299, 311]. patients [105, 7, 30, 16, 266, 88, 795, 575, 335, 640, 463, 141, 840]. pattern [289, 238]. patterns [100, 162, 98]. Peculiarity [437]. pedigrees [205]. penetrance [55]. penile [631]. penile-vaginal [631]. people [794]. percentage [771]. performance [450, 810, 760, 759, 532, 551, 309, 71, 78]. Perinatal [82]. period [410, 411, 810, 736, 522, 422, 687, 508, 285, 29, 322, 437]. permutation [96]. person [754]. personal [658, 40]. persons [98]. perspective [846, 179, 849, 848]. perturbations [334]. Peto [690]. Pharmaceutical [789, 258, 146]. phase [698, 313, 700, 315]. Philadelphia [607]. physician [835]. physicians [200]. Pike [690]. Pillemer [291]. pilot [839]. placebo [604, 687, 107, 782, 606]. placental [15]. Planning [176, 533]. Plasma [69, 795]. pneumoconiosis [123]. pneumonia [587]. point [196, 800, 524, 507, 804]. points [465]. Poisson [657, 721]. police [671]. Policy [634]. Polio [833]. pollution [760, 759]. polychotomous [227]. polynomial [569]. pooling [385]. population [353, 361, 222, 147, 438]. populations [100, 549, 602, 101]. possibility [133]. possibly [481]. Post [275]. Post-neonatal [275]. postal [332, 428]. postmarketing [585]. Potential [385]. Power [201, 412, 814, 741, 520, 345, 637, 290, 526, 83, 114, 357, 505, 119, 260]. Practical [444, 826, 56]. practically [314]. practice [151, 852, 834, 851, 51, 269]. pragmatic [603]. pre [338]. pre-assessment [338]. preclinical [273]. predict [564]. Predicted [100, 4]. Predicting [452, 795]. Prediction [113, 272, 575, 132, 311]. predictive [250, 249, 257, 330]. Preface [469, 624, 66, 216, 307, 340, 440, 458, 594, 677, 711, 792, 153, 829, 369]. preference [686]. preferences [121]. preliminary [715]. preliminary-test [715]. prenatal [54]. preoperative [450]. preparations [194]. Prescription [602]. presence [130, 61, 55, 486, 347, 614, 500]. present [72]. presentation [568, 373]. Press [291, 806, 456, 706, 529, 690, 305, 709, 786, 723, 151, 825, 620, 645, 807]. Press/IARC [645]. pressure [230, 475, 548, 84, 511]. prevalence [794, 775, 328]. prevented [589]. prevention [567, 629, 312, 493, 849]. Price [291, 789, 788, 806, 51, 560, 456, 708, 518, 706, 690, 125, 305, 709, 691, 364, 787, 561, 786, 790, 692, 723, 34, 363, 785, 825, 235, 516, 137, 693, 826, 707, 269, 253, 749, 645, 214, 807, 581, 517, 607, 529, 151, 234, 621, 150, 620, 619]. primary [464, 673]. principle [70]. Printing [789]. Pritchard [305]. Probabilistic [311]. probabilities [336, 434]. probability [449, 630, 62, 421, 631, 511]. probability-based [449]. probands [205]. problem [352, 662]. Problems [379, 445, 563, 612, 544, 335]. procedure [97, 790]. procedures [250, 177, 284, 532, 772, 552]. process [577, 495, 491, 318, 557]. Professional [825]. progesterone [69]. prognosis [189]. Prognostic [219, 220, 90, 335, 81, 221, 190, 132, 222, 564, 309]. programmes [720, 452]. progress [563]. progressive [509]. progressively [32]. project [174]. projections [626, 524]. prolonged [463]. proper [703]. proportion [698]. Proportional [538, 819, 87, 190, 265, 695, 111, 117, 696, 264, 263]. proportionate [347]. proportions [423, 604, 188, 251]. proposal [283]. proposed [544]. prospect [46]. Prospective [539, 295, 433, 210, 118, 211]. prospects [338]. protective [616]. protocol [789, 450, 316, 159]. protocols [304]. Providence [693]. provisional [662]. psoriasis [513]. psoriatic [513]. Psychiatric [286, 551]. psychiatry [794]. publication [342]. Publications [560, 707, 645, 619]. published [186]. Publishing [235, 269]. pulmonary [473, 760, 759, 640]. putative [544]. PVC [97].

quality [601, 108, 366, 69, 162, 129, 666, 844, 408]. quantitative [613, 357, 800]. quantity [569]. quartiles [779]. questioning [3]. questionnaire [332, 428]. questionnaires [786]. questions [842, 693]. Quetelet [81].

R [788, 806, 690, 151, 234, 707, 645]. radiation [323, 4, 639, 162]. radiographs [123]. radiological [805, 543]. Ralph [364]. random [429, 573, 375, 535, 685, 640]. Randomization [855, 355, 47, 344, 702, 605, 32, 86, 31]. randomized [337, 154, 218, 220, 316, 601, 392, 168, 398, 404, 448, 343, 610, 373, 200, 843, 172, 381, 840]. randomly [642]. range [795]. ranges [506, 298]. rank [203, 280]. ranks [802]. rare [415, 580]. rate [566, 412]. rater [191]. rates [720, 346, 636, 410, 411, 604, 509, 130, 110, 209, 29, 427, 798, 358, 721, 208, 251]. rating [786]. ratio [745, 426, 413, 702, 131, 103, 349]. rationale [655]. ratios [512, 466, 568, 27, 329, 554, 111, 813, 117, 241]. raying [565]. reaction [585]. reactions [20, 602, 454]. reader [825]. reading [333, 532]. reality [176]. really [107, 782, 606]. reanalysis [713]. receive [551]. receptor [559]. reciprocal [838]. record [709]. recordings [548]. records [428, 322, 671]. recruitment [176, 846]. recurrence [567, 58, 798]. recurrent [435]. reduce [596]. Reducing [56, 849]. reference [77, 361, 222, 70, 506, 124]. reflections [49, 75]. region [588]. register [454]. registers [142]. registration [73]. registries [164]. registry [507, 523, 804]. Regression [97, 265, 132, 726, 719, 739, 725, 819, 794, 642, 239, 190, 512, 353, 426, 697, 526, 148, 727, 89, 695, 107, 782, 240, 657, 754, 60, 212, 798, 696, 438, 606, 264, 763, 241, 137]. Rejoinder [761]. relapse [812]. relapsing [223]. related [346, 335, 752]. relating [262]. relationship [576, 120]. relationships [729]. Relative [55, 740, 745, 415, 242, 426, 712, 580, 554, 95, 493, 754]. Reliability [25, 242, 409, 431, 509]. Remarks [651, 674, 503, 853, 831, 650, 675]. reminiscence [40]. Remission [598]. renal [781, 335, 785]. renin [781]. Repeated [484, 489, 774, 478, 578, 337, 488, 114, 136, 474, 826, 57, 479, 481]. replacement [355]. Replicate [442]. reply [716]. Report [485]. reported [743]. Reporting [443, 283, 652, 654]. reports [73, 753, 428, 585, 366]. representation [569, 120]. Required [637, 16, 266]. requirements [337, 409, 554]. Research [790, 518, 317, 723, 166, 136, 129, 116, 150, 291, 560, 691, 269]. resected [220]. residents [822]. residuals [718]. resistance [134]. respondent [822]. responder [120]. response [774, 63, 484, 100, 302, 332, 109, 353, 604, 577, 726, 576, 492, 669, 767, 802, 438, 534, 194, 495]. responses [246]. resulting [31]. results [394, 295, 398, 284, 90, 852, 586, 767, 784, 851]. retest [242]. Retrospect [46]. retrospective [162, 746, 28]. Review [291, 581, 517, 789, 788, 806, 51, 560, 607, 456, 708, 518, 706, 529, 690, 125, 305, 709, 691, 364, 787, 561, 786, 790, 692, 723, 34, 363, 151, 234, 621, 785, 825, 235, 516, 137, 693, 826, 548, 707, 269, 150, 253, 620, 749, 619, 645, 214, 807, 365, 219, 757, 563, 444, 139, 213, 108, 83, 202, 408, 294]. reviewing [291]. reviews [824]. Rhode [693]. rhythm [359]. Richard [291]. ridge [148]. right [779, 500]. Rigoni [459]. Rigoni-Stern [459]. Rising [794]. Risk [496, 740, 745, 565, 239, 658, 242, 426, 549, 228, 329, 596, 516, 300, 525, 21, 495, 491, 90, 493, 754, 575, 28]. risks [712, 536, 554, 55, 544]. Robert [607, 787, 693]. Robustness [348]. ROC [780]. role [47, 160, 326, 402, 839]. Room [789]. Rosner [854]. Rothman [561]. Rotterdam [67, 414]. routine [231, 71]. Roy [305]. rubella [452, 54]. rule [822, 703]. rules [74, 396, 612]. run [841]. run-in [841]. runs [558]. rural [59].

S [788, 456, 690, 363]. S09 [789]. same [451, 584]. Sample [337, 728, 345, 409, 605, 727, 554, 505, 315, 346, 566, 698, 139, 213, 434, 201, 424, 521, 357, 414, 616, 260, 241, 855]. samples [348]. sampling [429, 82, 822, 103, 251]. San [217, 26]. sarcoidosis [640]. savage [32]. saving [667]. scale [849]. scaled [348]. scales [786]. scan [224, 362]. scheme [659]. schemes [69, 205]. School [789]. science [291]. Sciences [693]. Scientific [560, 707, 645, 387, 442, 619]. Scientists [459, 826]. sclerosing [142]. scores [94]. scoring [703, 272]. Scottish [144]. screen [88]. screen-detected [88]. Screening [673, 720, 273, 797, 4, 846, 629, 192, 775, 460]. seat [667]. second [799]. sectional [547]. selected [353, 253, 438]. Selecting [239, 257]. Selection [320, 840, 59, 686, 510]. selective [450]. self [686, 591]. self-delusion [591]. self-selection [686]. Sense [785]. Sensitivity [334, 330, 617, 776]. sensitization [785]. September [67, 217, 308, 441, 595, 793, 459]. sequence [693, 602]. sequences [196, 15]. Sequential [284, 302, 698, 854, 10, 772, 688, 76]. serial [109, 577, 528, 114, 535]. series [678, 188, 662, 8, 817, 663, 587, 331, 664]. serious [177]. seroconversion [719]. Serriffe [26]. services [428]. sets [415, 580, 744]. Seventh [441]. several [423, 295, 568, 262]. severe [309, 271]. Severity [19]. sex [454]. sexual [631]. SFr [137]. Shanghai [822]. shared [290]. shared-haplotypes [290]. Sheila [51, 785]. Shewhart [231, 70]. short [626, 189, 659]. short-term [626, 189]. should [394, 174, 843]. show [100]. Showa [437]. Sickle [837]. sided [618, 701]. signed [203, 280]. significance [278]. significantly [177]. similarity [802]. simple [17, 424, 172]. Simplified [96]. simulation [143, 267, 302, 278, 567, 433, 118, 211]. Simultaneous [336]. Single [822, 745, 140]. Single-stage [822]. six [377]. Sixth [308]. Size [855, 741, 337, 566, 345, 139, 213, 409, 201, 424, 605, 727, 521, 554, 357, 505, 18, 533, 147, 257, 260]. sizes [728, 434, 616, 315]. skin [219]. Small [346, 241]. smallest [201]. Smoking [690, 746, 490]. smoothed [558]. Smoothing [537]. SMRS [100]. social [274, 275, 206, 207]. Society [67, 217, 308, 441, 595, 788, 793, 693]. soft [807]. software [365, 365]. sojourn [720]. Some [445, 332, 465, 323, 736, 522, 49, 104, 480, 90, 427, 633, 314, 228, 48, 261, 844, 78, 260, 172, 693]. source [451]. sources [544]. Southampton [789]. space [135]. spaced [476]. spans [101]. sparse [436]. speakers [370, 470, 830]. Special [487, 124]. specialized [245]. specialties [364]. specific [328]. specificity [617, 776, 330]. splines [697]. spontaneous [326, 244]. sporadics [55]. spread [632, 633, 627]. square [146, 413, 349]. stability [725, 735]. Stablein [137]. stage [221, 810, 822, 686, 534, 335, 640]. stagei [219]. stages [731]. standard [551, 62]. standardization [6]. standardized [346, 131, 117]. standards [305]. Stanley [150]. Staquet [151]. state [448, 482, 552]. States [628]. statistic [413, 362]. Statistical [626, 734, 459, 485, 234, 304, 731, 116, 358, 544, 262, 28, 98, 625, 143, 267, 636, 97, 278, 456, 451, 316, 709, 323, 290, 289, 655, 528, 743, 617, 299, 548, 615, 107, 782, 319, 585, 586, 849, 784, 606, 310, 294, 780, 335, 560, 518, 787, 723, 34, 749, 619, 807]. statistician [3, 42]. statisticians [583]. Statistics [9, 662, 6, 341, 854, 855, 581, 708, 690, 707, 620, 214, 51, 691, 38, 150]. Statistik [365, 365]. Statistik-software [365, 365]. status [448, 610]. steering [601]. Stern [459]. Stillbirths [207]. stochastic [495, 491, 633]. Stopping [74, 612, 177, 433, 396]. strata [436]. Strategies [21, 736, 522, 132]. strategy [187]. stratification [533]. Stratified [85, 344, 31]. stratum [264]. stratum-covariate [264]. streamlined [852, 851]. Street [790, 116]. Strengths [392]. strictly [509, 703]. structured [498]. students [449]. Studies [485, 855, 819, 273, 442, 778, 709, 345, 323, 168, 409, 295, 431, 637, 160, 466, 156, 83, 201, 329, 730, 210, 283, 605, 497, 480, 554, 371, 118, 211, 505, 611, 753, 240, 159, 103, 296, 777, 748, 610, 427, 845, 767, 157, 493, 744, 119, 243, 56, 421, 700, 124, 245, 262, 539, 490, 752, 839, 288, 723]. Study [105, 545, 394, 346, 100, 302, 835, 229, 450, 316, 613, 135, 258, 425, 82, 683, 146, 488, 7, 473, 821, 228, 133, 83, 361, 148, 222, 555, 104, 447, 162, 781, 223, 86, 684, 18, 414, 59, 657, 454, 260, 72, 331, 463, 239, 82, 475, 489]. studying [842]. Stuttgart [365]. subacute [142]. subadditivity [638]. Subgroup [459]. subjective [257]. subjects [551]. subsequent [48, 358]. subset [398]. sudden [272, 358]. sugar [81]. suitable [90]. Sullivan [557]. Summary [674, 525, 27]. Summing [291]. superiority [815]. surface [673]. Surgeon [459]. surgery [168]. Surgical [684]. Surrogate [679, 681, 682, 680, 678]. Surveillance [672, 668, 143, 267, 327, 655, 585, 657, 654, 663, 659, 670, 660, 664]. survey [774, 590, 579, 822, 455, 26, 823, 455, 668]. Survival [778, 597, 336, 545, 496, 741, 87, 113, 613, 265, 742, 434, 637, 88, 89, 58, 137, 464, 287, 96, 90, 533, 91, 610, 14, 446, 120, 354, 523, 335, 460]. survivors [464]. Survivorship [133, 569]. Switzerland [708]. Sydney [788]. Sylvester [151]. Symposium [366]. symptomatic [187]. syndrome [508, 272, 358, 289]. syphilis [148, 664]. system [129, 309, 659, 79]. systematic [375, 373, 754]. systems [666, 670, 225].

T [708, 125]. table [569, 317, 536]. Tables [16, 266, 511, 745, 334, 727, 618, 416, 813, 619, 349, 204, 241]. taking [361, 447]. Tarone [645]. Taylor [826, 253]. Teacher [41]. Teaching [455, 278, 449, 26, 408]. technical [193]. technique [231]. techniques [489, 143, 267, 135, 415, 488, 580, 234, 749, 71]. Technology [595]. telescopic [822]. temporal [410, 411, 799, 486]. term [626, 316, 323, 17, 189, 14, 645, 511]. termination [463]. Test [735, 242, 195, 252, 279, 334, 290, 16, 266, 715, 289, 114, 348, 600, 424, 558, 11, 551, 32, 244, 96, 552, 800, 349, 544]. test-retest [242]. Testing [507, 804, 747, 728, 429, 521, 618, 227, 435, 686]. Tests [83, 346, 407, 741, 250, 472, 450, 854, 192, 655, 412, 532, 492, 775, 357, 703, 776, 564, 203, 280, 586, 813, 688, 784, 245]. their [492, 428]. Theoretical [363]. theory [310]. therapeutic [121, 20]. therapeutics [402]. therapies [156]. therapy [445, 683, 160, 584, 162, 684, 159, 432]. Thibodeau [235].

Third [67, 15]. three [230, 302, 672, 221, 135, 473, 352, 736, 522, 422, 845]. three-decision [352]. three-period [522, 422]. three-stage [221]. Time [817, 664, 720, 542, 135, 577, 130, 662, 772, 719, 652, 8, 94, 533, 435, 767, 663, 587, 72, 331]. time-dependent [435]. time-series [331]. times [739, 637, 464, 134, 446]. titration [425, 700]. tomography [331]. topics [487]. total [361, 414]. toxicity [729]. toxicology [124]. Transcript [373]. transfusion [300, 508]. transfusion-associated [300, 508]. transient [448]. transmission [453, 631]. transplantation [785, 57]. trauma [19]. Treatment [202, 302, 218, 187, 613, 375, 353, 810, 284, 443, 596, 366, 799, 58, 134, 285, 548, 140, 603, 852, 686, 614, 438, 310, 141]. treatment-covariate [218, 614]. treatment-induced [799]. treatments [699, 736, 164, 522, 10, 687, 701, 76, 812]. trees [742, 696]. trend [195, 252, 279]. trends [130, 667, 486, 437, 72, 641]. trent [588]. trial [789, 278, 337, 154, 174, 698, 187, 220, 303, 514, 30, 567, 790, 837, 433, 142, 448, 10, 422, 285, 301, 833, 852, 701, 686, 811, 257, 200, 843, 76]. trials [338, 788, 47, 121, 186, 636, 248, 246, 638, 835, 423, 218, 377, 387, 442, 529, 316, 601, 842, 177, 392, 850, 46, 139, 213, 847, 679, 854, 108, 48, 16, 74, 266, 810, 832, 383, 568, 17, 398, 396, 284, 151, 404, 313, 678, 443, 681, 49, 535, 366, 612, 202, 58, 462, 783, 729, 5, 179, 312, 343, 533, 574, 844, 852, 373, 493, 849, 682, 688, 75, 591, 685, 796, 700, 310, 311, 354, 848, 534, 836, 463, 315, 812, 851, 839, 680, 172, 381, 840, 619]. trials-with [284]. Tromsø [239]. True [312, 552]. truncation [464]. Trust [789]. Tsiatis [854]. tumour [11, 94, 798, 80]. twin [466, 730]. Two [773, 451, 577, 720, 149, 143, 267, 302, 146, 323, 604, 810, 736, 522, 102, 197, 348, 10, 687, 304, 733, 618, 301, 209, 116, 686, 670, 534, 640, 76, 812]. two-period [810, 687]. Two-process [577]. two-sided [618]. two-stage [810, 686, 534, 640]. two-treatment [810]. two-way [116]. Tying [166]. type [195, 252, 279, 123, 357]. types [100, 741]. typology [356].

U.K [560, 305]. U.K. [690]. U.S. [835, 200, 670]. ulcer [567, 812]. unbalanced [32]. Unbiased [58]. Uncertainties [277]. uncertainty [527]. unclassified [781]. uncomplicated [223]. underlying [576]. Understanding [803, 708, 214]. unequal [741, 424, 354]. unequal-sample-size [424]. unequally [476]. Unidirectional [721]. Unit [855, 344, 605]. United [628]. units [654]. univariate [114]. University [291, 67, 308, 441, 595, 806, 456, 706, 529, 690, 709, 786, 723, 151, 825, 620, 645, 807]. unknown [552, 145]. unmatched [525]. upper [320]. Use [18, 271, 354, 720, 472, 465, 715, 210, 678, 166, 424, 300, 841, 579, 19, 274, 275, 285, 99, 301, 203, 280, 753, 94, 206, 207, 493, 823, 322, 243, 796, 33, 820, 848]. useful [314]. uses [581, 654]. Using [773, 242, 850, 164, 159, 454, 569, 740, 239, 81, 16, 266, 433, 662, 464, 673, 590, 731, 564, 90, 320, 144, 193, 498, 818, 708]. utility [249, 671]. Uxbridge [790].

V [708]. VA [845]. vaccination [452, 45]. vaccine [616]. vaginal [631].

Valerie [150]. validating [421]. validation [192]. Validity [114, 746, 140]. value [803, 572, 417, 618, 259, 245, 627]. values [502, 642, 814, 250, 330]. variability [232, 442, 333]. variable [578, 576, 492, 15]. variables [578, 751, 492, 552, 320, 717]. Variance [436, 553, 643, 413, 826, 124, 510]. variances [196]. variation [735, 410, 411, 805, 543, 822, 123]. varying [101]. venous [781]. ventilatory [448]. Verlag [365]. versus [170, 555, 772, 312, 120, 547]. vi [709, 269]. via [190, 499]. victims [358]. view [475]. viewpoint [443]. viral [261]. Virginia [669]. virus [629, 719]. vitamin [836]. vivo [299]. vole [45]. Volume [723, 693]. voluntary [629]. vs [248].

W [706]. W8 [790]. Wahrendorf [645]. Wald [690]. Wales [253]. walk [535]. Walter [137, 365]. Waltham [581]. Wampler [137]. Ware [235]. way [116]. weight [15]. weights [124]. Wessex [789]. West [485, 669]. where [245, 72]. whether [572]. which [720, 459]. while [596]. wide [840]. widely [101]. Wilcoxon [741, 195, 252, 279, 412]. Wilcoxon-type [195, 252, 279]. Wiley [517, 518, 364, 787, 34, 363, 234, 516, 150, 253]. William [607]. win [371]. Winter [253]. wissenschaftliche [365]. withdrawal [500]. within [398, 754, 510]. within-individual [510]. within-person [754]. without [177]. women [81]. Woodward [691]. Woolson [787]. work [95]. worker [461, 263]. Workshop [485, 751]. worth [596]. worth-while [596]. would [854, 688].

x [291, 693, 214, 565]. X-raying [565]. xi [692, 619, 807]. xii [607, 708, 518, 125, 723]. xiii [825, 620]. xiv [517, 786, 826, 749]. xv [788, 305]. xxii [785]. xxiii [706, 235]. xxvi [529]. xxx [690].

years [575]. York [518, 34, 363, 621, 235, 137, 150]. Young [39].

Z [749]. Zelen [140]. zones [276].

References

Anonymous:1982:Ma

[1] Anonymous. Masthead. Statistics in Medicine, 1(1):fmi, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1982:E

[2] Theodore Colton, Laurence Freedman, and Tony Johnson. Editorial. Statistics in Medicine, 1(1):1–3, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finney:1982:QS

[3] D. J. Finney. The questioning statistician. Statistics in Medicine, 1(1): 5–13, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dubin:1982:EDM

[4] Neil Dubin. Effect of different mammographic radiation exposures on predicted benefits of screening for breast cancer. Statistics in Medicine, 1 (1):15–24, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McPherson:1982:CNI

[5] Klim McPherson. On choosing the number of interim analyses in clinical trials. Statistics in Medicine, 1(1):25–36, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woolson:1982:MHS

[6] Robert F. Woolson and Judy A. Bean. Mantel–Haenszel statistics and direct standardization. Statistics in Medicine, 1(1):37–39, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1982:LFA

[7] Ralph B. D’Agostino and Michael W. Pozen. The logistic function as an aid in the detection of acute coronary disease in emergency patients (a case study). Statistics in Medicine, 1(1):41–48, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marks:1982:MCB

[8] Ronald G. Marks. Measuring the correlation between time series of hormonal data. Statistics in Medicine, 1(1):49–57, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1982:SMJ

[9] Douglas G. Altman. Statistics in medical journals. Statistics in Medicine, 1(1):59–71, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1982:DSC

[10] D. R. Jones, C. E. Newman, and J. Whitehead. The design of a sequential clinical trial for the comparison of two lung cancer treatments. Statistics in Medicine, 1(1):73–82, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koziol:1982:DFT

[11] James A. Koziol and Donna A. Maxwell. A distribution-free test for paired growth curve analyses with application to an animal tumour immunotherapy experiment. Statistics in Medicine, 1(1):83–89, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Aa

[12] Anonymous. Announcement. Statistics in Medicine, 1(1):91, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Mb

[13] Anonymous. Masthead. Statistics in Medicine, 1(2):fmi, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1982:LTS

[14] Stuart J. Pocock, Sheila M. Gore, and Gillian R. Kerr. Long term survival analysis: The curability of breast cancer. Statistics in Medicine, 1(2):93– 104, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Susser:1982:TVA

[15] Mervyn Susser and Zena Stein. Third variable analysis: Application to causal sequences among nutrient intake, maternal weight, birthweight, placental weight, and gestation. Statistics in Medicine, 1(2):105–120, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1982:TNP

[16] L. S. Freedman. Tables of the number of patients required in clinical trials using the logrank test. Statistics in Medicine, 1(2):121–129, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gans:1982:SMB

[17] Daniel J. Gans. A simple method based on broken-line interpolation for displaying data from long-term clinical trials. Statistics in Medicine, 1 (2):131–137, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morton:1982:UDE

[18] R. Hugh Morton. Use of a designed experiment for the study of muscle fibre size. Statistics in Medicine, 1(2):139–143, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levy:1982:SMM

[19] Paul S. Levy, Jack Goldberg, Siu Hui, Janet Rothrock, Norman Iverson, and Trina Hosmer. Severity measurement in multiple trauma by use of ICDA conditions. Statistics in Medicine, 1(2):145–152, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finney:1982:DAR

[20] D. J. Finney. The detection of adverse reactions to therapeutic drugs. Statistics in Medicine, 1(2):153–161, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manton:1982:SAE

[21] Kenneth G. Manton, Eric Stallard, and Wilson Riggan. Strategies for analysing ecological health data: Models of the biological risk of individuals. Statistics in Medicine, 1(2):163–181, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knox:1982:DME

[22] E. G. Knox and R. Lancashire. Detection of minimal epidemics. Statistics in Medicine, 1(2):183–189, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Ab

[23] Anonymous. Apology. Statistics in Medicine, 1(2):i, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Mc

[24] Anonymous. Masthead. Statistics in Medicine, 1(3):fmi, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feigl:1982:RBC

[25] Polly Feigl, Lincoln Polissar, Warren W. Lane, and Vincent Guinee. Reliability of basic cancer patient data. Statistics in Medicine, 1(3):191–204, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macfarlane:1982:NMS

[26] S. B. J. Macfarlane and J. B. Moody. Neonatal mortality in San Serriffe: a game for teaching survey methods. Statistics in Medicine, 1(3):205– 216, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1982:IES

[27] Sander Greenland. Interpretation and estimation of summary ratios under heterogeneity. Statistics in Medicine, 1(3):217–227, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittemore:1982:SME

[28] Alice S. Whittemore. Statistical methods for estimating attributable risk from retrospective data. Statistics in Medicine, 1(3):229–243, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osmond:1982:APC

[29] C. Osmond and M. J. Gardner. Age, period and cohort models applied to cancer mortality rates. Statistics in Medicine, 1(3):245–259, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1982:EPC

[30] Allan Donner. The exclusion of patients from a clinical trial. Statistics in Medicine, 1(3):261–265, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palta:1982:MLL

[31] Mari Palta and Saeid B. Amini. Magnitude and likelihood of loss resulting from non-stratified randomization. Statistics in Medicine, 1(3):267–275, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesser:1982:EUR

[32] Martin L. Lesser. The effect of unbalanced randomization on the progressively censored savage test. Statistics in Medicine, 1(3):277–280, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shirley:1982:UBD

[33] Eryl A. C. Shirley. The use of background data in the analysis of covariance. Statistics in Medicine, 1(3):281–291, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gore:1982:BRS

[34] Sheila M. Gore. Book review: Statistical models and methods for lifetime data, Jerald F. Lawless, Wiley, New York,1982. Price £D27.25. No. of pages: 580. Statistics in Medicine, 1(3):293–294, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Ac

[35] Anonymous. Announcement. Statistics in Medicine, 1(3):295, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:Md

[36] Anonymous. Masthead. Statistics in Medicine, 1(4):fmi, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:1982:ABH

[37] I. D. Hill. Austin Bradford Hill — ancestry and early life. Statistics in Medicine, 1(4):297–300, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Himsworth:1982:BHS

[38] Sir Harold Himsworth. Bradford Hill and Statistics in Medicine. Statistics in Medicine, 1(4):301–303, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1982:BHY

[39] Benedict J. Duffy, Jr. A. B. H. — young at heart. Statistics in Medicine, 1(4):305–306, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneiderman:1982:PR

[40] Marvin A. Schneiderman. A personal reminiscence. Statistics in Medicine, 1(4):307–308, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rutstein:1982:TF

[41] David D. Rutstein. Teacher and friend. Statistics in Medicine, 1(4): 309, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schoolman:1982:CS

[42] Harold M. Schoolman. The clinician and the statistician. Statistics in Medicine, 1(4):311–316, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schilling:1982:BHC

[43] Richard Schilling. Bradford Hill’s contribution to occupational health. Statistics in Medicine, 1(4):317–324, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lilienfeld:1982:BHI

[44] Abraham M. Lilienfeld. Bradford Hill’s influence on epidemiology. Statistics in Medicine, 1(4):325–328, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutherland:1982:BVB

[45] Ian Sutherland. BCG and vole bacillus vaccination in adolescence and mortality from leukaemia. Statistics in Medicine, 1(4):329–335, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doll:1982:CTR

[46] Sir Richard Doll. Clinical trials: Retrospect and prospect. Statistics in Medicine, 1(4):337–344, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1982:RRC

[47] Peter Armitage. The role of randomization in clinical trials. Statistics in Medicine, 1(4):345–352, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleiss:1982:MCT

[48] Joseph L. Fleiss. Multicentre clinical trials: Bradford Hill’s contributions and some subsequent developments. Statistics in Medicine, 1(4):353–359, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holland:1982:CTS

[49] Walter W. Holland, Ellie Breeze, and Anthony V. Swan. Clinical trials: Some reflections. Statistics in Medicine, 1(4):361–368, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:B

[50] Anonymous. Bibliography. Statistics in Medicine, 1(4):369–375, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baum:1982:BRS

[51] Michael Baum and P. G. Smith. Book review: Statistics in practice, Sheila M. Gore and Douglas G. Altman, British Medical Association, London, 1982. No. of pages: 100. Price: £D7.00. Statistics in Medicine, 1(4):377– 378, ???? 1982. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1982:D

[52] Anonymous. Dedication. Statistics in Medicine, 1(4):i, ???? 1982. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Ma

[53] Anonymous. Masthead. Statistics in Medicine, 2(1):fmi, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knox:1983:EPI

[54] E. G. Knox. Epidemiology of prenatal infections: an extension of the congenital rubella model. Statistics in Medicine, 2(1):1–12, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Majumder:1983:RRD

[55] Partha P. Majumder, Ranajit Chakraborty, and Kenneth M. Weiss. Relative risks of diseases in the presence of incomplete penetrance and sporadics. Statistics in Medicine, 2(1):13–24, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schulz:1983:RCE

[56] Kenneth F. Schulz, Willard Cates, Jr., David A. Grimes, Richard M. Selik, and Carl W. Tyler, Jr. Reducing classification errors in cohort studies: The approach and a practical application. Statistics in Medicine, 2(1):25– 31, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:1983:ARB

[57] D. J. Spiegelhalter and P. G. I. Stovin. An analysis of repeated biopsies following cardiac transplantation. Statistics in Medicine, 2(1):33–40, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kay:1983:UAT

[58] Richard Kay and Martin Schumacher. Unbiased assessment of treatment effects on disease recurrence and survival in clinical trials. Statistics in Medicine, 2(1):41–58, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oler:1983:LLM

[59] Jacqueline Oler and J. Marvin Bentley. Log-linear model selection in a rural dental health study. Statistics in Medicine, 2(1):59–69, January– March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaefer:1983:BCM

[60] Robert L. Schaefer. Bias correction in maximum likelihood logistic regression. Statistics in Medicine, 2(1):71–78, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1983:ABH

[61] J. R. Green, Mui Kui Chiew, Heng Chin Low, and J. C. Woodrow. The association between HLA antigens and the presence of certain diseases. Statistics in Medicine, 2(1):79–85, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1983:SEP

[62] D. Machin, N. R. Dennis, P. A. Tippett, and V. Andrews. On the standard error of the probability of a particular diagnosis. Statistics in Medicine, 2(1):87–93, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1983:DAB

[63] Adelin Albert. Discriminant analysis based on multivariate response curves: a descriptive approach to dynamic allocation. Statistics in Medicine, 2(1):95–106, January–March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Aa

[64] Anonymous. Announcements. Statistics in Medicine, 2(1):107, January– March 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Mb

[65] Anonymous. Masthead. Statistics in Medicine, 2(2):fmi, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1983:P

[66] T. Colton, L. S. Freedman, and A. L. Johnson. Preface. Statistics in Medicine, 2(2):109, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:ISC

[67] Anonymous. International Society for Clinical Biostatistics Third International Meeting held at Erasmus University, Rotterdam from 12–15 September 1982. Statistics in Medicine, 2(2):111, April–June 1983. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finney:1983:BML

[68] D. J. Finney. Between medicine and law. Statistics in Medicine, 2(2): 113–121, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kay:1983:IIQ

[69] C. Kay, K. W. Kemp, A. B. J. Nix, R. J. Rowlands, G. V. Groom, D. W. Wilson, and K. Griffiths. The importance of internal quality control data in national external quality assessment schemes. plasma progesterone assays. Statistics in Medicine, 2(2):123–129, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leijnse:1983:DRM

[70] B. Leijnse. Are definitive (reference) methods and Shewhart’s principle the best approach towards accuracy? Statistics in Medicine, 2(2):131–140, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rowlands:1983:ACT

[71] R. J. Rowlands, K. Griffiths, K. W. Kemp, A. B. J. Nix, G. Richards, and D. W. Wilson. Application of cusum techniques to the routine monitoring of analytical performance in clinical laboratories. Statistics in Medicine, 2(2):141–145, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1983:MAL

[72] S. G. Thompson. A method of analysis of laboratory data in an epidemiological study where time trends are present. Statistics in Medicine, 2(2): 147–153, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeJonge:1983:DCR

[73] H. De Jonge. Deficiencies in clinical reports for registration of drugs. Statistics in Medicine, 2(2):155–166, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1983:SRC

[74] L. S. Freedman, D. Lowe, and P. Macaskill. Stopping rules for clinical trials. Statistics in Medicine, 2(2):167–174, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rumke:1983:CRC

[75] Chr. L. Rümke. Critical reflections on clinical trials. Statistics in Medicine, 2(2):175–181, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1983:ASC

[76] J. Whitehead, D. R. Jones, and S. H. Ellis. The analysis of a sequential clinical trial for the comparison of two lung cancer treatments. Statistics in Medicine, 2(2):183–190, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bezemer:1983:DRN

[77] P. D. Bezemer, J. C. Netelenbos, C. Mulder, J. A. Theune, I. H. Stamhuis, and J. P. Straus. Determining reference (‘normal’) limits in medicine: an application. Statistics in Medicine, 2(2):191–198, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmitz:1983:PLD

[78] P. I. M. Schmitz, J. D. F. Habbema, and J. Hermans. The performance of logistic discrimination on myocardial infarction data, in comparison with some other discriminant analysis methods. Statistics in Medicine, 2(2): 199–205, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:1983:ECD

[79] D. J. Spiegelhalter. Evaluation of clinical decision-aids, with an application to a system for dyspepsia. Statistics in Medicine, 2(2):207–216, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tate:1983:ATM

[80] Helen Tate. Assessing tumour markers. Statistics in Medicine, 2(2):217– 222, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1983:AUP

[81] M. J. Campbell. The analysis using genstat of anaemia, sugar intake and Quetelet’s index as prognostic indicators in women. Statistics in Medicine, 2(2):223–228, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayton:1983:LPM

[82] David G. Clayton. The Leicestershire Perinatal Mortality Study: a case study of multi-group discriminant analysis with complex sampling. Statistics in Medicine, 2(2):229–242, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1983:TIE

[83] Sander Greenland. Tests for interaction in epidemiologic studies: a review and a study of power. Statistics in Medicine, 2(2):243–251, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McManus:1983:BBP

[84] I. C. McManus. Bimodality of blood pressure levels. Statistics in Medicine, 2(2):253–258, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mantel:1983:SAM

[85] Nathan Mantel, Paolo Mocarelli, Alessandro Marocchi, Paolo Brambilla, and Roberto Baretta. Stratified analysis of multivariate clinical data: Application of a Mantel–Haenszel approach. Statistics in Medicine, 2(2): 259–266, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miettinen:1983:NRS

[86] Olli S. Miettinen. The need for randomization in the study of intended effects. Statistics in Medicine, 2(2):267–271, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bennett:1983:ASD

[87] Steve Bennett. Analysis of survival data by the proportional odds model. Statistics in Medicine, 2(2):273–277, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Habbema:1983:ASD

[88] J. D. F. Habbema, G. J. Van Oortmarssen, and D. J. Van Putten. An analysis of survival differences between clinically and screen-detected cancer patients. Statistics in Medicine, 2(2):279–285, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hudec:1983:FRM

[89] M. Hudec and H. Platz. Fitting regression models to censored survival data. Statistics in Medicine, 2(2):287–293, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marubini:1983:PFR

[90] E. Marubini, A. Morabito, and M. G. Valsecchi. Prognostic factors and risk groups: Some results given by using an algorithm suitable for censored survival data. Statistics in Medicine, 2(2):295–303, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oakes:1983:CMS

[91] David Oakes. Comparison of models for survival data. Statistics in Medicine, 2(2):305–311, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Ab

[92] Anonymous. Announcements. Statistics in Medicine, 2(2):313, April–June 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Mc

[93] Anonymous. Masthead. Statistics in Medicine, 2(3):fmi, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michalek:1983:ULS

[94] Joel E. Michalek and Daniel Mihalko. On the use of logrank scores in the analysis of litter-matched data on time to tumour appearance. Statistics in Medicine, 2(3):315–321, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See commentary [95].

Mantel:1983:CWM

[95] Nathan Mantel. Commentary on the work of Michalek and Mihalko relative to litter-matched data. Statistics in Medicine, 2(3):323–326, July– September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [94].

Marcuson:1983:SCM

[96] Richard Marcuson. Simplified computation of the multivariate permutation test for arbitrarily censored survival data. Statistics in Medicine, 2 (3):327–330, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1983:RAA

[97] Donald A. Berry and Terrance L. Fox. Regression analysis applied to PVC histories: a statistical procedure for evaluating antiarrhythmic drug efficacy. Statistics in Medicine, 2(3):331–343, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:1983:SAG

[98] M. Anthony Wong, Sheldon Simon, and Richard A. Olshen. Statistical analysis of gait patterns of persons with cerebral palsy. Statistics in Medicine, 2(3):345–354, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mantel:1983:CUM

[99] Nathan Mantel. Cautions on the use of medical databases. Statistics in Medicine, 2(3):355–362, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bell:1983:PMP

[100] C. M. Janine Bell and Derek A. Coleman. Predicted mortality patterns in cohort study populations exposed to different types of hazard: Can SMRS show a dose-response? Statistics in Medicine, 2(3):363–371, July– September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siler:1983:PMH

[101] William Siler. Parameters of mortality in human populations with widely varying life spans. Statistics in Medicine, 2(3):373–380, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartz:1983:CTM

[102] Arthur J. Hartz, Eldred E. Giefer, and Raymond G. Hoffmann. A comparison of two methods for calculating expected mortality. Statistics in Medicine, 2(3):381–386, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgenstern:1983:MDS

[103] Hal Morgenstern and Deborah M. Winn. A method for determining the sampling ratio in epidemiologic studies. Statistics in Medicine, 2(3):387– 396, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1983:SNP

[104] Gary G. Koch, Saul L. Gitomer, Lorie Skalland, and Maura E. Stokes. Some non-parametric and categorical data analyses for a change-over design study and discussion of apparent carry-over effects. Statistics in Medicine, 2(3):397–412, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Behar:1983:LFA

[105] S. Behar. The logistic function as an aid in the detection of acute coronary disease in emergency patients (a case study). Statistics in Medicine, 2 (3):413–416, July–September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1983:Ac

[106] Anonymous. Announcement. Statistics in Medicine, 2(3):416, July– September 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McDonald:1983:HMP

[107] Clement J. McDonald, Steven A. Mazzuca, and George P. McCabe, Jr. How much of the placebo ‘effect’ is really statistical regression? Statistics in Medicine, 2(4):417–427, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1983:MAQ

[108] P. M. Fayers and D. R. Jones. Measuring and analysing quality of life in cancer clinical trials: a review. Statistics in Medicine, 2(4):429–446, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chan:1983:AAR

[109] Linda S. Chan, Gary D. Overturf, and Jeffrey Selzer. Analysing antibody response based on data obtained from serial dilution methods. Statistics in Medicine, 2(4):447–454, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inskip:1983:MAA

[110] Hazel Inskip, Valerie Beral, Patricia Fraser, and John Haskey. Methods for age-adjustment of rates. Statistics in Medicine, 2(4):455–466, October– December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McDowall:1983:APM

[111] Michael McDowall. Adjusting proportional mortality ratios for the influence of extraneous causes of death. Statistics in Medicine, 2(4):467–475, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paddle:1983:IEI

[112] G. M. Paddle and B. Bennett. An investigation of an excess incidence of heart disease. Statistics in Medicine, 2(4):477–484, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:1983:PIS

[113] Ron Brookmeyer. Prediction intervals for survival data. Statistics in Medicine, 2(4):485–495, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hearne:1983:VPI

[114] Erwin M. Hearne III and Gary M. Clark. Validity and power investigation of a test for serial correlation in univariate repeated measures analysis. Statistics in Medicine, 2(4):497–504, October–December 1983. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Ma

[115] Anonymous. Masthead. Statistics in Medicine, 3(1):fmi, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moses:1984:SCC

[116] Lincoln Moses and Thomas A. Louis. Statistical consulting in clinical research: The two-way street. Statistics in Medicine, 3(1):1–5, January– March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roman:1984:CSP

[117] Eve Roman, Valerie Beral, Hazel Inskip, Michael McDowall, and Abe Adelstein. A comparison of standardized and proportional mortality ratios. Statistics in Medicine, 3(1):7–14, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lilienfeld:1984:LAE

[118] David E. Lilienfeld and David A. Pyne. The logistic analysis of epidemiologic prospective studies: Investigation by simulation. Statistics in Medicine, 3(1):15–26, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenbaum:1984:CDP

[119] Paul R. Rosenbaum. Case definition and power in case-control studies. Statistics in Medicine, 3(1):27–34, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:1984:NPG

[120] Richard Simon and Robert W. Makuch. A non-parametric graphical representation of the relationship between survival and the occurrence of an event: Application to responder versus non-responder bias. Statistics in Medicine, 3(1):35–44, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baskerville:1984:CTD

[121] J. C. Baskerville, J. H. Toogood, J. Mazza, and Barbara Jennings. Clinical trials designed to evaluate therapeutic preferences. Statistics in Medicine, 3(1):45–55, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sturt:1984:ALA

[122] Elizabeth Sturt. Analysis of linkage and association for diseases of genetic aetiology. Statistics in Medicine, 3(1):57–72, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Musch:1984:AKT

[123] David C. Musch, J. Richard Landis, Ian T. T. Higgins, John C. Gilson, and Robert N. Jones. An application of kappa-type analyses to interobserver variation in classifying chest radiographs for pneumoconiosis. Statistics in Medicine, 3(1):73–83, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shirley:1984:CBA

[124] Eryl A. C. Shirley and Peter Newnham. The choice between analysis of variance and analysis of covariance with special reference to the analysis of organ weights in toxicology studies. Statistics in Medicine, 3(1):85– 91, January–March 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:1984:BRB

[125] Joel E. Cohen. Book review: The biomathematics of malaria. The biomathematics of diseases NO. 1, Norman T. J. Bailey, Charles Griffin and Co., London, 1982. No of pages: xii + 210. Price: £16.50, $88.00. Statistics in Medicine, 3(1):93–95, January–March 1984. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Mb

[126] Anonymous. Masthead. Statistics in Medicine, 3(2):fmi, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:C

[127] Anonymous. Correction. Statistics in Medicine, 3(2):i, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1984:E

[128] Tony Johnson. Editorial. Statistics in Medicine, 3(2):97–99, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marinez:1984:EDQ

[129] Yolan N. Marinez, C. Alex McMahan, George M. Barnwell, and Herman S. Wigodsky. Ensuring data quality in medical research through an integrated data management system. Statistics in Medicine, 3(2):101–111, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gardner:1984:ITT

[130] M. J. Gardner and C. Osmond. Interpretation of time trends in disease rates in the presence of generation effects. Statistics in Medicine, 3(2): 113–130, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1984:BMD

[131] Sander Greenland. Bias in methods for deriving standardized morbidity ratio and attributable fraction estimates. Statistics in Medicine, 3(2): 131–141, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrell:1984:RMS

[132] Frank E. Harrell, Jr., Kerry L. Lee, Robert M. Califf, David B. Pryor, and Robert A. Rosati. Regression modelling strategies for improved prognostic prediction. Statistics in Medicine, 3(2):143–152, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldman:1984:SAW

[133] Anne I. Goldman. Survivorship analysis when cure is a possibility: a Monte Carlo study. Statistics in Medicine, 3(2):153–163, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimber:1984:ART

[134] Hilary Kimber and Martin Crowder. An analysis of resistance times to infection under treatment. Statistics in Medicine, 3(2):165–171, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1984:STT

[135] Rina Chen, Nathan Mantel, and Marcus A. Klingberg. A study of three techniques for time–space clustering in Hodgkin’s disease. Statistics in Medicine, 3(2):173–184, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keselman:1984:ARM

[136] H. J. Keselman and Joanne C. Keselman. The analysis of repeated measures designs in medical research. Statistics in Medicine, 3(2):185–195, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kay:1984:BRR

[137] Richard Kay. Book review: Regression analysis of survival data in cancer chemotherapy, Walter H. Carter, Jr., Galen L. Wampler and Donald M. Stablein, Marcel Dekker, New York and Basel, 1983. Price: SFr. 103. No. of pages: 209. Statistics in Medicine, 3(2):197, April–June 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Mc

[138] Anonymous. Masthead. Statistics in Medicine, 3(3):fmi, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1984:ASS

[139] Allan Donner. Approaches to sample size estimation in the design of clinical trials — a review. Statistics in Medicine, 3(3):199–214, July– September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McHugh:1984:VTD

[140] R. McHugh. Validity and treatment dilution in Zelen’s single consent design. Statistics in Medicine, 3(3):215–218, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanLier:1984:ATP

[141] Henk J. J. Van Lier, Philip Van Elteren, Andries J. Hoitsma, and Robert A. P. Koene. Allocation of treatment to patients when not all balancing factors are known to the co-ordinating centre. Statistics in Medicine, 3 (3):219–224, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoehler:1984:MRH

[142] Fred K. Hoehler, Nathan Mantel, Edmund Gehan, Esther Kahana, and Milton Alter. Medical registers as historical controls: Analysis of an open clinical trial of inosiplex in subacute sclerosing panencephalitis. Statistics in Medicine, 3(3):225–237, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barbujani:1984:CTS

[143] Guido Barbujani and Elisa Calzolari. Comparison of two statistical techniques for the surveillance of birth defects through a Monte Carlo simulation. Statistics in Medicine, 3(3):239–247, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickering:1984:CSI

[144] R. M. Pickering and J. F. Forbes. A classification of Scottish infants using latent class analysis. Statistics in Medicine, 3(3):249–259, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stoddard:1984:ACB

[145] Anne M. Stoddard. Automatic classification of biomedical information when classification error is unknown. Statistics in Medicine, 3(3):261– 271, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1984:CCS

[146] Christopher Cox and Christy Chuang. A comparison of chi-square partitioning and two logit analyses of ordinal pain data from a pharmaceutical study. Statistics in Medicine, 3(3):273–285, July–September 1984. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Regal:1984:GFB

[147] Ronald R. Regal and Ernest B. Hook. Goodness-of-fit based confidence intervals for estimates of the size of a closed population. Statistics in Medicine, 3(3):287–291, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hadgu:1984:ARR

[148] Alula Hadgu. An application of ridge regression analysis in the study of syphilis data. Statistics in Medicine, 3(3):293–299, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:CTM

[149] Anonymous. A comparison of two methods for calculating expected mortality. Statistics in Medicine, 3(3):301–302, July–September 1984. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1984:BRS

[150] M. K. Palmer. Book review: Statistics in medical research: Methods and issues, with applications in cancer research, Valerie Mike and Kenneth E. Stanley (Eds), Wiley, New York, 1983. no. of pages: 576. price: £35.65 ($49.90). Statistics in Medicine, 3(3):303, July–September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haybittle:1984:BRC

[151] John Haybittle. Book review: Cancer clinical trials: Methods and practice M. E. Buyse, M. J. Staquet and R. J. Sylvester (Eds.), Oxford University Press, Oxford, 1984. price $35. Statistics in Medicine, 3(3):303–304, July– September 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Md

[152] Anonymous. Masthead. Statistics in Medicine, 3(4):fmi, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedman:1984:P

[153] Lawrence Friedman, Richard Simon, Joel Verter, Janet Wittes, and Robert Wiites. Preface. Statistics in Medicine, 3(4):305, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1984:RCT

[154] Byron W. Brown, Jr. The randomized clinical trial. Statistics in Medicine, 3(4):307–311, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Da

[155] Anonymous. Discussion. Statistics in Medicine, 3(4):313–314, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gehan:1984:ETH

[156] Edmund A. Gehan. The evaluation of therapies: Historical control studies. Statistics in Medicine, 3(4):315–324, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinsky:1984:EHC

[157] Carl M. Pinsky. Experience with historical control studies in cancer immunotherapy. Statistics in Medicine, 3(4):325–329, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Db

[158] Anonymous. Discussion. Statistics in Medicine, 3(4):331–332, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moon:1984:UDP

[159] Thomas E. Moon, Stephen E. Jones, Gianni Bonadonna, Trevor J. Powles, Saul Rivkin, Aman Buzdar, and Eleanor Montague. Using a database of protocol studies to evaluate therapy: a breast cancer example. Statistics in Medicine, 3(4):333–339, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinstein:1984:ROS

[160] Alvan Feinstein. The role of observational studies in the evaluation of therapy. Statistics in Medicine, 3(4):341–345, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dc

[161] Anonymous. Discussion. Statistics in Medicine, 3(4):347–351, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kramer:1984:PCS

[162] Simon Kramer and James J. Diamond. The patterns of care study in clinical radiation therapy — a retrospective assessment of quality of care. Statistics in Medicine, 3(4):353–357, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dd

[163] Anonymous. Discussion. Statistics in Medicine, 3(4):359–360, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1984:UOD

[164] Sylvan B. Green and David P. Byar. Using observational data from registries to compare treatments: The fallacy of omnimetrics. Statistics in Medicine, 3(4):361–370, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:De

[165] Anonymous. Discussion. Statistics in Medicine, 3(4):371–373, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hlatky:1984:TCR

[166] Mark A. Hlatky, Kerry L. Lee, Frank E. Harrell, Jr., Robert M. Califf, David B. Pryor, Daniel B. Mark, and Robert A. Rosati. Tying clinical research to patient care by use of an observational database. Statistics in Medicine, 3(4):375–384, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Df

[167] Anonymous. Discussion. Statistics in Medicine, 3(4):385–387, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Detre:1984:NRS

[168] Katherine M. Detre. Non-randomized studies of coronary artery bypass surgery. Statistics in Medicine, 3(4):389–398, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelen:1984:C

[169] Marvin Zelen. Commentary. Statistics in Medicine, 3(4):399–400, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blackburn:1984:COV

[170] Henry Blackburn. Commentary: Observation versus experiment. Statistics in Medicine, 3(4):401–403, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dg

[171] Anonymous. Discussion. Statistics in Medicine, 3(4):405–408, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yusuf:1984:WDW

[172] Salim Yusuf, Rory Collins, and Richard Peto. Why do we need some large, simple randomized trials? Statistics in Medicine, 3(4):409–420, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dh

[173] Anonymous. Discussion. Statistics in Medicine, 3(4):421–422, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Canner:1984:HMD

[174] Paul L. Canner. How much data should be collected in a clinical trial? experience of the coronary drug project. Statistics in Medicine, 3(4): 423–432, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Di

[175] Anonymous. Discussion. Statistics in Medicine, 3(4):433, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:1984:PPR

[176] Joseph F. Collins, William O. Williford, David G. Weiss, Stephen F. Bingham, and C. James Klett. Planning patient recruitment: Fantasy and reality. Statistics in Medicine, 3(4):435–443, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1984:CES

[177] David L. Demets. Can early stopping procedures impact significantly on the efficiency of clinical trials without serious loss of information? Statistics in Medicine, 3(4):445–451, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dj

[178] Anonymous. Discussion. Statistics in Medicine, 3(4):453–454, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moertel:1984:IEC

[179] Charles G. Moertel. Improving the efficiency of clinical trials: a medical perspective. Statistics in Medicine, 3(4):455–465, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:Dk

[180] Anonymous. Discussion. Statistics in Medicine, 3(4):467–468, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:1984:C

[181] O. Dale Williams. Commentary. Statistics in Medicine, 3(4):469–471, October–December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meier:1984:C

[182] Paul Meier. Commentary. Statistics in Medicine, 3(4):473–475, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1984:A

[183] Anonymous. Announcement. Statistics in Medicine, 3(4):479, October– December 1984. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Ma

[184] Anonymous. Masthead. Statistics in Medicine, 4(1):fmi, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:C

[185] Anonymous. Correction. Statistics in Medicine, 4(1):i, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1985:MAI

[186] Colin B. Begg. A measure to aid in the interpretation of published clinical trials. Statistics in Medicine, 4(1):1–9, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charlton:1985:CTS

[187] J. R. H. Charlton, M. F. D’souza, M. Tooley, and R. Silver. A community trial strategy for evaluating treatment for symptomatic conditions. Statistics in Medicine, 4(1):11–21, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1985:NIE

[188] Chung-Cheng Hsieh. Note on interval estimation of the difference between proportions from correlated series. Statistics in Medicine, 4(1):23– 27, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hougaard:1985:DES

[189] Philip Hougaard and Erling Birk Madsen. Dynamic evaluation of shortterm prognosis after myocardial infarction. Statistics in Medicine, 4(1): 29–38, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1985:BIP

[190] Chen-Hsin Chen and Stephen L. George. The bootstrap and identification of prognostic factors via Cox’s proportional hazards regression model. Statistics in Medicine, 4(1):39–46, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1985:IRA

[191] Edward L. Korn and Sheldon Baumrind. Inter-rater agreement among orthodontists in a blocked experiment. Statistics in Medicine, 4(1):47– 52, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1985:EVG

[192] Manfred S. Green. The effect of validation group bias on screening tests for coronary artery disease. Statistics in Medicine, 4(1):53–61, January– March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schifflers:1985:BCA

[193] E. Schifflers, M. Smans, and C. S. Muir. Birth cohort analysis using irregular crosssectional data: a technical note. Statistics in Medicine, 4 (1):63–75, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Unkelbach:1985:DRA

[194] Hans Dieter Unkelbach and Thomas Wolf. Dose-response analysis of combination preparations. Statistics in Medicine, 4(1):77–85, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1985:WTTa

[195] Jack Cuzick. A Wilcoxon-type test for trend. Statistics in Medicine, 4(1): 87–90, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1985:AND

[196] Daniel Commenges and John Seal. The analysis of neuronal discharge sequences: Change-point estimation and comparison of variances. Statistics in Medicine, 4(1):91–104, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartz:1985:CTM

[197] A. J. Hartz, E. E. Giefer, and R. G. Hoffmann. A comparison of two methods for calculating expected mortality. Statistics in Medicine, 4(1):105– 109, January–March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Aa

[198] Anonymous. Announcement. Statistics in Medicine, 4(1):110, January– March 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Mb

[199] Anonymous. Masthead. Statistics in Medicine, 4(2):fmi, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stampfer:1985:FDA

[200] Meir J. Stampfer, Julie E. Buring, Walter Willett, Bernard Rosner, Kimberly Eberlein, and Charles H. Hennekens. The 2 × 2 factorial design: Its application to a randomized trial of aspirin and U.S. physicians. Statistics in Medicine, 4(2):111–116, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1985:PSS

[201] Sander Greenland. Power, sample size and smallest detectable effect determination for multivariate studies. Statistics in Medicine, 4(2):117–127, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalish:1985:TAM

[202] Leslie A. Kalish and Colin B. Begg. Treatment allocation methods in clinical trials: a review. Statistics in Medicine, 4(2):129–144, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maritz:1985:MUS

[203] J. S. Maritz. Models and the use of signed rank tests. Statistics in Medicine, 4(2):145–153, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1985:FCD

[204] David P. Smith. Formulae for cause-deleted life tables. Statistics in Medicine, 4(2):155–162, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Majumder:1985:CAB

[205] Partha P. Majumder. Comparison of ascertainment-bias correction schemes for pedigrees ascertained through multiple probands. Statistics in Medicine, 4(2):163–173, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murrells:1985:ULMa

[206] T. J. Murrells, T. M. F. Smith, J. C. Catford, and D. Machin. The use of logit models to investigate social and biological factors in infant mortality. i: Methodology. Statistics in Medicine, 4(2):175–187, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murrells:1985:ULMb

[207] T. J. Murrells, J. C. Catford, T. M. F. Smith, and D. Machin. The use of logit models to investigate social and biological factors in infant mortality. II: Stillbirths. Statistics in Medicine, 4(2):189–200, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsutakawa:1985:EBE

[208] Robert K. Tsutakawa, Gary L. Shoop, and Carl J. Marienfeld. Empirical Bayes estimation of cancer mortality rates. Statistics in Medicine, 4(2): 201–212, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miettinen:1985:CAT

[209] Olli Miettinen and Markku Nurminen. Comparative analysis of two rates. Statistics in Medicine, 4(2):213–226, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halperin:1985:ULM

[210] Max Halperin, Robert D. Abbott, William C. Blackwelder, Ronald Jacobowitz, K. K. Lan Gordon, Joel I. Verter, and Hans Wedel. On the use of the logistic model in prospective studies. Statistics in Medicine, 4(2): 227–235, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lilienfeld:1985:LAE

[211] D. E. Lilienfeld, D. A. Pyne, and James R. Anderson. The logistic analysis of epidemiologic prospective studies: Investigation by simulation. Statistics in Medicine, 4(2):237–242, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaefer:1985:BCM

[212] R. L. Schaefer and S. C. Richardson. Bias correction in maximum likelihood logistic regression. Statistics in Medicine, 4(2):243–245, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1985:ASS

[213] A. Donner and Robert Makuch. Approaches to sample size estimation in the design of clinical trials — a review. Statistics in Medicine, 4(2):247, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tate:1985:BRU

[214] Helen Tate. Book review: Understanding medical statistics. Leonard A. Goldstone. Heinemann Medical Books Ltd., London, 1983. No. of pages: x + 181. Price: £7.50. Statistics in Medicine, 4(2):249, April–June 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Mc

[215] Anonymous. Masthead. Statistics in Medicine, 4(3):fmi, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1985:P

[216] Ted Colton, Laurence Freedman, and Tony Johnson. Preface. Statistics in Medicine, 4(3):251, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:ISC

[217] Anonymous. International Society for Clinical Biostatistics Fifth International Meeting held at Palazzo dei Congressi, San Marino from 9–13 September 1984. Statistics in Medicine, 4(3):253, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1985:AAT

[218] David P. Byar. Assessing apparent treatment-covariate interactions in randomized clinical trials. Statistics in Medicine, 4(3):255–263, July– September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cascinelli:1985:PFS

[219] N. Cascinelli, E. Marubini, A. Morabito, and R. Bufalino. Prognostic factors for stagei melanoma of the skin: a review. Statistics in Medicine, 4(3):265–278, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chastang:1985:PFR

[220] Claude Chastang, Bernard Lebeau, Yves Charpak, and Guy Decroix. Prognostic factors from a randomized clinical trial in resected lung cancer. Statistics in Medicine, 4(3):279–285, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chastang:1985:CDA

[221] Claude Chastang, Philippe Travade, and Ariane Auquier. Critical discussion of the assessment of a three-stage prognostic classification for chronic lymphocytic leukemia. Statistics in Medicine, 4(3):287–293, July– September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:1985:CMC

[222] Catherine Hill, Agnes Laplanche, and Ali Rezvani. Comparison of the mortality of a cohort with the mortality of a reference population in a prognostic study. Statistics in Medicine, 4(3):295–302, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Micciolo:1985:CAS

[223] R. Micciolo, I. Vantini, G. Cavallini, W. Piubello, G. Talamini, L. Benini, and L. A. Scuro. Correspondence analysis in a study of the clinical evolution of uncomplicated chronic relapsing alcoholic pancreatitis. Statistics in Medicine, 4(3):303–309, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teather:1985:CAC

[224] D. Teather, B. A. Morton, G. H. du Boulay, K. M. Wills, D. Plummer, and P. R. Innocent. Computer assistance for C.T. scan interpretation and cerebral disease diagnosis. Statistics in Medicine, 4(3):311–315, July– September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Torasso:1985:KBE

[225] Pietro Torasso. Knowledge based expert systems for medical diagnosis. Statistics in Medicine, 4(3):317–325, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaldor:1985:LCA

[226] John Kaldor and David Clayton. Latent class analysis in chronic disease epidemiology. Statistics in Medicine, 4(3):327–335, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1985:HTP

[227] R. J. Marshall and E. M. Chisholm. Hypothesis testing in the polychotomous logistic model with an application to detecting gastrointestinal cancer. Statistics in Medicine, 4(3):337–344, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Filippazzo:1985:ASR

[228] M. G. Filippazzo, E. Aragona, M. Cottone, G. Dardanoni, F. Lanzarone, A. Marenghini, C. Patti, E. Sciarrino, R. Simonetti, F. Tine, and L. Pagliaro. Assessment of some risk factors for hepatocellular carcinoma: a case control study. Statistics in Medicine, 4(3):345–351, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1985:MAM

[229] M. J. Campbell. Multiplicative and additive models with external controls in a cohort study of cancer mortality. Statistics in Medicine, 4(3):353– 360, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bassein:1985:CTD

[230] Leona Bassein, Claudio Borghi, Francesco Vittorio Costa, Enrico Strocchi, Alessandra Mussi, and Ettore Ambrosioni. Comparison of three devices for measuring blood pressure. Statistics in Medicine, 4(3):361–368, July– September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boerma:1985:RCC

[231] G. J. M. Boerma. Are routine clinical chemistry analyses ‘in control’ according to Shewhart’s technique? Statistics in Medicine, 4(3):369–378, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bossi:1985:LMA

[232] Anna Bossi and Silvano Milani. Linear models for the analysis of variability in factorial designs: an application to anthropometric indicators. Statistics in Medicine, 4(3):379–386, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miettinen:1985:LE

[233] O. S. Miettinen, Guido Barbujani, and Elisa Calzolari. Letters to the editor. Statistics in Medicine, 4(3):387–391, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1985:BRM

[234] M. J. R. Healy. Book review: Mathematical methods in medicine, part I: Statistical and analytical techniques, D. Ingram and R. Bloch (eds), Wiley, Chichester, 1984. no of pages: 474. price: £29.95, $45.00. Statistics in Medicine, 4(3):393–394, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1985:BRB

[235] Tony Johnson. Book review: Biostatistics in clinical medicine, Joseph A. Ingelfinger, Frederick Mosteller, L. A. Thibodeau and James H. Ware, Macmillan Publishing Co., New York, 1983, No. of pages: xxiii + 316.

Price: £22.95. Statistics in Medicine, 4(3):394, July–September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Ab

[236] Anonymous. Announcement. Statistics in Medicine, 4(3):395, July– September 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Md

[237] Anonymous. Masthead. Statistics in Medicine, 4(4):fmi, October– December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1985:PIH

[238] David P. Smith and Carl Slater. The pattern of ischemic heart disease 1931–1980. Statistics in Medicine, 4(4):397–412, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brenn:1985:SRF

[239] Tormod Brenn and Egil Arnesen. Selecting risk factors: a comparison of discriminant analysis, logistic regression and Cox’s regression model using data from the Tromsø Heart Study. Statistics in Medicine, 4(4):413–423, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moolgavkar:1985:AAL

[240] Suresh H. Moolgavkar, Edward D. Lustbader, and David J. Venzon. Assessing the adequacy of the logistic regression model for matched casecontrol studies. Statistics in Medicine, 4(4):425–435, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1985:SSE

[241] S. D. Walter. Small sample estimation of log odds ratios from logistic regression and fourfold tables. Statistics in Medicine, 4(4):437–444, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayton:1985:UTR

[242] David Clayton. Using test-retest reliability data to improve estimates of relative risk: an application of latent class analysis. Statistics in Medicine, 4(4):445–455, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1985:UAM

[243] Bernard Rosner, Alvaro Muñoz, Ira Tager, Frank Speizer, and Scott Weiss. The use of an autoregressive model for the analysis of longitudinal data in epidemiologic studies. Statistics in Medicine, 4(4):457–467, October– December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levin:1985:CTH

[244] Bruce Levin and Jennie Kline. The cusum test of homogeneity with an application in spontaneous abortion epidemiology. Statistics in Medicine, 4(4):469–488, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shirley:1985:VST

[245] Eryl A. C. Shirley. The value of specialized tests in studies where ordered group means are expected. Statistics in Medicine, 4(4):489–496, October– December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1985:ODC

[246] Donald A. Berry and Larry M. Pearson. Optimal designs for clinical trials with dichotomous responses. Statistics in Medicine, 4(4):497–508, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Achcar:1985:ABA

[247] Jorge A. Achcar, Ron Brookmeyer, and William G. Hunter. An application of Bayesian analysis to medical follow-up data. Statistics in Medicine, 4(4):509–520, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1985:IAC

[248] Donald A. Berry. Interim analyses in clinical trials: Classical vs. Bayesian approaches. Statistics in Medicine, 4(4):521–526, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1985:ACU

[249] Edward L. Korn, Vernon K. Sondak, Carl A. Bertelsen, and David H. Kern. Analysis of the clinical utility of a predictive chemosensitivity assay. Statistics in Medicine, 4(4):527–534, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bennett:1985:TEP

[250] B. M. Bennett. On tests for equality of predictive values for t diagnostic procedures. Statistics in Medicine, 4(4):535–539, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1985:NOS

[251] S. D. Walter and H. Morgenstern. A note on optimal sampling for the comparison of proportions or rates. Statistics in Medicine, 4(4):541–542, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1985:WTTb

[252] Jack Cuzick. A Wilcoxon-type test for trend. Statistics in Medicine, 4(4): 543–547, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1985:BRA

[253] Stuart Pocock. Book review: Atlas of cancer mortality in England and Wales 1968–1978. M. J. Gardner, P. D. Winter, C. P. Taylor and E. D. Acheson, Wiley, Chichester, 1983. No. of pages: 128. Price: £47.25, $89.25. Atlas of mortality from selected diseases in England and Wales 1968–1978. M. J. Gardner, P. D. Winter and D. J. P. Barker, Wiley, Chichester, 1984. No. of pages: 108. Price: £59.00, $85.00. Statistics in Medicine, 4(4):549, October–December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1985:Ac

[254] Anonymous. Announcement. Statistics in Medicine, 4(4):550, October– December 1985. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Ma

[255] Anonymous. Masthead. Statistics in Medicine, 5(1):fmi, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:C

[256] Anonymous. Correction. Statistics in Medicine, 5(1):i, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:1986:PAS

[257] D. J. Spiegelhalter and L. S. Freedman. A predictive approach to selecting the size of a clinical trial, based on subjective clinical opinion. Statistics in Medicine, 5(1):1–13, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang:1986:NMO

[258] Christy Chuang and Alan Agresti. A new model for ordinal pain data from a pharmaceutical study. Statistics in Medicine, 5(1):15–20, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rindskopf:1986:VLC

[259] David Rindskopf and Wallace Rindskopf. The value of latent class analysis in medical diagnosis. Statistics in Medicine, 5(1):21–27, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1986:CNC

[260] Jeremy M. G. Taylor. Choosing the number of controls in a matched case-control study, some sample size, power and efficiency considerations. Statistics in Medicine, 5(1):29–36, January–February 1986. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helfenstein:1986:BJM

[261] Ulrich Helfenstein. Box–Jenkins modelling of some viral infectious diseases. Statistics in Medicine, 5(1):37–47, January–February 1986. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:1986:SMR

[262] D. C. Thomas, M. Goldberg, R. Dewar, and J. Siemiatycki. Statistical methods for relating several exposure factors to several diseases in case-heterogeneity studies. Statistics in Medicine, 5(1):49–60, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1986:CDP

[263] S. D. Walter. Cause-deleted proportional mortality analysis and the healthy worker effect. Statistics in Medicine, 5(1):61–71, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1986:ASC

[264] Peter F. Thall and John M. Lachin. Assessment of stratum-covariate interactions in Cox’s proportional hazards regression model. Statistics in Medicine, 5(1):73–83, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciampi:1986:RAC

[265] Antonio Ciampi, Sheilah A. Hogg, and Louis Kates. Regression analysis of censored survival data with the generalized F family — an alternative to the proportional hazards model. Statistics in Medicine, 5(1):85–96, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1986:TNP

[266] L. S. Freedman. Tables of the number of patients required in clinical trials using the logrank test. Statistics in Medicine, 5(1):97–98, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barbujani:1986:CTS

[267] G. Barbujani and E. Calzolari. Comparison of two statistical techniques for the surveillance of birth defects through a Monte Carlo simulation. Statistics in Medicine, 5(1):98–99, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Aa

[268] Anonymous. Announcement. Statistics in Medicine, 5(1):100, January– February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meade:1986:BRR

[269] T. W. Meade. Book review: Research into health and illness: Issues in design, analysis and practice, Andrew F. Long, Gower Publishing Co. Ltd., Aldershot, 1984. No. of pages: vi + 197. Price: £15.00. Statistics in Medicine, 5(1):101, January–February 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Mb

[270] Anonymous. Masthead. Statistics in Medicine, 5(2):fmi, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1986:UID

[271] G. D. Murray. Use of an international data bank to compare outcome following severe head injury in different centres. Statistics in Medicine, 5(2):103–112, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peters:1986:PSI

[272] Tim J. Peters and Jean Golding. Prediction of sudden infant death syndrome: an independent evaluation of four scoring methods. Statistics in Medicine, 5(2):113–126, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:1986:CCS

[273] R. Brookmeyer, N. E. Day, and S. Moss. Case-control studies for estimation of the natural history of preclinical disease from screening data. Statistics in Medicine, 5(2):127–138, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1986:ULMa

[274] D. Machin, T. J. Murrells, J. C. Catford, and T. M. F. Smith. The use of logit models to investigate social and biological factors in infant mortality. III: Neonatal mortality. Statistics in Medicine, 5(2):139–153, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1986:ULMb

[275] D. Machin, T. J. Murrells, J. C. Catford, and T. M. F. Smith. The use of logit models to investigate social and biological factors in infant mortality. IV: Post-neonatal mortality. Statistics in Medicine, 5(2):155– 169, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huel:1986:AGC

[276] Guy Huel, Jean-Francois Petiot, and Philippe Lazar. Algorithm for the grouping of contiguous geographical zones. Statistics in Medicine, 5(2): 171–181, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kay:1986:DII

[277] C. Kay, A. B. J. Nix, K. W. Kemp, R. J. Rowlands, G. Richards, G. V. Groom, K. Griffiths, and D. W. Wilson. Dose-interpolation of immunoassay data: Uncertainties associated with curve-fitting. Statistics in Medicine, 5(2):183–192, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:1986:CSC

[278] J. M. Bland. Computer simulation of a clinical trial as an aid to teaching the concept of statistical significance. Statistics in Medicine, 5(2):193– 197, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1986:WTT

[279] Jack Cuzick. A Wilcoxon-type test for trend. Statistics in Medicine, 5 (2):199, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maritz:1986:MUS

[280] J. S. Maritz. Models and the use of signed rank tests. Statistics in Medicine, 5(2):199–200, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Ab

[281] Anonymous. Announcements. Statistics in Medicine, 5(2):201–202, March–April 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Mc

[282] Anonymous. Masthead. Statistics in Medicine, 5(3):fmi, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:1986:PIR

[283] Walter W. Hauck and Sharon Anderson. A proposal for interpreting and reporting negative studies. Statistics in Medicine, 5(3):203–209, May– June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halpern:1986:STA

[284] Jerry Halpern and Byron Wm. Brown, Jr. Sequential treatment allocation procedures in clinical trials-with particular attention to the analysis of results for the biased coin design. Statistics in Medicine, 5(3):211–229, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Madsen:1986:UEB

[285] K. S. Madsen, J. P. Miller, and M. A. Province. The use of an extended baseline period in the evaluation of treatment in a longitudinal Duchenne muscular dystrophy trial. Statistics in Medicine, 5(3):231–241, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1986:PAC

[286] Per Kragh Andersen and Niels Kr. Rasmussen. Psychiatric admissions and choice of abortion. Statistics in Medicine, 5(3):243–253, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1986:CDM

[287] Edward L. Korn. Censoring distributions as a measure of follow-up in survival analysis. Statistics in Medicine, 5(3):255–260, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelen:1986:CCS

[288] M. Zelen and R. A. Parker. Case-control studies and Bayesian inference. Statistics in Medicine, 5(3):261–269, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grimson:1986:STC

[289] Roger C. Grimson and Susan Groshen. A statistical test for contact among individuals in an epidemic and a pattern among cases of Acquired Immune Deficiency Syndrome. Statistics in Medicine, 5(3):271–279, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ewens:1986:CSP

[290] W. J. Ewens. Considerations of statistical power in the shared-haplotypes test. Statistics in Medicine, 5(3):281–288, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1986:BRS

[291] Douglas G. Altman. Book review: Summing up. The science of reviewing research, Richard J. Light and David B. Pillemer, Harvard University Press, Cambridge, Mass., 1984. No. of pages: x + 191. Price: £6.25 (paperback); £14.95 (hardback). Statistics in Medicine, 5(3):289, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Ac

[292] Anonymous. Announcement. Statistics in Medicine, 5(3):290, May–June 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Md

[293] Anonymous. Masthead. Statistics in Medicine, 5(4):fmi, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1986:GER

[294] E. A. Thompson. Genetic epidemiology: a review of the statistical basis. Statistics in Medicine, 5(4):291–302, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dyer:1986:MCR

[295] Alan R. Dyer. A method for combining results from several prospective epidemiologic studies. Statistics in Medicine, 5(4):303–317, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moussa:1986:ADC

[296] Mohamed A. A. Moussa. Allocation designs in cohort and case-control studies. Statistics in Medicine, 5(4):319–326, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1986:CEM

[297] Niels Keiding and Michael Væth. Calculating expected mortality. Statistics in Medicine, 5(4):327–334, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1986:ENR

[298] Toshiro Tango. Estimation of normal ranges of clinical laboratory data. Statistics in Medicine, 5(4):335–346, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kost:1986:DOS

[299] Gerald J. Kost. Discrimination and other statistical intervals for the interpretation of in vivo patient monitoring data. Statistics in Medicine, 5(4): 347–354, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

KaplanMD:1986:UBD

[300] Jonathan E. Kaplan, Lawrence B. Schonberger, and Kung-Jong Lui. The use of the binomial distribution in establishing an association between high-risk donors and transfusion-associated AIDS. Statistics in Medicine, 5(4):355–362, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mapleson:1986:UGB

[301] W. W. Mapleson. The use of glim and the bootstrap in assessing a clinical trial of two drugs. Statistics in Medicine, 5(4):363–374, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benichou:1986:SST

[302] Jacques Benichou and Claude Chastang. A simulation study of three sequential methods for the comparison of two treatment groups when the response criterion is censored. Statistics in Medicine, 5(4):375–385, July– August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:1986:CTI

[303] P. Y. Chi, D. R. Bristol, and J. V. Castellana. A clinical trial with an interim analysis. Statistics in Medicine, 5(4):387–392, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Le:1986:SCT

[304] Chap T. Le. Statistical comparison of two handwashing protocols. Statistics in Medicine, 5(4):393–396, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1986:BRD

[305] T. J. Cole. Book review: The development and exploitation of empirical birthweight standards, Roy Carr-Hill and Colin Pritchard, Macmillan Press, Basingstoke, U.K., 1985. No. of pages: xv + 195. Price: £40.00. Statistics in Medicine, 5(4):397, July–August 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Me

[306] Anonymous. Masthead. Statistics in Medicine, 5(5):fmi, September– October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1986:P

[307] Ted Colton, Laurence Freedman, and Tony Johnson. Preface. Statistics in Medicine, 5(5):399, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:ISC

[308] Anonymous. International Society for Clinical Biostatistics Sixth International Meeting held at the University of Düsseldorf from 16–20 September 1985. Statistics in Medicine, 5(5):401, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1986:APC

[309] G. D. Murray, L. S. Murray, P. Barlow, G. M. Teasdale, and W. B. Jennett. Assessing the performance and clinical impact of a computerized prognostic system in severe head injury. Statistics in Medicine, 5(5):403–410, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simes:1986:ASD

[310] R. J. Simes. Application of statistical decision theory to treatment choices: Implications for the design and analysis of clinical trials. Statistics in Medicine, 5(5):411–420, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:1986:PPP

[311] D. J. Spiegelhalter. Probabilistic prediction in patient management and clinical trials. Statistics in Medicine, 5(5):421–433, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moon:1986:CTC

[312] Thomas E. Moon. Clinical trials of cancer prevention agents: True versus observed prevention effect. Statistics in Medicine, 5(5):435–439, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herson:1986:CPI

[313] Jay Herson and Stephen K. Carter. Calibrated phase II clinical trials in oncology. Statistics in Medicine, 5(5):441–447, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turri:1986:DPU

[314] M. Turri and G. Stein. The determination of practically useful doses of new drugs: Some methodological considerations. Statistics in Medicine, 5(5):449–457, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1986:SSP

[315] John Whitehead. Sample sizes for phase II and phase III clinical trials: an integrated approach. Statistics in Medicine, 5(5):459–464, September– October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chastang:1986:PAI

[316] Claude Chastang, Philippe Travade, Jacques Benichou, Guillaume Dighiero, and Jacques-Louis Binet. Patient accrual and interim statistical analysis in long-term randomized clinical trials: The French chronic lymphocytic leukemia CLL 80 protocol as a case study. Statistics in Medicine, 5(5):465–473, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farley:1986:LTM

[317] T. M. M. Farley. Life-table methods for contraceptive research. Statistics in Medicine, 5(5):475–489, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mau:1986:CPA

[318] Jochen Mau and Berthold Steinke. A counting process approach to the analysis of the course of non-Hodgkin’s lymphoma. Statistics in Medicine, 5(5):491–495, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McSherry:1986:IDB

[319] D. M. G. McSherry. Intelligent dialogue based on statistical models of clinical decision-making. Statistics in Medicine, 5(5):497–502, September– October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohmann:1986:SVU

[320] C. Ohmann, M. Künneke, R. Zaczyk, K. Thon, and W. Lorenz. Selection of variables using ‘independence Bayes’ in computer-aided diagnosis of upper gastrointestinal bleeding. Statistics in Medicine, 5(5):503–515, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1986:PMC

[321] R. J. Marshall. Partitioning methods for classification and decision making in medicine. Statistics in Medicine, 5(5):517–526, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:1986:APC

[322] Chris Robertson and Peter Boyle. Age, period and cohort models: The use of individual records. Statistics in Medicine, 5(5):527–538, September– October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Darby:1986:SRS

[323] Sarah C. Darby. Some recent statistical analyses of two long-term studies of exposure to ionizing radiation. Statistics in Medicine, 5(5):539–546, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wahrendorf:1986:CFC

[324] Jürgen Wahrendorf. The changing face of cancer epidemiology. Statistics in Medicine, 5(5):547–553, September–October 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:Mf

[325] Anonymous. Masthead. Statistics in Medicine, 5(6):fmi, November– December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gladen:1986:RHA

[326] Beth C. Gladen. On the role of ‘habitual aborters’ in the analysis of spontaneous abortion. Statistics in Medicine, 5(6):557–564, November– December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gallus:1986:SMC

[327] G. Gallus, C. Mandelli, M. Marchi, and G. Radaelli. On surveillance methods for congenital malformations. Statistics in Medicine, 5(6):565– 571, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Podgor:1986:EIA

[328] Marvin J. Podgor and M. Cristina Leske. Estimating incidence from age-specific prevalence for irreversible diseases with differential mortality. Statistics in Medicine, 5(6):573–578, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1986:ARR

[329] Sander Greenland. Adjustment of risk ratios in case-base studies (hybrid epidemiologic designs). Statistics in Medicine, 5(6):579–584, November– December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taube:1986:SSP

[330] Adam Taube. Sensitivity, specificity and predictive values: a graphical approach. Statistics in Medicine, 5(6):585–591, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsouros:1986:ATS

[331] Agis D. Tsouros and Robert J. Young. Applications of time-series analysis: a case study on the impact of computer tomography. Statistics in Medicine, 5(6):593–606, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cartwright:1986:SEF

[332] Ann Cartwright. Some experiments with factors that might affect the response of mothers to a postal questionnaire. Statistics in Medicine, 5 (6):607–617, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hertzberg:1986:MVA

[333] Vicki Stover Hertzberg and Lloyd D. Fisher. A model for variability in arteriographic reading. Statistics in Medicine, 5(6):619–627, November– December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dupont:1986:SFE

[334] William D. Dupont. Sensitivity of Fisher’s exact test to minor perturbations in 2 × 2 contingency tables. Statistics in Medicine, 5(6):629–635, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Umen:1986:PFM

[335] Andrew J. Umen and Chap T. Le. Prognostic factors, models and related statistical problems in the survival of end-stage renal disease patients on hemodialysis. Statistics in Medicine, 5(6):637–652, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Afifi:1986:SNP

[336] A. A. Afifi, R. M. Elashoff, and J. J. Lee. Simultaneous non-parametric confidence intervals for survival probabilities from censored data. Statistics in Medicine, 5(6):653–662, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bloch:1986:SSR

[337] Daniel A. Bloch. Sample size requirements and the cost of a randomized clinical trial with repeated measurements. Statistics in Medicine, 5(6): 663–667, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1986:EPA

[338] Anonymous. On expert pre-assessment of prospects in clinical trials. Statistics in Medicine, 5(6):669, November–December 1986. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Ma

[339] Anonymous. Masthead. Statistics in Medicine, 6(1):fmi, January– February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1987:Pa

[340] Ted Colton, Laurence Freedman, and Tony Johnson. Preface. Statistics in Medicine, 6(1):1–2, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wulff:1987:WDD

[341] Henrik R. Wulff, Bjørn Andersen, Preben Brandenhoff, and Flemming Guttler. What do doctors know about statistics? Statistics in Medicine, 6(1):3–10, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simes:1987:CPB

[342] Robert John Simes. Confronting publication bias: a cohort design for meta-analysis. Statistics in Medicine, 6(1):11–29, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:1987:ECM

[343] Timothy M. Morgan and Robert M. Elashoff. Effect of covariate measurement error in randomized clinical trials. Statistics in Medicine, 6 (1):31–41, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1987:ADA

[344] Allan Donner and Alan Donald. Analysis of data arising from a stratified design with the cluster as unit of randomization. Statistics in Medicine, 6 (1):43–52, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Connett:1987:SSP

[345] John E. Connett, Judith A. Smith, and Richard B. McHugh. Sample size and power for pair-matched case-control studies. Statistics in Medicine, 6 (1):53–59, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bean:1987:SSB

[346] Judy A. Bean, Curtis G. Wiltse, and Robert E. Woolson. Small sample behaviour of hypothesis tests related to indirect standardized rates: a Monte Carlo study. Statistics in Medicine, 6(1):61–70, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockette:1987:EPM

[347] Howard E. Rockette and Vincent C. Arena. Evaluation of the proportionate mortality index in the presence of multiple comparisons. Statistics in Medicine, 6(1):71–77, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heeren:1987:RTI

[348] Timothy Heeren and Ralph D’Agostino. Robustness of the two independent samples t-test when applied to ordinal scaled data. Statistics in Medicine, 6(1):79–90, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1987:MLE

[349] W. J. Shih. Maximum likelihood estimation and likelihood ratio test for square tables with missing data. Statistics in Medicine, 6(1):91–97, January–February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Aa

[350] Anonymous. Announcements. Statistics in Medicine, 6(1):99, January– February 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Mb

[351] Anonymous. Masthead. Statistics in Medicine, 6(2):fmi, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:1987:TDP

[352] John D. Emerson and David Tritchler. The three-decision problem in medical decision making. Statistics in Medicine, 6(2):101–112, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curnow:1987:CRA

[353] R. N. Curnow. Correcting for regression in assessing the response to treatment in a selected population. Statistics in Medicine, 6(2):113–117, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sposto:1987:UUA

[354] Richard Sposto and Mark D. Krailo. Use of unequal allocation in survival trials. Statistics in Medicine, 6(2):119–125, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abel:1987:MRR

[355] U. Abel. Modified replacement randomization. Statistics in Medicine, 6 (2):127–135, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wastell:1987:NAC

[356] D. G. Wastell and R. Gray. The numerical approach to classification: a medical application to develop a typology for facial pain. Statistics in Medicine, 6(2):137–146, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Linnet:1987:CQD

[357] Kristian Linnet. Comparison of quantitative diagnostic tests: Type I error, power, and sample size. Statistics in Medicine, 6(2):147–158, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spicer:1987:SAH

[358] C. C. Spicer, C. J. Lawrence, and D. P. Southall. Statistical analysis of heart rates in subsequent victims of sudden infant death syndrome. Statistics in Medicine, 6(2):159–166, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1987:FNM

[359] Joel B. Greenhouse, Robert E. Kass, and Ruey S. Tsay. Fitting nonlinear models with ARMA errors to biological rhythm data. Statistics in Medicine, 6(2):167–183, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karlberg:1987:MHG

[360] Johan Karlberg. On the modelling of human growth. Statistics in Medicine, 6(2):185–192, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1987:BIA

[361] Sander Greenland. Bias in indirectly adjusted comparisons due to taking the total study population as the reference group. Statistics in Medicine, 6(2):193–195, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wallenstein:1987:ADS

[362] Sylvan Wallenstein and Norman Neff. An approximation for the distribution of the scan statistic. Statistics in Medicine, 6(2):197–207, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1987:BRT

[363] Sander Greenland. Book review: Theoretical epidemiology, Olli S. Miettinen, Wiley, New York, 1985. No. of pages: 359, Price: $32.95. Statistics in Medicine, 6(2):209–211, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dendy:1987:BRM

[364] P. P. Dendy. Book review: Mathematical methods in medicine. Part II. Applications in clinical specialties, edited by David Ingram and Ralph Bloch, Wiley, Chichester, 1986. No. of pages: 533. Price: £29.95. Statistics in Medicine, 6(2):211–212, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittkowski:1987:BRS

[365] Knut M. Wittkowski. Book review: Statistik-software: 3. Konferenz über die wissenschaftliche Anwendung von Statistik-software. Walter Lehmacher and Allmut Hörmann (eds), Gustav Fischer Verlag, Stuttgart, 1985. No. of pages: 402. DM 58.00. Statistics in Medicine, 6(2):212, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houghton:1987:SMQ

[366] J. Houghton and R. A’Hern. Symposium on methodology and quality assurance in cancer clinical trials: Cancer treatment reports, number 10, 1039–1233. Statistics in Medicine, 6(2):213, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Ab

[367] Anonymous. Announcements. Statistics in Medicine, 6(2):215, March 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Mc

[368] Anonymous. Masthead. Statistics in Medicine, 6(3):fmi, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yusuf:1987:P

[369] Salim Yusuf, Richard Simon, and Susan S. Ellenberg. Preface. Statistics in Medicine, 6(3):217–218, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:LSD

[370] Anonymous. List of speakers and discussants. Statistics in Medicine, 6(3): 219–220, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Light:1987:AEI

[371] Richard J. Light. Accumulating evidence from independent studies: What we can win and what we can lose. Statistics in Medicine, 6(3):221–228, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Da

[372] Anonymous. Discussion. Statistics in Medicine, 6(3):229–231, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peto:1987:WDW

[373] Richard Peto. Why do we need systematic overviews of randomized trials? (transcript of an oral presentation, modified by the editors). Statistics in Medicine, 6(3):233–240, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Db

[374] Anonymous. Discussion. Statistics in Medicine, 6(3):241–244, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:1987:ALB

[375] Rory Collins, Richard Gray, Jon Godwin, and Richard Peto. Avoidance of large biases and large random errors in the assessment of moderate treatment effects: The need for systematic overviews. Statistics in Medicine, 6(3):245–250, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dc

[376] Anonymous. Discussion. Statistics in Medicine, 6(3):251–254, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Canner:1987:OSC

[377] Paul L. Canner. An overview of six clinical trials of aspirin in coronary heart disease. Statistics in Medicine, 6(3):255–263, April–May 1987. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dd

[378] Anonymous. Discussion. Statistics in Medicine, 6(3):265–267, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1987:PMI

[379] Robert E. Wittes. Problems in the medical interpretation of overviews. Statistics in Medicine, 6(3):269–276, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:De

[380] Anonymous. Discussion. Statistics in Medicine, 6(3):277–280, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yusuf:1987:OMM

[381] Salim Yusuf. Obtaining medically meaningful answers from an overview of randomized clinical trials. Statistics in Medicine, 6(3):281–286, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Df

[382] Anonymous. Discussion. Statistics in Medicine, 6(3):287–294, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Furberg:1987:LOC

[383] Curt D. Furberg and Timothy M. Morgan. Lessons from overviews of cardiovascular trials. Statistics in Medicine, 6(3):295–303, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dg

[384] Anonymous. Discussion. Statistics in Medicine, 6(3):305–306, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedman:1987:PPO

[385] Michael A. Friedman. Potential pooling opportunities: Cancer. Statistics in Medicine, 6(3):307–312, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dh

[386] Anonymous. Discussion. Statistics in Medicine, 6(3):313–314, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chalmers:1987:MACa

[387] Thomas C. Chalmers, Howard Levin, Henry S. Sacks, Dinah Reitman, Jayne Berrier, and Raguraman Nagalingam. Meta-analysis of clinical trials as a scientific discipline. i: Control of bias and comparison with large cooperative trials. Statistics in Medicine, 6(3):315–325, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Di

[388] Anonymous. Discussion. Statistics in Medicine, 6(3):327–328, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meier:1987:C

[389] Paul Meier. Commentary. Statistics in Medicine, 6(3):329–331, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Norton:1987:C

[390] Larry Norton. Commentary. Statistics in Medicine, 6(3):333–335, April– May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dj

[391] Anonymous. Discussion. Statistics in Medicine, 6(3):337–340, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1987:MCR

[392] David L. Demets. Methods for combining randomized clinical trials: Strengths and limitations. Statistics in Medicine, 6(3):341–348, April– May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dk

[393] Anonymous. Discussion. Statistics in Medicine, 6(3):349–350, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:1987:ISD

[394] Kent R. Bailey. Inter-study differences: How should they influence the interpretation and analysis of results? Statistics in Medicine, 6(3):351– 358, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dl

[395] Anonymous. Discussion. Statistics in Medicine, 6(3):359–360, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1987:EOE

[396] Stephanie J. Green, Thomas R. Fleming, and Scott Emerson. Effects on overviews of early stopping rules for clinical trials. Statistics in Medicine, 6(3):361–367, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dm

[397] Anonymous. Discussion. Statistics in Medicine, 6(3):369, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelber:1987:IRS

[398] Richard D. Gelber and Aron Goldhirsch. Interpretation of results from subset analyses within overviews of randomized clinical trials. Statistics in Medicine, 6(3):371–378, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dn

[399] Anonymous. Discussion. Statistics in Medicine, 6(3):379–380, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedges:1987:C

[400] Larry V. Hedges. Commentary. Statistics in Medicine, 6(3):381–385, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Do

[401] Anonymous. Discussion. Statistics in Medicine, 6(3):387–388, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:1987:ROC

[402] Richard Simon. The role of overviews in cancer therapeutics. Statistics in Medicine, 6(3):389–393, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dp

[403] Anonymous. Discussion. Statistics in Medicine, 6(3):395–396, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hennekens:1987:IOR

[404] Charles H. Hennekens, Julie E. Buring, and Patricia R. Hebert. Implications of overviews of randomized trials. Statistics in Medicine, 6(3): 397–402, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Dq

[405] Anonymous. Discussion. Statistics in Medicine, 6(3):403–409, April–May 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Md

[406] Anonymous. Masthead. Statistics in Medicine, 6(4):fmi, June 1987. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1987:BAD

[407] Colin B. Begg. Biases in the assessment of diagnostic tests. Statistics in Medicine, 6(4):411–423, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nix:1987:IQC

[408] A. B. J. Nix, R. J. Rowlands, K. W. Kemp, D. W. Wilson, and K. Griffiths. Internal quality control in clinical chemistry: a teaching review. Statistics in Medicine, 6(4):425–440, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1987:SSR

[409] Allan Donner and Michael Eliasziw. Sample size requirements for reliability studies. Statistics in Medicine, 6(4):441–448, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayton:1987:MTVa

[410] D. Clayton and E. Schifflers. Models for temporal variation in cancer rates. i: Age-period and age-cohort models. Statistics in Medicine, 6(4):

449–467, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayton:1987:MTVb

[411] D. Clayton and E. Schifflers. Models for temporal variation in cancer rates. II: Age-period-cohort models. Statistics in Medicine, 6(4):469–481, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halpern:1987:CRM

[412] Jerry Halpern and Byron Wm. Brown, Jr. Cure rate models: Power of the logrank and generalized Wilcoxon tests. Statistics in Medicine, 6(4):483– 489, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donald:1987:AMH

[413] Alan Donald and Allan Donner. Adjustments to the Mantel–Haenszel chi-square statistic and odds ratio variance estimator when the data are clustered. Statistics in Medicine, 6(4):491–499, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mulder:1987:ATM

[414] Paul G. H. Mulder and Anton L. Hempenius. Analysis of total mortality in the Rotterdam sample of the Kaunas–Rotterdam intervention study (KRIS). Statistics in Medicine, 6(4):501–516, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1987:RES

[415] Rina Chen. The relative efficiency of the sets and the cusum techniques in monitoring the occurrence of a rare event. Statistics in Medicine, 6 (4):517–525, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newman:1987:FCD

[416] Stephen C. Newman. Formulae for cause-deleted life tables. Statistics in Medicine, 6(4):527–528, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutchinson:1987:VLC

[417] T. P. Hutchinson, K. Y. Tang, and D. Rindskopf. The value of latent class analysis in medical diagnosis. Statistics in Medicine, 6(4):529–531, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Ac

[418] Anonymous. Announcement. Statistics in Medicine, 6(4):533, June 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Me

[419] Anonymous. Masthead. Statistics in Medicine, 6(5):fmi, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:C

[420] Anonymous. Correction. Statistics in Medicine, 6(5):i, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shewmon:1987:PII

[421] D. Alan Shewmon. The probability of inevitability: The inherent impossibility of validating criteria for brain death or ‘irreversibility’ through clinical studies. Statistics in Medicine, 6(5):535–553, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1987:MBD

[422] Byron Jones and Michael G. Kenward. Modelling binary data from a three-period cross-over trial. Statistics in Medicine, 6(5):555–564, July– August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1987:IEC

[423] Marc Buyse and Louise M. Ryan. Issues of efficiency in combining proportions of deaths from several clinical trials. Statistics in Medicine, 6(5): 565–576, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1987:SMS

[424] F. Y. Hsieh. A simple method of sample size calculation for unequalsample-size designs that use the logrank or t-test. Statistics in Medicine, 6(5):577–581, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang:1987:ATS

[425] Christy Chuang. The analysis of a titration study. Statistics in Medicine, 6(5):583–590, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dixon:1987:MCR

[426] Dennis O. Dixon and George W. Divine. Multiple comparisons for relative risk regression: Extension of the k-ratio method. Statistics in Medicine, 6(5):591–597, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palta:1987:SCA

[427] Mari Palta and Thomas Cook. Some considerations in the analysis of rates of change in longitudinal studies. Statistics in Medicine, 6(5):599– 611, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martin:1987:MMS

[428] Claudia J. Martin. Monitoring maternity services by postal questionnaire: Congruity between mothers’ reports and their obstetric records. Statistics in Medicine, 6(5):613–627, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chakraborty:1987:NRS

[429] Ranajit Chakraborty and Craig L. Hanis. Non-random sampling in human genetics: Estimation of familial correlations, model testing, and interpretation. Statistics in Medicine, 6(5):629–646, July–August 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Mf

[430] Anonymous. Masthead. Statistics in Medicine, 6(6):fmi, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eliasziw:1987:CFA

[431] Michael Eliasziw and Allan Donner. A cost-function approach to the design of reliability studies. Statistics in Medicine, 6(6):647–655, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:1987:EEC

[432] Steven M. Snapinn. Evaluating the efficacy of a combination therapy. Statistics in Medicine, 6(6):657–665, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gebski:1987:MDA

[433] Val Gebski, Don McNeil, Alan Coates, and John Forbes. Monitoring distributional assumptions and early stopping for a prospective clinical trial using Monte Carlo simulation. Statistics in Medicine, 6(6):667–678, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dorey:1987:ESS

[434] Frederick J. Dorey and Edward L. Korn. Effective sample sizes for confidence intervals for survival probabilities. Statistics in Medicine, 6(6): 679–687, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Niland:1987:HTT

[435] Joyce C. Niland, Duncan Thomas, and Stanley P. Azen. Hypothesis testing of time-dependent recurrent events. Statistics in Medicine, 6(6):689– 699, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1987:VEA

[436] Sander Greenland. Variance estimators for attributable fraction estimates consistent in both large strata and sparse data. Statistics in Medicine, 6(6):701–708, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:1987:APC

[437] Toshiro Tang and Shiusuke Kurashina. Age, period and cohort analysis of trends in mortality from major diseases in Japan, 1955 to 1979: Peculiarity of the cohort born in the early Showa Era. Statistics in Medicine, 6(6): 709–726, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1987:CRA

[438] S. J. Senn and R. N. Curnow. Correcting for regression in assessing the response to treatment in a selected population. Statistics in Medicine, 6(6): 727–728, September 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Mg

[439] Anonymous. Masthead. Statistics in Medicine, 6(7):fmi, October– November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1987:Pb

[440] Ted Colton, Laurence Freedman, and Tony Johnson. Preface. Statistics in Medicine, 6(7):729, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:ISC

[441] Anonymous. International Society for Clinical Biostatistics. Seventh International Meeting held at University College, Cardiff from 15–19 September 1986. Statistics in Medicine, 6(7):731, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chalmers:1987:MACb

[442] Thomas C. Chalmers, Jayne Berrier, Henry S. Sacks, Howard Levin, Dinah Reitman, and Raguraman Nagalingam. Meta-analysis of clinical trials as a scientific discipline. II: Replicate variability and comparison of studies that agree and disagree. Statistics in Medicine, 6(7):733–744, October– November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1987:RCT

[443] Jørgen Hilden. Reporting clinical trials from the viewpoint of a patient’s choice of treatment. Statistics in Medicine, 6(7):745–752, October– November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1987:PAD

[444] David L. Demets. Practical aspects in data monitoring: a brief review. Statistics in Medicine, 6(7):753–760, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anthony:1987:SMP

[445] Honor M. Anthony. Some methodological problems in the assessment of complementary therapy. Statistics in Medicine, 6(7):761–771, October– November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumacher:1987:IHC

[446] Martin Schumacher, Manfred Olschewski, and Claudia Schmoor. The impact of heterogeneity on the comparison of survival times. Statistics in Medicine, 6(7):773–784, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kramar:1987:MAT

[447] Andrew Kramar, Marie-Hélène Pejovic, and Daniel Chassagne. A method of analysis taking into account competing events: Application to the study of digestive complications following irradiation for cervical cancer. Statistics in Medicine, 6(7):785–794, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jais:1987:TSM

[448] Jean-Philippe Jais, Claude Chastang, and Jean-Claude Raphael. A transient state model for analysing ventilatory status in a randomized clinical trial. Statistics in Medicine, 6(7):795–803, October–November 1987. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crichton:1987:PBA

[449] Nicola J. Crichton and Peter A. Emerson. A probability-based aid for teaching medical students a logical approach to diagnosis. Statistics in Medicine, 6(7):805–811, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charpak:1987:DSE

[450] Yves Charpak, Christian Bléry, and Claude Chastang. Designing a study for evaluating a protocol for the selective performance of preoperative tests. Statistics in Medicine, 6(7):813–822, October–November 1987. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Busse:1987:FDC

[451] H. Busse, E. Bergmann, and K. Bergmann. Fluoride and dental caries: Two different statistical approaches to the same data source. Statistics in Medicine, 6(7):823–842, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeBoo:1987:PDE

[452] Th. M. De Boo, J. A. M. Van Druten, and A. D. Plantinga. Predicting the dynamic effects of rubella vaccination programmes. Statistics in Medicine, 6(7):843–851, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collett:1987:MEI

[453] D. Collett and M. S. Lye. Modelling the effect of intervention on the transmission of malaria in East Malaysia. Statistics in Medicine, 6(7): 853–861, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpson:1987:UAR

[454] Judy M. Simpson, D. N. Bateman, and M. D. Rawlins. Using the adverse reactions register to study the effects of age and sex on adverse drug reactions. Statistics in Medicine, 6(7):863–867, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1987:CHS

[455] P. A. Lewis and M. Charny. The Cardiff Health Survey: Teaching survey methodology by participation. Statistics in Medicine, 6(7):869–874, October–November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:1987:BRS

[456] J. M. Bland. Book review: The statistical consultant in action, edited by D. J. Hand and B. S. Everitt, Cambridge University Press, 1987. No. of pages: 189, Price: £20 ($34.50). Statistics in Medicine, 6(7):875, October– November 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1987:Mh

[457] Anonymous. Masthead. Statistics in Medicine, 6(8):fmi, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1987:Pc

[458] Ted Colton, Laurence Freedman, and Tony Johnson. Preface. Statistics in Medicine, 6(8):877, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeStavola:1987:SFA

[459] Biancaw De Stavola. Statistical facts about cancers on which Doctor Rigoni-Stern based his contribution to the Surgeons’ Subgroup of the IV Congress of the Italian Scientists on 23 September 1842. Statistics in Medicine, 6(8):881–884, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1987:ESC

[460] S. D. Walter and L. W. Stitt. Evaluating the survival of cancer cases detected by screening. Statistics in Medicine, 6(8):885–900, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bell:1987:MHW

[461] C. M. J. Bell and D. A. Coleman. Models of the healthy worker effect in industrial cohorts. Statistics in Medicine, 6(8):901–909, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenward:1987:ADC

[462] Michael G. Kenward and Byron Jones. The analysis of data from 2 × 2 cross-over trials with baseline measurements. Statistics in Medicine, 6(8): 911–926, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanEs:1987:ETC

[463] Gerrit-Anne Van Es, Jan G. P. Tijssen, Jacobus Lubsen, and Roel van Strik. Early termination of clinical trials with prolonged observation of individual patients: a case study. Statistics in Medicine, 6(8):927–937, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1987:CAS

[464] Niels Keiding, Thomas Bayer, and Steen Watt-Boolsen. Confirmatory analysis of survival data using left truncation of the life times of primary survivors. Statistics in Medicine, 6(8):939–944, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crichton:1987:SPU

[465] Nicola J. Crichton, John G. Fryer, and Clive C. Spicer. Some points on the use of ‘independent Bayes’ to diagnose acute abdominal pain. Statistics in Medicine, 6(8):945–959, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franti:1987:IED

[466] C. E. Franti, A. D. Wiggins, and J. F. Kraus. Improved estimates of discordance ratios in twin studies. Statistics in Medicine, 6(8):961–970, December 1987. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Ma

[467] Anonymous. Masthead. Statistics in Medicine, 7(1-2):fmi, January– February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1988:E

[468] Ted Colton, Laurence Freedman, and Tony Johnson. Editorial. Statistics in Medicine, 7(1-2):i–ii, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:P

[469] Anonymous. Preface. Statistics in Medicine, 7(1-2):1–2, January– February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:LSD

[470] Anonymous. List of speakers and discussants. Statistics in Medicine, 7 (1-2):3–5, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breslow:1988:CI

[471] Norman Breslow. Chairman’s introduction. Statistics in Medicine, 7(1-2): 7, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buist:1988:ULF

[472] A. Sonia Buist and William M. Vollmer. The use of lung function tests in identifying factors that affect lung growth and aging. Statistics in Medicine, 7(1-2):11–18, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diem:1988:CST

[473] John E. Diem and John R. Liukkonen. A comparative study of three methods for analysing longitudinal pulmonary function data. Statistics in Medicine, 7(1-2):19–28, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:1988:GMA

[474] Thomas A. Louis. General methods for analysing repeated measures. Statistics in Medicine, 7(1-2):29–45, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lauer:1988:LVB

[475] Ronald M. Lauer and William R. Clarke. A longitudinal view of blood pressure during childhood: The Muscatine Study. Statistics in Medicine, 7(1-2):47–57, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1988:AMA

[476] B. Rosner and A. Muñoz. Autoregressive modelling for the analysis of longitudinal data with unequally spaced examinations. Statistics in Medicine, 7(1-2):59–71, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1988:AGC

[477] P. David Wilson. Autoregressive growth curves and Kalman filtering. Statistics in Medicine, 7(1-2):73–86, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:DMA

[478] Anonymous. Discussion. methods for analysing repeated measurements. Statistics in Medicine, 7(1-2):87–92, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ware:1988:IAR

[479] James H. Ware, Stuart Lipsitz, and Frank E. Speizer. Issues in the analysis of repeated categorical outcomes. Statistics in Medicine, 7(1-2):95–107, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landis:1988:SGM

[480] J. Richard Landis, Michael E. Miller, Charles S. Davis, and Gary G. Koch. Some general methods for the analysis of categorical data in longitudinal studies. Statistics in Medicine, 7(1-2):109–137, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:1988:ARM

[481] L. J. Wei and D. O. Stram. Analysing repeated measurements with possibly missing observations by modelling marginal distributions. Statistics in Medicine, 7(1-2):139–148, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalbfleisch:1988:LAM

[482] J. D. Kalbfleisch and J. F. Lawless. Likelihood analysis of multi-state models for disease incidence and mortality. Statistics in Medicine, 7(1-2): 149–160, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeger:1988:C

[483] Scott L. Zeger. Commentary. Statistics in Medicine, 7(1-2):161–168, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:DRC

[484] Anonymous. Discussion. repeated categorical response. Statistics in Medicine, 7(1-2):169–173, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dwyer:1988:RWS

[485] James H. Dwyer and Manning Feinleib. Report of a Workshop on Statistical Issues in Longitudinal Studies of Health Held in Berlin (West), 23– 26 June 1986. Statistics in Medicine, 7(1-2):177–183, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Province:1988:FAP

[486] M. A. Province and D. C. Rao. Familial aggregation in the presence of temporal trends. Statistics in Medicine, 7(1-2):185–198, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:DST

[487] Anonymous. Discussion. special topics. Statistics in Medicine, 7(1-2): 199–202, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cupples:1988:CBR

[488] L. Adrienne Cupples, Ralph B. D’Agostino, Keaven Anderson, and William B. Kannel. Comparison of baseline and repeated measure covariate techniques in the Framingham heart study. Statistics in Medicine, 7(1-2):205–218, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abbott:1988:CCB

[489] Robert D. Abbott. Commentary on ‘Comparison of baseline and repeated measure covariate techniques in the Framingham heart study’. Statistics in Medicine, 7(1-2):219–222, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittemore:1988:ECS

[490] Alice S. Whittemore. Effect of cigarette smoking in epidemiological studies of lung cancer. Statistics in Medicine, 7(1-2):223–238, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manton:1988:MIM

[491] Kenneth G. Manton, Max A. Woodbury, and Eric Stallard. Models of the interaction of mortality and the evolution of risk factor distribution: a general stochastic process formulation. Statistics in Medicine, 7(1-2): 239–256, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lagakos:1988:EMM

[492] S. W. Lagakos. Effects of mismodelling and mismeasuring explanatory variables on tests of their association with a response variable. Statistics in Medicine, 7(1-2):257–274, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prentice:1988:AUR

[493] Ross L. Prentice and Steven G. Self. Aspects of the use of relative risk models in the design and analysis of cohort studies and prevention trials. Statistics in Medicine, 7(1-2):275–287, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halperin:1988:C

[494] Max Halperin. Commentary. Statistics in Medicine, 7(1-2):289–292, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manton:1988:CRD

[495] Kenneth G. Manton, Max A. Woodbury, and Eric Stallard. Commentary. Response to Dr Halperin on ‘Models of the interaction of mortality and the evolution of risk factor distribution: a general stochastic process formulation’. Statistics in Medicine, 7(1-2):293–295, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:DRF

[496] Anonymous. Discussion. risk factors and survival. Statistics in Medicine, 7(1-2):297–302, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laird:1988:MDL

[497] Nan M. Laird. Missing data in longitudinal studies. Statistics in Medicine, 7(1-2):305–315, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluchter:1988:AIM

[498] Mark D. Schluchter. Analysis of incomplete multivariate data using linear models with structured covariance matrices. Statistics in Medicine, 7(1-2):317–324, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woodbury:1988:EHM

[499] Max A. Woodbury, Kenneth G. Manton, and Anatoli I. Yashin. Estimating hidden morbidity via its effect on mortality and disability. Statistics in Medicine, 7(1-2):325–336, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:1988:ACP

[500] Margaret C. Wu and Kent Bailey. Analysing changes in the presence of informative right censoring caused by death and withdrawal. Statistics in Medicine, 7(1-2):337–346, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:1988:C

[501] Roderick J. A. Little. Commentary. Statistics in Medicine, 7(1-2):347– 355, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:DAD

[502] Anonymous. Discussion. analysis of data with missing values. Statistics in Medicine, 7(1-2):357–360, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breslow:1988:CCR

[503] Norman Breslow. Chairman’s concluding remarks. Statistics in Medicine, 7(1-2):361, January–February 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mb

[504] Anonymous. Masthead. Statistics in Medicine, 7(3):fmi, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lubin:1988:SSP

[505] Jay H. Lubin, Mitchell H. Gail, and Abby G. Ershow. Sample size and power for case-control studies when exposures are continuous. Statistics in Medicine, 7(3):363–376, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Merkouriou:1988:ERR

[506] Savvas Merkouriou and Douglas Dix. Estimating reference ranges in clinical pathology: an objective approach. Statistics in Medicine, 7(3):377– 385, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpkin:1988:TCP

[507] J. M. Simpkin and D. Y. Downham. Testing for a change-point in registry data with an example on hypospadias. Statistics in Medicine, 7(3):387– 393, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1988:MBA

[508] Kung-Jong Lui, Thomas A. Peterman, Dale N. Lawrence, and James R. Allen. A model-based approach to characterize the incubation period of paediatric transfusion-associated acquired immunodeficiency syndrome. Statistics in Medicine, 7(3):395–401, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Espeland:1988:ADR

[509] Mark A. Espeland, William C. Murphy, and Dennis H. Leverett. Assessing diagnostic reliability and estimating incidence rates associated with a strictly progressive disease: Dental caries. Statistics in Medicine, 7(3): 403–416, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1988:CSB

[510] P. David Wilson. Correcting for selection bias in estimation of withinindividual variance. Statistics in Medicine, 7(3):417–422, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1988:PLT

[511] P. David Wilson and J. Richard Hebel. The probability that long-term mean diastolic blood pressure is extreme: Tables for black individuals. Statistics in Medicine, 7(3):423–433, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1988:MRM

[512] Christopher Cox. Multinomial regression models based on continuation ratios. Statistics in Medicine, 7(3):435–441, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1988:IAN

[513] Jack Green, Mandouh Montasser, Heng Chin Low, and J. C. Woodrow. Investigation of the associations of a number of HLA antigens with psoriasis and psoriatic arthritis. Statistics in Medicine, 7(3):443–450, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:1988:CTI

[514] P. Y. Chi, D. R. Bristol, and J. V. Castellana. A clinical trial with an interim analysis. Statistics in Medicine, 7(3):451, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Aa

[515] Anonymous. Announcement. Statistics in Medicine, 7(3):452, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1988:BRL

[516] David R. Jones. Book review: Living with risk: The BMA guide. Wiley, Chichester, 1987. Number of pages: 192; Price: £14.95/$29.60. Statistics in Medicine, 7(3):453–454, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1988:BRD

[517] Per Kragh Andersen. Book review: The design and analysis of clinical experiments, J. L. Fleiss, Wiley, Chichester, 1986. Number of pages: xiv + 432, price: £38.35/$55.00. Statistics in Medicine, 7(3):454, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1988:BRS

[518] M. J. Campbell. Book review: Statistical designs for research, by Leslie Kish, Wiley, New York, 1987. No. of pages: xii + 267; Price: £33.70 ($48.95). Statistics in Medicine, 7(3):455, March 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mc

[519] Anonymous. Masthead. Statistics in Medicine, 7(4):fmi, April 1988. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blumenson:1988:LPO

[520] Leslie E. Blumenson. Loss of power from an optimistic alternative hypothesis. Statistics in Medicine, 7(4):457–466, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1988:NSS

[521] Peter A. Lachenbruch. A note on sample size computation for testing interactions. Statistics in Medicine, 7(4):467–469, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hafner:1988:SAS

[522] Kerry B. Hafner, Gary G. Koch, and Andrew T. Canada. Some analysis strategies for three-period changeover designs with two treatments. Statistics in Medicine, 7(4):471–481, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tallis:1988:ASD

[523] G. M. Tallis, P. Leppard, and T. J. O’Neill. The analysis of survival data from a central cancer registry with passive follow-up. Statistics in Medicine, 7(4):483–490, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schulman:1988:DEM

[524] J. Schulman, S. Selvin, and D. W. Merrill. Density equalized map projections: a method for analysing clustering around a fixed point. Statistics in Medicine, 7(4):491–505, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuritz:1988:SAR

[525] Stephen J. Kuritz and J. Richard Landis. Summary attributable risk estimation from unmatched case-control data. Statistics in Medicine, 7 (4):507–517, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1988:EDP

[526] Manfred S. Green. Evaluating the discriminatory power of a multiple logistic regression model. Statistics in Medicine, 7(4):519–524, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:1988:UML

[527] E. Lesaffre and A. Albert. An uncertainty measure in logistic discrimination. Statistics in Medicine, 7(4):525–533, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hamilton:1988:DSA

[528] Martin A. Hamilton and Michael G. Rinaldi. Descriptive statistical analyses of serial dilution data. Statistics in Medicine, 7(4):535–544, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaplin:1988:BRC

[529] Caroline Chaplin. Book review: Clinical trials: Design, conduct and analysis, Curtis L. Meinert, Oxford University Press, 1986. No. of pages xxvi + 469. price: £50.00. Statistics in Medicine, 7(4):545, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Ab

[530] Anonymous. Announcements. Statistics in Medicine, 7(4):547, April 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Md

[531] Anonymous. Masthead. Statistics in Medicine, 7(5):fmi, May 1988. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1988:MRP

[532] Peter A. Lachenbruch. Multiple reading procedures: The performance of diagnostic tests. Statistics in Medicine, 7(5):549–557, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moussa:1988:PSS

[533] Mohamed A. A. Moussa. Planning the size of survival time clinical trials with allowance for stratification. Statistics in Medicine, 7(5):559–569, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1988:OTS

[534] Peter F. Thall, Richard Simon, Susan S. Ellenberg, and Richard Shrager. Optimal two-stage designs for clinical trials with binary response. Statistics in Medicine, 7(5):571–579, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hopper:1988:RWM

[535] John L. Hopper and Graeme P. Young. A random walk model for evaluating clinical trials involving serial observations. Statistics in Medicine, 7(5):581–590, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Khamis:1988:CTE

[536] H. J. Khamis. Contingency table estimation of genetic parameters and disease risks. Statistics in Medicine, 7(5):591–600, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segal:1988:SME

[537] Mark R. Segal, Scott T. Weiss, Frank E. Speizer, and Ira B. Tager. Smoothing methods for epidemiologic analysis. Statistics in Medicine, 7(5):601–611, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobson:1988:PHM

[538] Annette J. Dobson. Proportional hazards models for average data for groups. Statistics in Medicine, 7(5):613–618, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1988:PES

[539] John Whitehead and Cameron Dorse. Prospective epidemiological studies involving paired organs. Statistics in Medicine, 7(5):619–625, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Ac

[540] Anonymous. Announcement. Statistics in Medicine, 7(5):627, May 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Me

[541] Anonymous. Masthead. Statistics in Medicine, 7(6):fmi, June 1988. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1988:HET

[542] Douglas G. Altman and J. Patrick Royston. The hidden effect of time. Statistics in Medicine, 7(6):629–637, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kjaersgaard-Andersen:1988:NME

[543] Per Kjærsgaard-Andersen, Finn Christensen, Steen Asmus Schmidt, Niels Wisbech Pedersen, and Bent Jøensen. A new method of estimation of interobserver variation and its application to the radiological assessment of osteoarthrosis in hip joints. Statistics in Medicine, 7(6):639–647, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stone:1988:IEE

[544] Richard A. Stone. Investigations of excess environmental risks around putative sources: Statistical problems and a proposed test. Statistics in Medicine, 7(6):649–660, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1988:MMS

[545] Per Kragh Andersen. Multistate models in survival analysis: a study of nephropathy and mortality in diabetes. Statistics in Medicine, 7(6):661– 670, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1988:MLF

[546] T. Timothy Chen. Modelling the liver function data in diabetes. Statistics in Medicine, 7(6):671–683, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vollmer:1988:LVC

[547] William M. Vollmer, Larry R. Johnson, Lynn E. McCamant, and A. Sonia Buist. Longitudinal versus cross-sectional estimation of lung function decline-further insights. Statistics in Medicine, 7(6):685–696, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marler:1988:BRS

[548] Matthew R. Marler, Rolf G. Jacob, John P. Lehoczky, and Alvin P. Shapiro. Book review: The statistical analysis of treatment effects in 24-hour ambulatory blood pressure recordings. Statistics in Medicine, 7 (6):697–716, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1988:EAR

[549] Jonas H. Ellenberg. The extrapolation of attributable risk to new populations. Statistics in Medicine, 7(6):717–725, June 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mf

[550] Anonymous. Masthead. Statistics in Medicine, 7(7):fmi, July 1988. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavori:1988:IAP

[551] Philip W. Lavori and Martin B. Keller. Improving the aggregate performance of psychiatric diagnostic methods when not all subjects receive the standard test. Statistics in Medicine, 7(7):727–737, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1988:IVE

[552] N. J. D. Nagelkerke, V. Fidler, and M. Buwalda. Instrumental variables in the evaluation of diagnostic test procedures when the true disease state is unknown. Statistics in Medicine, 7(7):739–744, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1988:VEE

[553] Sander Greenland. Variance estimation for epidemiologic effect estimates under misclassification. Statistics in Medicine, 7(7):745–757, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lemeshow:1988:SSR

[554] Stanley Lemeshow, David W. Hosmer, Jr., and Janelle Klar. Sample size requirements for studies estimating odds ratios or relative risks. Statistics in Medicine, 7(7):759–764, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirji:1988:EVA

[555] Karim F. Hirji, Robert M. Elashoff, Dan H. Moore II, and Deborah E. Bennett. Exact versus asymptotic analysis for a matched case-control study. Statistics in Medicine, 7(7):765–772, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1988:CIC

[556] James M. Robins. Confidence intervals for causal parameters. Statistics in Medicine, 7(7):773–785, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newman:1988:MPI

[557] Stephen C. Newman. A Markov process interpretation of Sullivan’s index of morbidity and mortality. Statistics in Medicine, 7(7):787–794, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katz:1988:CER

[558] Barry P. Katz. A computer enhanced runs test for smoothed data. Statistics in Medicine, 7(7):795–804, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunn:1988:ODD

[559] G. Dunn. ‘Optimal’ designs for drug, neurotransmitter and hormone receptor assays. Statistics in Medicine, 7(7):805–815, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1988:BRS

[560] Colin B. Begg. Book review: Statistical methods in medical research. P. Armitage and G. Berry, Blackwell Scientific Publications, Oxford, U.K., 1987. No. of pages: 559. Price £22.50. Statistics in Medicine, 7(7):817– 818, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1988:BRM

[561] John C. Duffy. Book review: Modern epidemiology. Kenneth J. Rothman, Little, Brown and Co, Boston, 1986. No. of pages: 358. Price: £33.90. Statistics in Medicine, 7(7):818, July 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mg

[562] Anonymous. Masthead. Statistics in Medicine, 7(8):fmi, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayton:1988:AEH

[563] David Clayton. The analysis of event history data: a review of progress and outstanding problems. Statistics in Medicine, 7(8):819–841, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Makuch:1988:IPF

[564] Robert W. Makuch, Philip S. Rosenberg, and James Mulshine. Identifying prognostic factors in binary outcome data: an application using liver function tests and age to predict liver metastases. Statistics in Medicine, 7(8):843–856, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:1988:NCC

[565] J. F. Bithell and C. A. Stiller. A new calculation of the carcinogenic risk of obstetric X-raying. Statistics in Medicine, 7(8):857–864, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bohning:1988:CIE

[566] Dankmar Böhning. Confidence interval estimation of a rate and the choice of sample size. Statistics in Medicine, 7(8):865–875, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elashoff:1988:DCT

[567] Janet D. Elashoff, Gary G. Koch, and George Y. H. Chi. Designing a clinical trial to demonstrate prevention of ulcer recurrence: Modelling simulation approaches. Statistics in Medicine, 7(8):877–888, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galbraith:1988:NGP

[568] R. F. Galbraith. A note on graphical presentation of estimated odds ratios from several clinical trials. Statistics in Medicine, 7(8):889–894, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anson:1988:PDC

[569] Jon Anson. The parameters of death: a consideration of the quantity of information in a life table using a polynomial representation of the survivorship curve. Statistics in Medicine, 7(8):895–912, August 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mh

[570] Anonymous. Masthead. Statistics in Medicine, 7(9):fmi, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:EA

[571] Anonymous. Editorial announcement. Statistics in Medicine, 7(9):913, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hayes:1988:MAW

[572] R. J. Hayes. Methods for assessing whether change depends on initial value. Statistics in Medicine, 7(9):915–927, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:1988:CIP

[573] S. C. Choi and D. M. Stablein. Comparing incomplete paired binomial data under non-random mechanisms. Statistics in Medicine, 7(9):929–939, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1988:CBC

[574] Gordon D. Murray and Janet G. Findlay. Correcting for the bias caused by drop-outs in hypertension trials. Statistics in Medicine, 7(9):941–946, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tiret:1988:POA

[575] L. Tiret, F. Hatton, J. M. Desmonts, and G. Vourc’h. Prediction of outcome of anaesthesia in patients over 40 years: a multifactorial risk index. Statistics in Medicine, 7(9):947–954, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Irwig:1988:ERR

[576] L. M. Irwig and H. T. Groeneveld. Exposure-response relationship for a dichotomized response when the continuous underlying variable is not measured. Statistics in Medicine, 7(9):955–964, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dersimonian:1988:TPM

[577] Rebecca Dersimonian and Stuart G. Baker. Two-process models for discrete-time serial categorical response. Statistics in Medicine, 7(9):965– 974, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blackwood:1988:LVM

[578] Larry Blackwood. Latent variable models for the analysis of medical data with repeated measures of binary variables. Statistics in Medicine, 7(9): 975–981, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leclerc:1988:CAL

[579] A. Leclerc, D. Luce, F. Lert, J. F. Chastang, and P. Logeay. Correspondence analysis and logistic modelling: Complementary use in the analysis of a health survey among nurses. Statistics in Medicine, 7(9):983–995, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gallus:1988:RES

[580] G. Gallus and Rina Chen. The relative efficiency of the sets and the cusum techniques in monitoring the occurrence of a rare event. Statistics in Medicine, 7(9):997–999, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1988:BRM

[581] Douglas G. Altman. Book review: Medical uses of statistics. John C. Bailar, III, and Frederick Mosteller (eds), Nejm books, Waltham mass., 1986. no. of pages: 425. price: £20. Statistics in Medicine, 7(9):1001, September 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mi

[582] Anonymous. Masthead. Statistics in Medicine, 7(10):fmi, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marks:1988:IBS

[583] Ronald G. Marks, Elizabeth K. Dawson-Saunders, John C. Bailar, Bruce B. Dan, and Joyce A. Verran. Interactions between statisticians and biomedical journal editors. Statistics in Medicine, 7(10):1003–1011, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1988:CED

[584] L. S. Freedman and B. L. De Stavola. Comparing the effects of different durations of the same therapy. Statistics in Medicine, 7(10):1013–1021, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Norwood:1988:SMP

[585] Paula K. Norwood and Allan R. Sampson. A statistical methodology for postmarketing surveillance of adverse drug reaction reports. Statistics in Medicine, 7(10):1023–1030, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:1988:IIR

[586] R. A. Parker and R. B. Rothenberg. Identifying important results from multiple statistical tests. Statistics in Medicine, 7(10):1031–1043, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1988:AMT

[587] Donna F. Stroup, Stephen B. Thacker, and Joy L. Herndon. Application of multiple time series analysis to the estimation of pneumonia and influenza mortality by age 1962–1983. Statistics in Medicine, 7(10):1045– 1059, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milner:1988:FDA

[588] P. C. Milner. Forecasting the demand on accident and emergency departments in health districts in the trent region. Statistics in Medicine, 7(10): 1061–1072, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1988:END

[589] Chung-Cheng Hsieh and Stephen D. Walter. The effect of non-differential exposure misclassification on estimates of the attributable and prevented fraction. Statistics in Medicine, 7(10):1073–1085, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1988:ENC

[590] Edward L. Korn and Barry I. Graubard. Examining neighbourhood confounding in a survey: an example using the national health and nutrition examination survey II. Statistics in Medicine, 7(10):1087–1098, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1988:CTC

[591] S. J. Senn. Cross-over trials, carry-over effects and the art of selfdelusion. Statistics in Medicine, 7(10):1099–1101, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Ad

[592] Anonymous. Announcement. Statistics in Medicine, 7(10):1103, October 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mj

[593] Anonymous. Masthead. Statistics in Medicine, 7(11):fmi, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1988:P

[594] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 7(11):1105, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:ISC

[595] Anonymous. International Society for Clinical Biostatistics, Eighth International Meeting held at Chalmers University of Technology, Gothenburg from 7–11 September 1987p. Statistics in Medicine, 7(11):1107, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holme:1988:DTM

[596] Ingar Holme. Drug treatment of mild hypertension to reduce the risk of CHD: Is it worth-while? Statistics in Medicine, 7(11):1109–1120, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1988:HSA

[597] Odd O. Aalen. Heterogeneity in survival analysis. Statistics in Medicine, 7(11):1121–1137, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:1988:RDE

[598] Gerta Rücker and Dorle Messerer. Remission duration: an example of interval-censored observations. Statistics in Medicine, 7(11):1139–1145, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flahault:1988:MIE

[599] A. Flahault, S. Letrait, P. Blin, S. Hazout, J. Ménarés, and A. J. Valleron. Modelling the 1985 influenza epidemic in France. Statistics in Medicine, 7(11):1147–1155, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Howel:1988:TDC

[600] Denise Howel. A test to detect clustering applied to muscle fibres. Statistics in Medicine, 7(11):1157–1164, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chilvers:1988:IQD

[601] C. E. D. Chilvers, P. M. Fayers, L. S. Freedman, R. M. Greenwood, D. Machin, N. Palmer, and A. J. Westlake. Improving the quality of data in randomized clinical trials: The compact computer package. compact steering committee. Statistics in Medicine, 7(11):1165–1170, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petri:1988:PSA

[602] H. Petri, H. C. W. De Vet, J. Naus, and J. Urquhart. Prescription sequence analysis: a new and fast method for assessing certain adverse reactions of prescription drugs in large populations. Statistics in Medicine, 7(11): 1171–1175, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1988:EPE

[603] R. G. Newcombe. Explanatory and pragmatic estimates of the treatment effect when deviations from allocated treatment occur. Statistics in Medicine, 7(11):1179–1186, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:1988:OPR

[604] Simon J. Day. Optimal placebo response rates for comparing two binomial proportions. Statistics in Medicine, 7(11):1187–1194, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1988:SSF

[605] F. Y. Hsieh. Sample size formulae for intervention studies with the cluster as unit of randomization. Statistics in Medicine, 7(11):1195–1201, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [855].

Senn:1988:HMP

[606] S. J. Senn. How much of the placebo ‘effect’ is really statistical regression? Statistics in Medicine, 7(11):1203, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1988:BRE

[607] G. Berry. Book review: Essentials of biostatistics, Robert C. Elston and William D. Johnson, F. A. Davis company, Philadelphia, 1987. no. of pages xii + 307. price: £19.65. Statistics in Medicine, 7(11):1204, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Ae

[608] Anonymous. Announcement. Statistics in Medicine, 7(11):1205, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1988:Mk

[609] Anonymous. Masthead. Statistics in Medicine, 7(12):fmi, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ozminkowski:1988:IOS

[610] Ronald J. Ozminkowski, Paul M. Wortman, and Dietrich W. Roloff. Inborn/outborn status and neonatal survival: a meta-analysis of nonrandomized studies. Statistics in Medicine, 7(12):1207–1221, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1988:BAC

[611] R. J. Marshall. Bayesian analysis of case-control studies. Statistics in Medicine, 7(12):1223–1230, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1988:SRE

[612] Michael D. Hughes and Stuart J. Pocock. Stopping rules and estimation problems in clinical trials. Statistics in Medicine, 7(12):1231–1242, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chastang:1988:QSB

[613] Claude Chastang, David Byar, and Steven Piantadosi. A quantitative study of the bias in estimating the treatment effect caused by omitting a balanced covariate in survival models. Statistics in Medicine, 7(12): 1243–1255, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1988:NPA

[614] M. Schemper. Non-parametric analysis of treatment-covariate interaction in the presence of censoring. Statistics in Medicine, 7(12):1257–1266, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mau:1988:SAC

[615] Jochen Mau. A statistical assessment of clinical equivalence. Statistics in Medicine, 7(12):1267–1277, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oneill:1988:SSE

[616] Robert T. O’neill. On sample sizes to estimate the protective efficacy of a vaccine. Statistics in Medicine, 7(12):1279–1288, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1988:BSS

[617] Richard H. Jones and Maureen W. McClatchey. Beyond sensitivity, specificity and statistical independence. Statistics in Medicine, 7(12):1289– 1295, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:1988:DOS

[618] Christopher J. Lloyd. Doubling the one-sided p-value in testing independence in 2 × 2 tables against a two-sided alternative. Statistics in Medicine, 7(12):1297–1306, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1988:BRS

[619] Patrick Royston. Book review: Statistical tables for the design of clinical trials. David Machin and Michael J. Campbell, Blackwell scientific publications, Oxford, 1987. no. of pages: xi + 210. price: £30. Statistics in Medicine, 7(12):1307, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:1988:BRI

[620] R. J. Prescott. Book review: An introduction to medical statistics. Martin Bland, Oxford University Press, 1987. no. of pages: xiii + 365. price: £9.95. Statistics in Medicine, 7(12):1307–1308, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1988:BRM

[621] M. J. R. Healy. Book review: Multivariate interpretation of clinical laboratory data. Adelin Albert and Eugene K. Harris, Marcel Dekker Inc., New York, 1987. no. of pages: 328. price: $69.75. Statistics in Medicine, 7(12):1308–1309, December 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Ma

[622] Anonymous. Masthead. Statistics in Medicine, 8(1):fmi, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:EN

[623] Anonymous. Editors’ note. Statistics in Medicine, 8(1):i, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:1989:P

[624] Ron Brookmeyer. Preface. Statistics in Medicine, 8(1):1–2, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeger:1989:SMM

[625] Scott L. Zeger, Lai-Chu See, and Peter J. Diggle. Statistical methods for monitoring the AIDS epidemic. Statistics in Medicine, 8(1):3–21, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:1989:SMS

[626] Ron Brookmeyer and Anne Damiano. Statistical methods for short-term projections of AIDS incidence. Statistics in Medicine, 8(1):23–34, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deGruttola:1989:VAI

[627] Victor de Gruttola and Stephen W. Lagakos. The value of AIDS incidence data in assessing the spread of HIV infection. Statistics in Medicine, 8 (1):35–43, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1989:MHI

[628] Jeremy M. G. Taylor. Models for the HIV infection and AIDS epidemic in the United States. Statistics in Medicine, 8(1):45–58, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1989:DPM

[629] Mitchell H. Gail, Dale Preston, and Steve Piantadosi. Disease prevention models of voluntary confidential screening for human immunodeficiency virus (HIV). Statistics in Medicine, 8(1):59–81, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eisenberg:1989:NPP

[630] Bennett Eisenberg. The number of partners and the probability of HIV infection. Statistics in Medicine, 8(1):83–92, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wiley:1989:HPH

[631] James A. Wiley, Stephen J. Herschkorn, and Nancy S. Padian. Heterogeneity in the probability of HIV transmission per sexual contact: The case of male-to-female transmission in penile-vaginal intercourse. Statistics in Medicine, 8(1):93–102, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bongaarts:1989:MSH

[632] John Bongaarts. A model of the spread of HIV infection and the demographic impact of AIDS. Statistics in Medicine, 8(1):103–120, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:1989:SSM

[633] W. Y. Tan and H. Hsu. Some stochastic models of AIDS spread. Statistics in Medicine, 8(1):121–136, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brandt:1989:PIM

[634] Edward N. Brandt. Policy implications of modelling of the AIDS epidemic. Statistics in Medicine, 8(1):137–139, January 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mb

[635] Anonymous. Masthead. Statistics in Medicine, 8(2):fmi, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berlin:1989:CSM

[636] Jesse A. Berlin, Nan M. Laird, Henry S. Sacks, and Thomas C. Chalmers. A comparison of statistical methods for combining event rates from clinical trials. Statistics in Medicine, 8(2):141–151, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emrich:1989:RDP

[637] Lawrence J. Emrich. Required duration and power determinations for historically controlled studies of survival times. Statistics in Medicine, 8 (2):153–160, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brittain:1989:FDC

[638] Erica Brittain and Janet Wittes. Factorial designs in clinical trials: The effects of non-compliance and subadditivity. Statistics in Medicine, 8(2):

161–171, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:1989:IAE

[639] E. S. Gilbert. Issues in analysing the effects of occupational exposure to low levels of radiation. Statistics in Medicine, 8(2):173–187, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vacek:1989:ATS

[640] Pamela M. Vacek, Ruth M. Mickey, and Dianne Y. Bell. Application of a two-stage random effects model to longitudinal pulmonary function data from sarcoidosis patients. Statistics in Medicine, 8(2):189–200, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verdecchia:1989:MEC

[641] Arduino Verdecchia, Riccardo Capocaccia, Viviana Egidi, and Antonio Golini. A method for the estimation of chronic disease morbidity and trends from mortality data. Statistics in Medicine, 8(2):201–216, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azen:1989:EPM

[642] Stanley P. Azen, Michael van Guilder, and Mary Ann Hill. Estimation of parameters and missing values under a regression model with nonnormally distributed and non-randomly incomplete data. Statistics in Medicine, 8(2):217–228, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katz:1989:VEM

[643] Barry P. Katz and Siu L. Hui. Variance estimation for medical decision analysis. Statistics in Medicine, 8(2):229–241, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dupont:1989:LE

[644] W. D. Dupont. Letter to the editor. Statistics in Medicine, 8(2):243–245, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:1989:BRD

[645] Louise M. Ryan. Book review: The design and analysis of long-term animal experiments. J. J. Gart, D. Krewski, P. N. Lee, R. E. Tarone and J. Wahrendorf. IARC Scientific Publications 79, Oxford University Press/ IARC, 1986. No. of pages: 219. Price: £20. Statistics in Medicine, 8(2):

246, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seldrup:1989:A

[646] J. Seldrup. Announcement. Statistics in Medicine, 8(2):247, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:C

[647] Anonymous. Corrections. Statistics in Medicine, 8(2):248, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Aa

[648] Anonymous. Acknowledgement. Statistics in Medicine, 8(2):249, February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mc

[649] Anonymous. Masthead. Statistics in Medicine, 8(3):fmi, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mason:1989:OR

[650] James O. Mason. Opening remarks. Statistics in Medicine, 8(3):253– 254, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1989:R

[651] Theodore Colton. Remarks. Statistics in Medicine, 8(3):255–256, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1989:AMM

[652] Kung-Jong Lui and Richard K. Rudy. An application of a mathematical model to adjust for time lag in case reporting. Statistics in Medicine, 8 (3):259–262, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conn:1989:MBA

[653] Judith M. Conn, Kung-Jong Lui, and Daniel L. McGee. A model-based approach to the imputation of missing data: Home injury incidences. Statistics in Medicine, 8(3):263–266, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raubertas:1989:ADS

[654] Richard F. Raubertas. An analysis of disease surveillance data that uses the geographic locations of the reporting units. Statistics in Medicine, 8 (3):267–271, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hall:1989:ASD

[655] David B. Hall. Analysis of surveillance data: a rationale for statistical tests with comments on confidence intervals and statistical models. Statistics in Medicine, 8(3):273–278, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1989:C

[656] Donna F. Stroup. Commentary. Statistics in Medicine, 8(3):279–281, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:1989:ASD

[657] R. A. Parker. Analysis of surveillance data with Poisson regression: a case study. Statistics in Medicine, 8(3):285–294, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:1989:CBP

[658] Samuel P. Caudill, S. Jay Smith, and Gerald R. Cooper. Cholesterol-based personal risk assessment in coronary heart disease. Statistics in Medicine, 8(3):295–309, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shore:1989:SSB

[659] David Lee Shore and Dana Quade. A surveillance system based on a short memory scheme. Statistics in Medicine, 8(3):311–322, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1989:DAO

[660] Donna F. Stroup, G. David Williamson, Joy L. Herndon, and John M. Karon. Detection of aberrations in the occurrence of notifiable diseases surveillance data. Statistics in Medicine, 8(3):323–329, March 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lane:1989:C

[661] Mike Lane. Commentary. Statistics in Medicine, 8(3):331–332, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katzoff:1989:ATS

[662] Myron Katzoff. The application of time series forecasting methods to an estimation problem using provisional mortality statistics. Statistics in Medicine, 8(3):335–341, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schnell:1989:TSA

[663] Daniel Schnell, Akbar Zaidi, and Gladys Reynolds. A time series analysis of gonorrhea surveillance data. Statistics in Medicine, 8(3):343–352, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zaidi:1989:TSA

[664] Akbar A. Zaidi, Daniel J. Schnell, and Gladys H. Reynolds. Time series analysis of syphilis surveillance data. Statistics in Medicine, 8(3):353– 362, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thacker:1989:C

[665] Stephen B. Thacker. Commentary. Statistics in Medicine, 8(3):363, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Massey:1989:MDQ

[666] James T. Massey and Keith L. Hoffman. Monitoring data quality through comparisons between data systems. Statistics in Medicine, 8(3):367–377, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McGee:1989:ETE

[667] Daniel L. McGee and Phil Rhodes. Estimating trends in the effectiveness of seat belts in saving lives, 1975–1985. Statistics in Medicine, 8(3):379– 385, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpson:1989:SAN

[668] Gloria Simpson and Paul J. Placek. Surveillance aspects of the 1988 National Maternal and Infant Health Survey. Statistics in Medicine, 8 (3):387–388, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Middleton:1989:EMR

[669] Dannie C. Middleton and Roger B. Trent. Emergency medical response to occupational locations in West Virginia. Statistics in Medicine, 8(3): 388, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stout-Wiegand:1989:FOI

[670] Nancy Stout-Wiegand. Fatal occupational injuries in U.S. industries, 1984: Comparison of two national surveillance systems. Statistics in Medicine, 8(3):389–390, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rokaw:1989:CUN

[671] William M. Rokaw, James A. Mercy, and Jack C. Smith. Comparability and utility of national homicide data from death certificates and police records. Statistics in Medicine, 8(3):390, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brackbill:1989:SOC

[672] Robert Brackbill and Lawrence R. Murphy. Surveillance of occupational cardiovascular disease through the linkage of three data bases. Statistics in Medicine, 8(3):391, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knutson:1989:SHB

[673] Lisa Knutson and David Hall. Screening of hepatitis B surface antigen carriers for primary hepatocellular carcinoma using alpha-fetoprotein. Statistics in Medicine, 8(3):391–392, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berkelman:1989:SR

[674] Ruth L. Berkelman. Summary remarks. Statistics in Medicine, 8(3):393– 395, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reynolds:1989:CR

[675] Gladys H. Reynolds. Closing remarks. Statistics in Medicine, 8(3):397– 400, March 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Md

[676] Anonymous. Masthead. Statistics in Medicine, 8(4):fmi, April 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1989:Pa

[677] Theodore Colton, Laurence Freedman, Tony Johnson, and David MaChin. Preface. Statistics in Medicine, 8(4):401, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herson:1989:USE

[678] Jay Herson. The use of surrogate endpoints in clinical trials (an introduction to a series of four papers). Statistics in Medicine, 8(4):403–404, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1989:SEC

[679] Susan S. Ellenberg and J. Michael Hamilton. Surrogate endpoints in clinical trials: Cancer. Statistics in Medicine, 8(4):405–413, April 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1989:SEC

[680] Janet Wittes, Edward Lakatos, and Jeffrey Probstfield. Surrogate endpoints in clinical trials: Cardiovascular diseases. Statistics in Medicine, 8(4):415–425, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:1989:SEC

[681] Argye Hillis and Daniel Seigel. Surrogate endpoints in clinical trials: Ophthalmologic disorders. Statistics in Medicine, 8(4):427–430, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prentice:1989:SEC

[682] Ross L. Prentice. Surrogate endpoints in clinical trials: Definition and operational criteria. Statistics in Medicine, 8(4):431–440, April 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colditz:1989:HSD

[683] Graham A. Colditz, James N. Miller, and Frederick Mosteller. How study design affects outcomes in comparisons of therapy. i: Medical. Statistics in Medicine, 8(4):441–454, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:1989:HSD

[684] James N. Miller, Graham A. Colditz, and Frederick Mosteller. How study design affects outcomes in comparisons of therapy. II: Surgical. Statistics in Medicine, 8(4):455–466, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1989:CIR

[685] S. J. Senn. Covariate imbalance and random allocation in clinical trials. Statistics in Medicine, 8(4):467–475, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:1989:TST

[686] Gerta Rücker. A two-stage trial design for testing treatment, self-selection and treatment preference effects. Statistics in Medicine, 8(4):477–485, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1989:TPC

[687] Gary G. Koch, Ingrid A. Amara, Byron W. Brown, Jr., Theodore Colton, and Dennis B. Gillings. A two-period crossover design for the comparison of two active treatments and placebo. Statistics in Medicine, 8(4):487– 504, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1989:IGS

[688] Gary L. Rosner and Anastasios A. Tsiatis. The impact that group sequential tests would have made on ECOG clinical trials. Statistics in Medicine, 8(4):505–516, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [854].

Niland:1989:LE

[689] Joyce C. Niland, Duncan Thomas, and Stanley P. Azen. Letters to the editor. Statistics in Medicine, 8(4):517–519, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charlton:1989:BRU

[690] Anne Charlton. Book review: U.K. Smoking statistics. N. Wald, S. Kiryluk, S. Darby, R. Doll, M. Pike, and R. Peto, Oxford University Press, Oxford, 1988. No. of pages: xxx + 197. Price: £30. Statistics in Medicine, 8(4):521, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cross:1989:BRS

[691] K. W. Cross. Book review: Statistics for health management and research. Mark Woodward and Lesley M. A. Francis, Edward Arnold, London, 1988. No. of pages: 346. Price: £13.95. Statistics in Medicine, 8(4):522, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Everitt:1989:BRA

[692] B. S. Everitt. Book review: Applied categorical data analysis. Daniel H. Freeman, Jr., Marcel Dekker Inc, 1987. No. of pages: xi + 318. Price: £54.85 ($69.75). Statistics in Medicine, 8(4):523, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirkwood:1989:BRS

[693] T. B. L. Kirkwood. Book review: Some mathematical questions in biology: DNA sequence analysis. Volume 17, Lectures on Mathematics in the Life Sciences, Robert M. Miura, (Ed.), American Mathematical Society, Providence, Rhode Island, 1986. No. of pages: x + 124. Price: $29. Statistics in Medicine, 8(4):523–524, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Me

[694] Anonymous. Masthead. Statistics in Medicine, 8(5):fmi, May 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ingram:1989:ECP

[695] Deborah D. Ingram and Joel C. Kleinman. Empirical comparisons of proportional hazards and logistic regression models. Statistics in Medicine, 8(5):525–538, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segal:1989:CEP

[696] Mark R. Segal and Daniel A. Bloch. A comparison of estimated proportional hazards models and regression trees. Statistics in Medicine, 8(5): 539–550, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Durrleman:1989:FRM

[697] Sylvain Durrleman and Richard Simon. Flexible regression models with cubic splines. Statistics in Medicine, 8(5):551–561, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:1989:BSP

[698] Myron N. Chang, Harry S. Wieand, and Vivian T. Chang. The bias of the sample proportion following a group sequential phase II clinical trial. Statistics in Medicine, 8(5):563–570, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francom:1989:LLM

[699] S. F. Francom, C. Chuang-Stein, and J. R. Landis. A log-linear model for ordinal data to characterize differential change among treatments. Statistics in Medicine, 8(5):571–582, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1989:ATS

[700] Weichung Joseph Shih, A. Lawrence Gould, and Irving K. Hwang. The analysis of titration studies in phase III clinical trials. Statistics in Medicine, 8(5):583–591, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodary:1989:HEE

[701] Chantal Rodary, Catherine Com-Nougue, and Marie-France Tournade. How to establish equivalence between treatments: a one-sided clinical trial in paediatric oncology. Statistics in Medicine, 8(5):593–598, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1989:ECO

[702] Allan Donner and Walter Hauck. Estimation of a common odds ratio in paired-cluster randomization designs. Statistics in Medicine, 8(5):599– 607, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Linnet:1989:ADT

[703] Kristian Linnet. Assessing diagnostic tests by a strictly proper scoring rule. Statistics in Medicine, 8(5):609–618, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Makuch:1989:GKN

[704] Robert W. Makuch, Philip S. Rosenberg, and Gwendolyn Scott. Goodman and Kruskal’s λ: a new look at an old measure of association. Statistics in Medicine, 8(5):619–631, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:LE

[705] Anonymous. Letter to the editor. Statistics in Medicine, 8(5):633, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carstensen:1989:BRE

[706] Bendix Carstensen. Book review: European community atlas of ‘avoidable death’. W. W. Holland (Ed), Oxford University Press, Oxford, 1988. No. of pages: xxiii + 356. Price: £40. Statistics in Medicine, 8(5):635, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1989:BRE

[707] J. N. S. Matthews. Book review: Essential medical statistics. Betty R. Kirkwood, Blackwell Scientific Publications, Oxford, 1988. No. of pages:

234. Price: £9.95. Statistics in Medicine, 8(5):636, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:1989:BRU

[708] J. Martin Bland. Book review: Using and understanding medical statistics. (2nd edition) D. F. Matthews and V. T. Farewell, Karger, Basel, Switzerland, 1988. No. of pages: xii + 228. Price £18.20 ($26.75). Statistics in Medicine, 8(5):636–637, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coles:1989:BRH

[709] E. C. Coles. Book review: Handbook of record linkage methods for health and statistical studies, administration, and business. Howard B. Newcombe, Oxford University Press, Oxford, 1988. No. of pages: vi + 210. Price: £22.50. Statistics in Medicine, 8(5):638, May 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mf

[710] Anonymous. Masthead. Statistics in Medicine, 8(6):fmi, June 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1989:Pb

[711] Theodore Colton, Laurence Freedman, Tony Johnson, and David MaChin. Preface. Statistics in Medicine, 8(6):639, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dupont:1989:CRR

[712] William D. Dupont. Converting relative risks to absolute risks: a graphical approach. Statistics in Medicine, 8(6):641–651, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michalek:1989:CRE

[713] Joel E. Michalek, Daniel Mihalko, and Ram C. Tripathi. Cautions on the reanalysis of epidemiologic databases. Statistics in Medicine, 8(6):653– 664, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mantel:1989:C

[714] Nathan Mantel. Comment. Statistics in Medicine, 8(6):665–667, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1989:CCU

[715] Sander Greenland. Comment: Cautions in the use of preliminary-test estimators. Statistics in Medicine, 8(6):669–673, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michalek:1989:AR

[716] Joel E. Michalek, Daniel Mihalko, and Ram C. Tripathi. Authors’ reply. Statistics in Medicine, 8(6):675–677, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1989:CCI

[717] James Robins. The control of confounding by intermediate variables. Statistics in Medicine, 8(6):679–701, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elandt-Johnson:1989:GGR

[718] Regina C. Elandt-Johnson and Fraser B. Smith. Graphical generalized residuals in fitting distributions: Applications to epidemiological follow-up data. Statistics in Medicine, 8(6):703–723, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lassauzet:1989:RMT

[719] Marie-Liesse G. Lassauzet, Wesley O. Johnson, and Mark C. Thurmond. Regression models for time to seroconversion following experimental bovine leukaemia virus infection. Statistics in Medicine, 8(6):725–741, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alexander:1989:EST

[720] Freda E. Alexander. Estimation of sojourn time distributions and false negative rates in screening programmes which use two modalities. Statistics in Medicine, 8(6):743–755, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suissa:1989:UMC

[721] Samy Suissa and Rachid Salmi. Unidirectional multiple comparisons of Poisson rates. Statistics in Medicine, 8(6):757–764, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:1989:LE

[722] Bent Jørgensen and Per Kjærsgaard-Andersen. Letter to the editor. Statistics in Medicine, 8(6):765–767, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1989:BRS

[723] Mitchell H. Gail. Book review: Statistical methods in cancer research. Volume II. The design and analysis of cohort studies. N. E. Breslow and N. E. Day, Oxford University Press for International Agency for Research on Cancer, 1987. No of pages: xii + 406. Price: £30. Statistics in Medicine, 8(6):769–770, June 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mg

[724] Anonymous. Masthead. Statistics in Medicine, 8(7):fmi, July 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1989:BIS

[725] Douglas G. Altman and Per Kragh Andersen. Bootstrap investigation of the stability of a Cox regression model. Statistics in Medicine, 8(7):771– 783, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hastie:1989:ROC

[726] T. J. Hastie, J. L. Botha, and C. M. Schnitzler. Regression with an ordered categorical response. Statistics in Medicine, 8(7):785–794, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1989:SST

[727] F. Y. Hsieh. Sample size tables for logistic regression. Statistics in Medicine, 8(7):795–802, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bristol:1989:SSC

[728] David R. Bristol. Sample sizes for constructing confidence intervals and testing hypotheses. Statistics in Medicine, 8(7):803–811, July 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mather:1989:MED

[729] Frances J. Mather, Richard M. Simon, Gary M. Clark, and Daniel D. Von Hoff. A method for the evaluation of dose-toxicity relationships in clinical trials. Statistics in Medicine, 8(7):813–823, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1989:CMT

[730] Sander Greenland. On correcting for misclassification in twin studies and other matched-pair studies. Statistics in Medicine, 8(7):825–829, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1989:SAS

[731] Ira M. Longini, Jr., W. Scott Clark, Robert H. Byers, John W. Ward, William W. Darrow, George F. Lemp, and Herbert W. Hethcote. Statistical analysis of the stages of HIV infection using a Markov model. Statistics in Medicine, 8(7):831–843, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1989:EEE

[732] James M. Robins and Sander Greenland. Estimability and estimation of excess and etiologic fractions. Statistics in Medicine, 8(7):845–859, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:1989:CIT

[733] Lynette L-Y Lim and John Whitehead. Comparison of the information in two lung function experiments. Statistics in Medicine, 8(7):861–870, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cologne:1989:SCL

[734] John B. Cologne, Paul M. Mendelman, and Donald O. Chaffin. Statistical comparison of ligand-binding kinetics. Statistics in Medicine, 8(7):871– 881, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:1989:TBB

[735] Shein-Chung Chow and Jun Shao. Test for batch-to-batch variation in stability analysis. Statistics in Medicine, 8(7):883–890, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gough:1989:SAS

[736] K. Gough and K. B. Hafner. Some analysis strategies for three period changeover designs with two treatments. Statistics in Medicine, 8(7):891– 892, July 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mh

[737] Anonymous. Masthead. Statistics in Medicine, 8(8):fmi, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1989:MME

[738] M. J. R. Healy. Measuring measuring errors. Statistics in Medicine, 8 (8):893–906, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1989:LRM

[739] Odd O. Aalen. A linear regression model for the analysis of life times. Statistics in Medicine, 8(8):907–925, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barlow:1989:BCI

[740] William E. Barlow and Wei-Huei Sun. Bootstrapped confidence intervals for the Cox model using a linear relative risk form. Statistics in Medicine, 8(8):927–935, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beltangady:1989:EUC

[741] Mohan S. Beltangady and Ralph F. Frankowski. Effect of unequal censoring on the size and power of the logrank and Wilcoxon types of tests for survival data. Statistics in Medicine, 8(8):937–945, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1989:EST

[742] Roger B. Davis and James R. Anderson. Exponential survival trees. Statistics in Medicine, 8(8):947–961, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:1989:CRA

[743] S. H. Heisterkamp, J. C. Jager, E. J. Ruitenberg, J. A. M. Van Druten, and A. M. Downs. Correcting reported aids incidence: a statistical approach. Statistics in Medicine, 8(8):963–976, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raboud:1989:EGA

[744] J. Raboud and N. E. Breslow. Efficiency gains from the addition of controls to matched sets in cohort studies. Statistics in Medicine, 8(8):977–985, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1989:CML

[745] Stan Becker. A comparison of maximum likelihood and Jewell’s estimators of the odds ratio and relative risk in single 2 × 2 tables. Statistics in Medicine, 8(8):987–996, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machlin:1989:VMA

[746] Steven R. Machlin, Joel C. Kleinman, and Jennifer H. Madans. Validity of mortality analysis based on retrospective smoking information. Statistics in Medicine, 8(8):997–1009, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelterman:1989:TCM

[747] Daniel Zelterman. Testing for central mixtures of compartment model parameters. Statistics in Medicine, 8(8):1011–1021, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nurminen:1989:BAC

[748] Markku Nurminen, Pertti Mutanen, and Roger J. Marshall. Bayesian analysis of case-control studies. Statistics in Medicine, 8(8):1023–1024, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:1989:BRS

[749] Gillian M. Raab. Book review: Statistical techniques in bioassay. Z. Govindarajulu, Karger, Basel, 1988, No. of pages: xiv + 166. Price: £31.40 ($46). Statistics in Medicine, 8(8):1025, August 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mi

[750] Anonymous. Masthead. Statistics in Medicine, 8(9):fmi, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1989:IEV

[751] David P. Byar and Mitchell H. Gail. Introduction. errors-in-variables workshop. Statistics in Medicine, 8(9):1027–1029, September 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willett:1989:OIR

[752] Walter Willett. An overview of issues related to the correction of nondifferential exposure measurement error in epidemiologic studies. Statistics in Medicine, 8(9):1031–1040, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1989:UDM

[753] James R. Marshall. The use of dual or multiple reports in epidemiologic studies. Statistics in Medicine, 8(9):1041–1049, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1989:CLR

[754] B. Rosner, W. C. Willett, and D. Spiegelman. Correction of logistic regression relative risk estimates and confidence intervals for systematic within-person measurement error. Statistics in Medicine, 8(9):1051–1069, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Da

[755] Anonymous. Discussion 1. Statistics in Medicine, 8(9):1071–1073, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caroll:1989:CAG

[756] Raymond J. Caroll. Covariance analysis in generalized linear measurement error models. Statistics in Medicine, 8(9):1075–1093, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1989:RMM

[757] T. Timothy Chen. A review of methods for misclassified categorical data in epidemiology. Statistics in Medicine, 8(9):1095–1106, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Db

[758] Anonymous. Discussion 2. Statistics in Medicine, 8(9):1107–1108, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasabelnaby:1989:IAP

[759] Nancy A. Hasabelnaby, James H. Ware, and Wayne A. Fuller. Indoor air pollution and pulmonary performance: Investigating errors in exposure assessment. Statistics in Medicine, 8(9):1109–1126, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gleser:1989:CIA

[760] Leon Jay Gleser. Commentary on “Indoor air pollution and pulmonary performance: Investigating errors in exposure assessment”. Statistics in Medicine, 8(9):1127–1131, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasabelnaby:1989:RCL

[761] Nancy A. Hasabelnaby, James H. Ware, and Wayne A. Fuller. Rejoinder to comments by Leon Jay Gleser. Statistics in Medicine, 8(9):1133–1135, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Dc

[762] Anonymous. Discussion 3. Statistics in Medicine, 8(9):1137–1138, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tosteson:1989:MEM

[763] Tor D. Tosteson, Leonard A. Stefanski, and Daniel W. Schafer. A measurement-error model for binary and ordinal regression. Statistics in Medicine, 8(9):1139–1147, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Dd

[764] Anonymous. Discussion 4. Statistics in Medicine, 8(9):1149, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:1989:ACC

[765] Ben G. Armstrong, Alice S. Whittemore, and Geoff R. Howe. Analysis of case-control data with covariate measurement error: Application to diet and colon cancer. Statistics in Medicine, 8(9):1151–1163, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:De

[766] Anonymous. Discussion 5. Statistics in Medicine, 8(9):1165–1166, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:1989:FRC

[767] Margaret Sullivan Pepe, Steven G. Self, and Ross L. Prentice. Further results on covariate measurement errors in cohort studies with time to response data. Statistics in Medicine, 8(9):1167–1178, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Df

[768] Anonymous. Discussion 6. Statistics in Medicine, 8(9):1179, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Ab

[769] Anonymous. Announcement. Statistics in Medicine, 8(9):1181, September 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mj

[770] Anonymous. Masthead. Statistics in Medicine, 8(10):fmi, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaiser:1989:ABC

[771] Lee Kaiser. Adjusting for baseline: Change or percentage change? Statistics in Medicine, 8(10):1183–1190, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:1989:GSP

[772] K. K. Gordon Lan and David L. Demets. Group sequential procedures: Calendar versus information time. Statistics in Medicine, 8(10):1191– 1198, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1989:UAM

[773] Mark P. Becker. Using association models to analyse agreement data: Two examples. Statistics in Medicine, 8(10):1199–1207, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:1989:SMR

[774] Alan Agresti. A survey of models for repeated ordered categorical response data. Statistics in Medicine, 8(10):1209–1224, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lew:1989:EPB

[775] Robert A. Lew and Paul S. Levy. Estimation of prevalence on the basis of screening tests. Statistics in Medicine, 8(10):1225–1230, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1989:DCE

[776] Kung-Jong Lui. A discussion on the conventional estimator of sensitivity and specificity in multiple tests. Statistics in Medicine, 8(10):1231–1240, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nurminen:1989:AEC

[777] Markku Nurminen. Analysis of epidemiologic case-base studies for binary data. Statistics in Medicine, 8(10):1241–1254, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cnaan:1989:SAN

[778] Avital Cnaan and Louise Ryan. Survival analysis in natural history studies of disease. Statistics in Medicine, 8(10):1255–1268, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elashoff:1989:NCL

[779] R. M. Elashoff, J. J. Lee, and A. A. Afifi. A note on confidence limits for quartiles with right censored data. Statistics in Medicine, 8(10):1269– 1276, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1989:SAR

[780] M. L. Thompson and W. Zucchini. On the statistical analysis of ROC curves. Statistics in Medicine, 8(10):1277–1290, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McLachlan:1989:MMP

[781] G. J. McLachlan and R. D. Gordon. Mixture models for partially unclassified data: a case study of renal venous renin in hypertension. Statistics in Medicine, 8(10):1291–1300, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McDonald:1989:HMP

[782] Clement J. McDonald and George P. McCabe. How much of the placebo ‘effect’ is really statistical regression? Statistics in Medicine, 8(10):1301– 1302, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1989:CBC

[783] J. A. Lewis, Gordon D. Murray, and Janet G. Findlay. Correcting for the bias caused by dropouts in hypertension trials. Statistics in Medicine, 8(10):1302–1304, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1989:IIR

[784] P. Royston, R. A. Parker, and R. B. Rothenberg. Identifying important results from multiple statistical tests. Statistics in Medicine, 8(10):1304– 1306, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1989:BRR

[785] M. J. R. Healy. Book review: Renal transplantation: Sense and sensitization. Sheila M. Gore and Benjamin A. Bradley, Kluwer, Dordrecht, 1988. No. of pages: xxii + 311. Price: £63 or $106. Statistics in Medicine, 8(10): 1307–1308, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ebrahim:1989:BRM

[786] Shah Ebrahim. Book review: Measuring health: a guide to rating scales and questionnaires. Ian McDowell and Claire Newell, Oxford University Press, 1987, No. of pages: xiv + 342. Price: £35. Statistics in Medicine, 8(10):1308–1309, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dore:1989:BRS

[787] Caroline J. Doré. Book review: Statistical methods for the analysis of biomedical data. Robert F. Woolson, Wiley, Chichester, 1987. No. of pages:

513. Price: £45. Statistics in Medicine, 8(10):1309–1310, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:BRGa

[788] Anonymous. Book review: Guidelines for clinical trials in cancer (2nd edition) R. J. Fisher, A. S. Coates and J. H. Colebatch (Eds), Clinical Oncological Society of Australia, GPO Box 4708, Sydney, NSW 2001. 1987. No. of pages: xv + 124. Price: $12 Australian. Statistics in Medicine, 8(10):1310–1311, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:BRC

[789] Anonymous. Book review: Clinical trial protocol notes (2nd Printing), Wessex Medico-Pharmaceutical Group. Wessex Medical School Trust, Room LF67, Southampton General Hospital S09 4XY. 1988. No. of pages:

115. Price: £4. Statistics in Medicine, 8(10):1310–1311, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:1989:BRC

[790] Stephen Evans. Book review: Clinical trial procedure — notes for doctors. A. J. Munro, The Association for Clinical Research. 11 Uxbridge Street, Kensington, London W8 7TQ. 1988. No. of pages: 72, Price: £9.25. Statistics in Medicine, 8(10):1310–1311, October 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mk

[791] Anonymous. Masthead. Statistics in Medicine, 8(11):fmi, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1989:Pc

[792] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 8(11):1313, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:ISC

[793] Anonymous. International Society for Clinical Biostatistics Ninth International Meeting held at Innsbruck, Austria from 30 August–2 September 1988. Statistics in Medicine, 8(11):1315, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1989:OLR

[794] Deborah Ashby, Christopher R. West, and David Ames. The ordered logistic regression model in psychiatry: Rising prevalence of dementia in old people’s homes. Statistics in Medicine, 8(11):1317–1326, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Racine-Poon:1989:PRP

[795] A. Racine-Poon and J. P. Dubois. Predicting the range of plasma carbamazepine concentrations in patients with epilepsy. Statistics in Medicine, 8(11):1327–1337, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1989:UBC

[796] S. J. Senn. The use of baselines in clinical trials of bronchodilators. Statistics in Medicine, 8(11):1339–1350, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crichton:1989:CAS

[797] Nicola J. Crichton and John P. Hinde. Correspondence analysis as a screening method for indicants for clinical diagnosis. Statistics in Medicine, 8(11):1351–1362, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seeber:1989:RAT

[798] G. U. H. Seeber. On the regression analysis of tumour recurrence rates. Statistics in Medicine, 8(11):1363–1369, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaldor:1989:ETE

[799] J. M. Kaldor and N. E. Day. Estimation of temporal effects in treatmentinduced second cancer. Statistics in Medicine, 8(11):1371–1380, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:1989:CCP

[800] Helmut Schäfer. Constructing a cut-off point for a quantitative diagnostic test. Statistics in Medicine, 8(11):1381–1391, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ulm:1989:AAD

[801] K. Ulm, G. Schmidt, P. Barthel, and A. Linzmaier. Assessing antiarrhythmic drug efficacy: Multivariate considerations. Statistics in Medicine, 8 (11):1393–1400, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schulgen:1989:MSR

[802] Gabi Schulgen. A measure of similarity for response curves based on ranks. Statistics in Medicine, 8(11):1401–1411, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boen:1989:UVM

[803] James R. Boen. Understanding P -value misuse. Statistics in Medicine, 8(11):1413–1414, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Worsley:1989:TCP

[804] K. J. Worsley, D. Y. Downham, and J. M. Simpkin. Testing for a changepoint in registry data with an example on hypospadias. Statistics in Medicine, 8(11):1414–1416, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dixon:1989:NME

[805] M. F. Dixon and N. C. Smeeton. A new method of estimation of interobserver variation and its application to the radiological assessment of osteoarthrosis in hip joints. Statistics in Medicine, 8(11):1416–1417, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:BRGb

[806] Anonymous. Book review: Glim: an introduction. M. J. R. Healy. Oxford University Press, Oxford, 1988. No. of pages: 130 + ix. Price: £20.00. Statistics in Medicine, 8(11):1418–1419, November 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1989:BRS

[807] Susan R. Wilson. Book review: Statistical modelling in glim. M. Aitkin, D. Anderson, B. Francis and J. Hinde. Oxford University Press, Oxford, 1989. No. of pages: 374 + xi. Price: £35.00 (hard cover); £17.50 (soft cover). Statistics in Medicine, 8(11):1418–1419, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Ac

[808] Anonymous. Announcement. Statistics in Medicine, 8(11):1420, November 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Ml

[809] Anonymous. Masthead. Statistics in Medicine, 8(12):fmi, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freeman:1989:PTS

[810] P. R. Freeman. The performance of the two-stage analysis of twotreatment, two-period crossover trials. Statistics in Medicine, 8(12):1421– 1432, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seigel:1989:GIC

[811] Daniel Seigel and Roy C. Milton. Grading of images in a clinical trial. Statistics in Medicine, 8(12):1433–1438, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1989:ARC

[812] John Whitehead. The analysis of relapse clinical trials, with application to a comparison of two ulcer treatments. Statistics in Medicine, 8(12): 1439–1454, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:1989:CTH

[813] S. R. Paul and A. Donner. A comparison of tests of homogeneity of odds ratios in K 2×2 tables. Statistics in Medicine, 8(12):1455–1468, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnard:1989:AGP

[814] G. A. Barnard. On alleged gains in power from lower P -values. Statistics in Medicine, 8(12):1469–1477, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gibson:1989:SDC

[815] Janet M. Gibson and John E. Overall. The superiority of a drug combination over each of its components. Statistics in Medicine, 8(12):1479–1484, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:1989:CCC

[816] Ayala Cohen. Comparison of correlated correlations. Statistics in Medicine, 8(12):1485–1495, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Schnell:1989:TSA

[817] Beverly Martinez-Schnell and Akbar Zaidi. Time series analysis of injuries. Statistics in Medicine, 8(12):1497–1508, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1989:EHA

[818] Toshiro Tango. Estimation of haemophilia-associated AIDS incidence in Japan using individual dates of diagnosis. Statistics in Medicine, 8(12): 1509–1514, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Annesi:1989:ELR

[819] Isabella Annesi, Thierry Moreau, and Joseph Lellouch. Efficiency of the logistic regression and Cox proportional hazards models in longitudinal studies. Statistics in Medicine, 8(12):1515–1521, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanek:1989:CNU

[820] Edward J. Stanek III, Susan S. Shetterley, Lindsay H. Allen, Gretel H. Pelto, and Adolfo Chavez. A cautionary note on the use of autoregressive models in analysis of longitudinal data. Statistics in Medicine, 8(12): 1523–1528, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1989:MMT

[821] S. W. Duffy, T. E. Rohan, and N. E. Day. Misclassification in more than one factor in a case-control study: a combination of Mantel–Haenszel and maximum likelihood approaches. Statistics in Medicine, 8(12):1529–1536, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levy:1989:SSC

[822] Paul S. Levy, Elena S. H. Yu, William T. Liu, Siu-Chi Wong, Ming-Yuan Zhang, Zheng-Yu Wang, and Robert Katzman. Single-stage cluster sampling with a telescopic respondent rule: a variation motivated by a survey of dementia in elderly residents of shanghai. Statistics in Medicine, 8(12):1537–1544, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1989:CAL

[823] S. C. Richardson and A. Leclerc. Correspondence analysis and logistic medelling: Complementary use in the analysis of a health survey among nurses. Statistics in Medicine, 8(12):1545–1547, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1989:ER

[824] Douglas G. Altman, P. J. Solomon, Niels G. Becker, and Niels Keiding. Extended reviews. Statistics in Medicine, 8(12):1549–1556, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1989:BRP

[825] Jørgen Hilden. Book review: Professional judgement. A reader in clinical decision making. Jack Dowie and Arthur Elstein, Cambridge University Press, Cambridge, 1988. No. of pages: xiii + 565. Price £37.50 (hardback), £13.50 (paperback). Statistics in Medicine, 8(12):1557–1558, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:1989:BRM

[826] Thomas A. Louis. Book review: Multivariate analysis of variance and repeated measures, a practical approach for behavioural scientists. D. J. Hand and C. C. Taylor, Chapman and Hall, London, 1987, No. of. pages: xiv + 262. Price: £25.00 (hard), £13.95 (paper). Statistics in Medicine, 8(12):1559, December 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1989:Mm

[827] Anonymous. Masthead. Statistics in Medicine, 9(1-2):fmi, January– February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1989:E

[828] Theodore Colton, Laurence Freedman, Tony Johnson, and David Machin. Editorial. Statistics in Medicine, 9(1-2):i, January–February 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Held:1989:P

[829] Peter Held, Janet Wittes, Jeanette Duggan, and Salim Yusuf. Preface. Statistics in Medicine, 9(1-2):1, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Held:1989:LSD

[830] Peter Held. List of speakers and discussants. Statistics in Medicine, 9 (1-2):3–4, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lenfant:1989:IR

[831] Claude Lenfant. Introductory remarks. Statistics in Medicine, 9(1-2): 5, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedewald:1989:CCT

[832] William T. Friedewald. Costs of clinical trials and the need for efficiency: a brief overview. Statistics in Medicine, 9(1-2):9–12, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meier:1989:PTE

[833] Paul Meier. Polio trial: an early efficient clinical trial. Statistics in Medicine, 9(1-2):13–16, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tognoni:1989:CGC

[834] Gianni Tognoni, Maria Grazia Franzosi, Silvio Garattini, Aldo Maggioni, Antonio Lotto, Francesco Mauri, and Fausto Rovelli. The case of gissi in changing the attitudes and practice of Italian cardiologists. Statistics in Medicine, 9(1-2):17–27, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buring:1989:CEC

[835] Julie E. Buring and Charles H. Hennekens. Cost and efficiency in clinical trials: The U.S. physicians’ health study. Statistics in Medicine, 9(1-2):29–33, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tielsch:1989:CEC

[836] James M. Tielsch and Keith P. West, Jr. Cost and efficiency considerations in community-based trials of vitamin A in developing countries. Statistics in Medicine, 9(1-2):35–43, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gaston:1989:SCA

[837] Marilyn Hughes Gaston and Joel Verter. Sickle cell anaemia trial. Statistics in Medicine, 9(1-2):45–51, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1989:FRC

[838] David P. Byar. Factorial and reciprocal control designs. Statistics in Medicine, 9(1-2):55–64, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1989:RIP

[839] Janet Wittes and Erica Brittain. The role of internal pilot studies in increasing the efficiency of clinical trials. Statistics in Medicine, 9(1-2):65– 72, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yusuf:1989:SPR

[840] Salim Yusuf, Peter Held, K. K. Teo, and Elizabeth Rudin Toretsky. Selection of patients for randomized controlled trials: Implications of wide or narrow eligibility criteria. Statistics in Medicine, 9(1-2):73–86, January– February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lang:1989:URE

[841] Janet M. Lang. The use of a run-in to enhance compliance. Statistics in Medicine, 9(1-2):87–95, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1989:EMS

[842] C. E. Davis. Efficient means of studying ancillary questions in clinical trials. Statistics in Medicine, 9(1-2):97–100, January–February 1989. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verter:1989:HMD

[843] Joel Verter. How much data should we collect in a randomized clinical trial? Statistics in Medicine, 9(1-2):103–113, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neaton:1989:EES

[844] James D. Neaton, Alain G. Duchene, Kenneth H. Svendsen, and Deborah Wentworth. An examination of the efficiency of some quality assurance methods commonly employed in clinical trials. Statistics in Medicine, 9(1-2):115–124, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peduzzi:1989:ECL

[845] Peter Peduzzi, Pamela Hartigan, and Gary Johnson. An evaluation of central laboratories in three VA cooperative studies. Statistics in Medicine, 9(1-2):125–136, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eaton:1989:CPS

[846] Terry Eaton and Craig M. Pratt. A clinic’s perspective on screening, recruitment and data collection. Statistics in Medicine, 9(1-2):137–144, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1989:DPC

[847] Susan S. Ellenberg. Discussion of papers on cost and efficiency of data collection in clinical trials. Statistics in Medicine, 9(1-2):145–151, January– February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Terrin:1989:EUE

[848] Michael L. Terrin. Efficient use of endpoints in clinical trials: a clinical perspective. Statistics in Medicine, 9(1-2):155–160, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prentice:1989:OEE

[849] Ross L. Prentice. Opportunities for enhancing efficiency and reducing cost in large scale disease prevention trials: a statistical perspective. Statistics in Medicine, 9(1-2):161–172, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Detsky:1989:UCE

[850] Allan S. Detsky. Using cost-effectiveness analysis to improve the efficiency of allocating funds to clinical trials. Statistics in Medicine, 9(1-2):173–184, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilhelmsen:1989:ARS

[851] Lars Wilhelmsen. The applicability of the results of streamlined trials to clinical practice. Statistics in Medicine, 9(1-2):185–191, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oconnell:1989:ARS

[852] John B. O’Connell and Jay W. Mason. The applicability of results of streamlined trials to clinical practice: The myocarditis treatment trial. Statistics in Medicine, 9(1-2):193–197, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hennekens:1989:CRC

[853] Charles Hennekens. Concluding remarks of the Chairman. Statistics in Medicine, 9(1-2):199, January–February 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Facey:1990:LEI

[854] Karen M. Facey and John Whitehead. Letter to the Editor: The impact that group sequential tests would have made on ECOG clinical trials. Gary Rosner and Anastasios A. Tsiatis, Statistics in Medicine, 8, 505– 516 (1989). Statistics in Medicine, 9(7):853–854, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [688].

Hsieh:1997:CSS

[855] F. Y. Hsieh. Correction: Sample Size Formulae for Intervention Studies with the Cluster as Unit of Randomization, by F. Y. Hsieh, Statistics in Medicine, 7, 1195–1201 (1988). Statistics in Medicine, 16(11):1300, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [605].